### v1

# One Model for All Tasks: Leveraging Efficient World Models in Multi-Task Planning

Abstract
In heterogeneous multi-task learning, tasks not only exhibit diverse observation and action spaces but also vary substantially in intrinsic difficulty. While conventional multi-task world models like UniZero excel in single-task settings, we find that when handling large-scale heterogeneous environments, gradient conflicts and the loss of model plasticity often constrain their sample and computational efficiency. In this work, we address these challenges from two perspectives: the single learning iteration and the overall learning process. First, we investigate the impact of key design spaces on extending UniZero to multi-task planning. We find that a Mixture-of-Experts (MoE) architecture provides the most substantial performance gains by mitigating gradient conflicts, leading to our proposed model, ScaleZero. Second, to dynamically balance the computational load across the learning process, we introduce an online, LoRA-based dynamic parameter scaling (DPS) strategy. This strategy progressively integrates LoRA adapters in response to task-specific progress, enabling adaptive knowledge retention and parameter expansion. Empirical evaluations on standard benchmarks such as Atari, DMControl (DMC), and Jericho demonstrate that ScaleZero, relying exclusively on online reinforcement learning with one model, attains performance on par with specialized single-task baselines. Furthermore, when augmented with our dynamic parameter scaling strategy, our method achieves competitive performance while requiring only 80% of the single-task environment interaction steps. These findings underscore the potential of ScaleZero for effective large-scale multi-task learning. Our code is available at https://github.com/opendilab/LightZero.
1 Introduction
Unified world models (Reed et al., 2022; Gallouédec et al., 2024b; Lee et al., 2022b) represent a significant step towards generalist agents, offering a single, cohesive framework for multi-modal perception, long-horizon prediction, and decision-making by learning a shared latent space representation. The true power of these models is unlocked when combined with search-based planning algorithms like Monte Carlo Tree Search (MCTS) (Silver et al., 2017; Schrittwieser et al., 2019), which can perform efficient lookahead within this compact latent space. MCTS excels at dynamically balancing exploration and exploitation in complex decision spaces, a combination that has achieved well-established success in homogeneous task domains like board games (Ye et al., 2021; Silver et al., 2017; Schrittwieser et al., 2019). The natural next frontier is to extend this success to heterogeneous multi-task reinforcement learning (MTRL) (Yu et al., 2024), aiming to build a single, generalist agent that can master a wide array of diverse environments. However, this ambition is impeded by a formidable obstacle: training a single shared model on diverse tasks with potentially conflicting dynamics and objectives is notoriously difficult. While prior work (Fernando et al., 2023; Shi et al., 2023) has attempted to mitigate inter-task interference, these methods have been predominantly studied in supervised learning settings. The challenges and dynamics of multi-task planning within a shared world model remain largely unexplored.
Our investigation into this specific domain of multi-task planning identifies and addresses two critical, intertwined obstacles that are not fully resolved by existing approaches: (1) Representational bottlenecks and plasticity collapse: In diverse multi-task settings, a shared model is susceptible to gradient dominance from simpler, faster-converging tasks (Cho et al., 2024). This imbalance leads to representation interference (Yu et al., 2020; Bejnordi et al., 2024), where the learning signals for more complex tasks are suppressed. The consequence is a progressive loss of network plasticity (Dohare et al., 2024; Todorov et al., 2025)the fundamental ability of a model to adjust its parameters to learn from new data–imposing a hard ceiling on the model’s overall learning capacity and leading to performance collapse on challenging tasks. (2) Static resource allocation: Conventional architectures employ a uniform, one-size-fits-all forward pass, applying the same computational effort to every task irrespective of its intrinsic difficulty. This static strategy results in profound computational inefficiency, as resources are squandered on already-mastered tasks instead of being directed toward those that require further learning.
To dissect these issues, we first conduct a rigorous empirical diagnosis in Figure 1. By deploying a baseline UniZero model across eight canonical Atari games, we observe a stark failure pattern. While the model rapidly masters simple tasks like Pong, it exhibits initial progress followed by a catastrophic performance collapse on complex, visually distinct games like Seaquest and ChopperCommand. We provide the quantitative evidence linking this failure to specific internal model dynamics: an uncontrolled inflation of the latent state norm and a corresponding spike in the dormant neuron ratio within the Transformer backbone, signaling a descent into a "frozen," non-adaptive state.
Motivated by this diagnosis, we undertake a principled exploration of the UniZero design space across the axes of input, model, and optimization from the the single learning iteration perspective. This systematic investigation, which evaluates task embeddings, encoder architectures (ResNet (He et al., 2016) vs. ViT (Dosovitskiy et al., 2020)), latent normalization schemes (Hansen et al., 2023), sparse Mixture-of-Experts (MoE) backbones, and gradient rectification methods(He et al., 2020), culminates in a new, powerful model we term ScaleZero. Validated on a comprehensive suite of online RL benchmarks—spanning 26 Atari games, 18 DeepMind Control tasks, and 4 Jericho environments with different observation and action space—ScaleZero consistently matches and often surpasses the performance of specific single-task expert agents.
Finally, to resolve the challenge of computational inefficiency, we introduce Dynamic Parameter Scaling (DPS) from the overall learning process perspective, a novel online mechanism that couples model capacity to learning progress. DPS adaptively curates the set of active tasks based on real-time return feedback and orchestrates a phased expansion of model capacity by injecting lightweight LoRA adapters. By strategically freezing previously trained parameters, DPS creates a curriculum of model complexity that directs computational resources where they are most needed. Our experiments demonstrate that when augmented with our dynamic parameter scaling strategy, our method achieves performance nearly on par with single-task agents while reducing total environment interactions by around 20%, offering a superior trade-off between final performance and computational budget.
Our main contributions are summarized as follows: (1) We provide the quantitative diagnosis of plasticity collapse in unified world models within heterogeneous MTRL, establishing a concrete link between performance degradation and internal network dynamics. (2) We conduct a systematic architectural exploration that yields ScaleZero, a unified world model that demonstrates exceptional performance and generalization across distinct tasks from Atari (diverse visual challenges), DMControl (state-based continuous control), and Jericho (text-based compositional reasoning) benchmark. (3) We propose Dynamic Parameter Scaling (DPS), an adaptive training strategy that dynamically allocates model capacity and computational resources, reducing total environment interactions by around 20%.
2 Background
2.1 Unified World Models in MCTS-Based Reinforcement Learning
Model-based reinforcement learning, particularly methods employing MCTS for planning within a learned latent space, has set the standard in complex sequential decision-making domains (Schrittwieser et al., 2019; Antonoglou et al., 2021; Ye et al., 2021). These methods learn a world model composed of three core components: (i) a representation model, $z_{t}=h_{\theta}(o_{t-H+1:t},a_{t-H+1:t-1})$, which encodes an observation history into a latent state, where $H$ is the training sequence length; (ii) a dynamics model, $(\hat{z}_{t+1},\hat{r}_{t})=g_{\theta}(z_{t},a_{t})$, predicting the subsequent latent state and reward; and (iii) a prediction model, $(p_{t},v_{t})=f_{\theta}(z_{t})$, which estimates a policy and value to guide the MCTS planner. Recent architectures like UniZero (Pu et al., 2024) push this paradigm further by unifying these three components into a single Transformer-based architecture. This monolithic design allows for end-to-end optimization via a composite loss function that aligns the model’s predictions with MCTS-derived targets for policy ($\pi_{t}$) and value ($\hat{v}_{t}$), alongside objectives for reward and state reconstruction:
$\mathcal{L}_{\text{UniZero}}=\sum_{t=0}^{H-1}\left(\mathcal{L}_{\text{value}}(v_{t},\hat{v}_{t})+\mathcal{L}_{\text{policy}}(p_{t},\pi_{t})+\mathcal{L}_{\text{reward}}(\hat{r}_{t},r_{t})+\mathcal{L}_{\text{dynamics}}(\hat{z}_{t+1},\operatorname{sg}(z_{t+1}))\right),$ (1)
where $\operatorname{sg}(\cdot)$ denotes the stop-gradient operator. While parameter-efficient, this unified training scheme creates a critical vulnerability in heterogeneous settings: the shared parameters are subject to a single, aggregated gradient signal, which can be dominated by easier tasks (Ma et al., 2023), thereby impeding learning on more complex ones and leading to plasticity collapse.
2.2 Gradient Conflict in Multi-Task Reinforcement Learning
In MTRL, we consider a distribution of $K$ distinct tasks, $\mathcal{T}=\{\tau_{1},\dots,\tau_{K}\}$. The objective is to learn a single set of shared parameters $\theta$ for a task-conditioned policy $\pi_{\theta}(a|s,k)$ that maximizes expected returns across all tasks, where $k\in\{1,\dots,K\}$ is the task identifier. Given a batch of data from all tasks, the total loss is the sum of the individual task losses: $\mathcal{L}_{\text{MTRL}}(\theta)=\sum_{k=1}^{K}\mathcal{L}^{k}_{\text{UniZero}},$ where $\mathcal{L}^{k}_{\text{UniZero}}$ is the loss computed on data from task $k$. The learning dynamics are driven by the total gradient $G_{\text{total}}=\nabla_{\theta}\mathcal{L}_{\text{MTRL}}$, which is the sum of per-task gradients: $G_{\text{total}}=\sum_{k=1}^{K}g_{k}$, where $g_{k}=\nabla_{\theta}\mathcal{L}^{k}_{\text{UniZero}}$. The core problem, known as gradient conflict, arises when these per-task gradients are misaligned. For any two tasks $i$ and $j$, the conflict can be quantified by the cosine similarity of their gradients: $\cos(g_{i},g_{j})=\frac{g_{i}\cdot g_{j}}{\|g_{i}\|\|g_{j}\|}$. A negative value indicates that an update improving performance on task $i$ will degrade performance on task $j$. In a monolithic architecture where all parameters $\theta$ are shared, frequent negative similarities lead to destructive interference (Sodhani et al., 2021). This forces the model into a suboptimal compromise, creating a representational bottleneck and impeding mastery of tasks with conflicting or complex objectives (Hansen et al., 2023).
2.3 Sparse and Parameter-Efficient Architectures
A significant challenge in developing unified models for heterogeneous environments is overcoming issues such as representational bottlenecks, plasticity collapse, and the rigidity imposed by static resource allocation. To address these challenges, the literature has explored several sparse and parameter-efficient architectures that operate at different perspectives. One prominent approach involves replacing the standard feed-forward network in a Transformer block (Vaswani et al., 2017) with a sparse Mixture-of-Experts architecture (Shazeer et al., 2017; Liu et al., 2024a). This model comprises $M$ expert networks and a trainable sparse gating function, $G(x)$, which selects a small top-k subset of experts for each input. The forward pass is defined as: $\mathrm{MoE}(x)=\sum_{i=1}^{M}G_{i}(x)\cdot\mathrm{Expert}_{i}(x).$ This conditional computation creates distinct representational pathways for different inputs and significantly improve model performance and efficiency with fewer resources. For parameter-efficient adaptation of large models, Low-Rank Adaptation (LoRA) (Hu et al., 2022) has become a widely adopted technique. LoRA adapts a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times m}$ by learning a low-rank update, $\Delta W=BA$, where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times m}$, with the rank $r\ll\min(d,m)$. The effective weight matrix then operates on an input $x$ as: $W(x)=\bigl{(}W_{0}+\alpha\,BA\bigr{)}x,$ where a learnable scalar $\alpha$ controls the magnitude of the update. By freezing the original weights $W_{0}$ and training only the compact low-rank factors $(A,B)$, LoRA facilitates efficient adaptation and isolates task-specific modifications. In this context, our work fuses these ideas from two perspectives. At the single learning iteration level, a gating function sparsely selects LoRA modules, treating them as experts. At the overall learning process level, we incrementally add new LoRA modules for new tasks—freezing prior ones—to efficiently scale model capacity.
3 Method
This section systematically explores the core challenges that unified world models face in heterogeneous multi-task learning. We begin in Section 3.1 by quantitatively diagnosing the phenomenon of plasticity collapse in shared-parameter models under a multi-task paradigm, uncovering its underlying mechanisms. Subsequently, in Section 3.2, we systematically dissect the design space of the UniZero model. Through a series of ablation studies, we converge on a new architecture, termed ScaleZero, which effectively mitigates plasticity degradation. Finally, to address the inefficiency of static resource allocation, we introduce Dynamic Parameter Scaling (DPS) in Section 3.3. This strategy adaptively expands the model’s capacity in response to task-specific learning progress, thereby achieving a superior trade-off between computational efficiency and final performance.
3.1 Diagnosing Plasticity Collapse in Multi-Task Learning
Figure 1: Diagnosing plasticity collapse in the baseline UniZero model on a multi-task Atari benchmark. While simple tasks like Pong and Hero show stable learning, complex tasks such as Seaquest and ChopperCommand suffer a catastrophic performance collapse in later training (Top). This performance failure is precisely correlated with a sharp spike in the dormant neuron ratio of the Transformer backbone (Bottom Left) and an uncontrolled inflation of the latent state norm (Bottom Right), empirically validating the link between external performance and internal network dynamics.
Metrics. In MTRL settings, shared-parameter models are susceptible to a freezing phenomenon, where parts of the network become saturated and lose their ability to adapt to new information. This plasticity collapse stems from the inherent heterogeneity across tasks in terms of their features, difficulty, and informational content, ultimately constraining the model’s overall learning potential. To quantitatively measure this degradation in network adaptability, we utilize two key metrics: (1) Dormant Neuron Ratio (Sokar et al., 2023): This metric measures the proportion of neurons in a given layer whose activation magnitude falls below a predefined threshold $\epsilon$ (set to 0.025 in this work) over a time series. A high dormant ratio indicates that a significant portion of the network’s pathways are inactive and no longer contribute to information processing. $\text{DormantRatio}(l)=\frac{1}{N_{l}}\sum_{i=1}^{N_{l}}\mathbf{1}\!\left(|h_{i}^{l}|\leq\epsilon\right),$ where $h_{i}^{l}$ is the activation of the $i$-th neuron in layer $l$, and $N_{l}$ is the total number of neurons in that layer. (2) Latent State Norm: This metric computes the average L2 norm of the latent state vectors $z_{t}$ produced by the encoder over a time series. An abnormally inflated latent norm is often correlated with training instability and slower learning. As our model processes sequential data in mini-batches, we average this norm across both the training sequence length $T$ and the mini-batch size $B$.
Analysis. Our experiments on a multi-task suite of 8 Atari games (see Appendix A) reveal a clear failure pattern: while simple tasks (e.g., Pong) converge stably, complex tasks (e.g., Seaquest) suffer a catastrophic performance collapse late in training (Figure 1). This collapse precisely corresponds to a sharp increase in the dormant neuron ratio and an uncontrolled inflation of the latent state norm as shown in bottom of Figure 1. In addition, we have added the trend of the feature effective rank Dohare et al. (2024) and training return in Appendix Figure 6, which also shows a marked decline in the performance degradation range. Thus, we attribute this plasticity loss to two intertwined factors: (1) Gradient Competition, where simpler, faster-converging tasks generate dominant gradients that suppress signals from more complex tasks, causing critical neural pathways to "go dormant"; and (2) Representation Interference, where a single, shared model struggles to encode diverse task dynamics, creating a representational bottleneck that constricts the solution space.
Motivated by this diagnosis, we propose to mitigate these issues by introducing task-differentiated mechanisms from two distinct perspectives. First, from the perspective of a single learning iteration, we investigate modifications that resolve interference within each forward and backward pass, leading to the design of ScaleZero (Section 3.2). Second, from the perspective of the overall learning process, we introduce a novel training strategy that dynamically allocates model capacity over time, we named as Dynamic Parameter Scaling strategy (Section 3.3).
3.2 Exploring the Design Space of UniZero
(a) Design Space of UniZero for Multitask learning
(b) Dynamic Parameter Scaling
Figure 2: (a) A systematic exploration of the UniZero design space across five axes: task conditioning, encoder architecture, latent normalization, backbone design, and multitask optimization. This investigation informs the design of our proposed ScaleZero model. (b) A conceptual diagram of Dynamic Parameter Scaling (DPS), our adaptive training strategy. DPS progressively expands model capacity by injecting lightweight LoRA adapters in stages, triggered by learning progress, to direct computational resources toward unsolved tasks while preserving existing knowledge.
Figure 3: Performance impact of architectural modifications on the Atari8 multi-task benchmark. This ablation study across the UniZero design space reveals that replacing the standard Transformer backbone with a Mixture-of-Experts architecture yields the most significant and consistent performance gains. In contrast, other interventions, with the partial exception of SimNorm, provide marginal or inconsistent benefits. These results underscore the centrality of the MoE’s conditional computation in overcoming the limitations of a shared, dense backbone.
Figure 4: Detailed analysis correlating performance with plasticity metrics across model variants in Figure 3. On two representative hard tasks (ChopperCommand and Seaquest), the performance collapse of baseline models directly correlates with catastrophic plasticity loss—evidenced by a sharp rise in the dormant neuron ratio and latent norm. In stark contrast, the MoE-based model sustains high returns by keeping the dormant ratio near-zero and the latent norm stable. This confirms that the MoE’s superior performance stems directly from its success in mitigating the diagnosed plasticity issues. Analyses for the remaining six environments, which show consistent trends, are deferred in Figure 8.
Based on the diagnosis above, we conducted a systematic decomposition and evaluation of the UniZero model’s design space to identify architectural components that can effectively counteract plasticity loss. We explored five key dimensions, as illustrated in Figure 2 (left): task information, encoder, latent normalization, backbone, and optimization strategy. The experimental details is showed in Appendix B. Figure 3 provides a overall comparison of the performance contributions from each design dimension. A key observation is that the MoE backbone yields the most significant performance gains. In contrast, other modifications, with the partial exception of SimNorm, fail to deliver stable performance improvements. To investigate the mechanisms behind these results, Figure 4 illustrates the relationship between task returns and plasticity-related metrics (i.e., transformer dormant ratio and latent norm) for two representative tasks, ChopperCommand and Seaquest. Analyses for the remaining six environments are provided in Appendix B. The key findings from this exploration are detailed below:
Explicit Task Conditioning. We experimented with concatenating a learnable task embedding indexed by task ID to the state representation to provide an explicit task identity. While this accelerated initial convergence on simpler tasks, its long-term benefit for complex tasks was marginal, and it failed to significantly improve internal plasticity metrics (Figure 4). This suggests that shallow task identifiers are insufficient to resolve deep-seated representational conflicts in online RL setting.
Encoder Architecture (ResNet vs. ViT). We compared a standard ResNet encoder against a ViT to assess the impact of encoder architecture. In our experiments on the Atari8 benchmark, the choice of encoder did not yield a significant performance advantage. We hypothesize that this is due to the limited number of tasks and the relatively low visual heterogeneity across the environments in this specific benchmark. However, we anticipate that the ViT’s superior capacity and scalability will become more critical in larger-scale, more heterogeneous multi-task settings.
Latent Normalization (LayerNorm vs. SimNorm). To suppress the inflation of the latent norm, we compared a standard LayerNorm (Ba et al., 2016) against a SimNorm (Hansen et al., 2023) to assess the impact of latent normalization. SimNorm (Hansen et al., 2023) normalize the latent representation by projecting $z$ into $L$ fixed-dimensional simplices using a softmax operation, which clamps the L2 norm of latent vectors within a predefined threshold. Although SimNorm effectively stabilized training and prevented performance collapse, its hard constraint appeared to curtail the expressive power of the representations, highlighting a trade-off between stability and representational capacity.
Mixture-of-Experts Backbone. To fundamentally address plasticity loss in the Transformer backbone, we replaced its standard feed-forward networks (FFNs) with MoE layers. In an MoE, a gating network sparsely routes each input token to a subset of "expert" sub-networks. This conditional computation mechanism decouples the dense computational pathways into multiple, parallel, specialized ones. This design: (1) reduces parameter collision, as different tasks can leverage different combinations of experts, mitigating representation interference; and (2) mitigates gradient conflict, as gradient updates for different tasks are largely directed to the distinct sets of activated experts. As shown in Figures 4 and 4, the MoE backbone delivered the most significant performance improvements on complex tasks like Seaquest and maintained a relatively low dormant ratio throughout training, effectively controlling the latent norm.
Multi-Task Gradient Correction. We implemented a dynamic gradient re-weighting scheme inspired by MoCo (Fernando et al., 2023) to directly address gradient interference. Despite yielding gains on some tasks, its significant computational overhead and inconsistent efficacy across the heterogeneous task suite led us to explore a more efficient gradient correction in future.
ScaleZero. Based on this exploration, we integrated the most impactful components to form the ScaleZero architecture. Its core design combines: (1) a ViT Encoder for powerful feature extraction; (2) an MoE Backbone to fundamentally address plasticity collapse through sparse, conditional computation; and (3) Standard Layer Normalization on the latent states to ensure stability without sacrificing expressive power. As validated in Section 4, ScaleZero provides a novel and effective blueprint for building generalist world models.
3.3 Dynamic Parameter Scaling for Balanced Multi-Task Learning
In heterogeneous MTRL, static, one-size-fits-all model architectures are suboptimal. They tend to waste computational resources on tasks that have already been mastered and often lack the necessary plasticity to effectively tackle the remaining, more difficult tasks. This can lead to inefficient training and risks negative transfer. To address this, we propose Dynamic Parameter Scaling (DPS), a strategy illustrated in Figure 2 (right). The core insight of DPS is that a model’s effective capacity should dynamically align with the current set of unsolved tasks. This approach instantiates a "curriculum of model complexity," adaptively expanding the model with new parametric resources directed precisely at the most challenging tasks still in progress.
Formalism and Mechanism. Let the set of tasks be $\mathcal{T}=\{\tau_{i}\}_{i=1}^{N}$ and the parameters of the shared backbone network be $\theta_{B}$. We define a task $\tau_{i}$ as "solved" when its performance metric, $\text{Metric}(\tau_{i})$, meets or exceeds a predefined threshold $\varepsilon_{i}$. At any time $t$, the set of all unsolved tasks is denoted as the active set $\mathcal{U}_{t}\subseteq\mathcal{T}$. To conserve computational resources, both data collection and model training are exclusively performed on the tasks within this active set. Once a task is considered solved, it is removed from $\mathcal{U}_{t}$, and we cease all related training activities for it. The training protocol is divided into $S+1$ stages ($s=0,1,\dots,S$). It begins with Stage 0 (Warm-up), an initial period of $T_{0}$ iterations where only the backbone parameters $\theta_{B}$ are trained. Each subsequent Stage $s\geq 1$ (Expansion) introduces an independent Low-Rank Adaptation (LoRA) module, $\Delta\theta_{s}$. This module consists of a trainable matrix pair $\Delta\theta_{s}=B_{s}A_{s}$ (where $B_{s}\in\mathbb{R}^{d\times r},A_{s}\in\mathbb{R}^{r\times k}$ with rank $r\ll d,k$) and a learnable scalar scaling factor $\alpha_{s}\in\mathbb{R}$.
Adaptive Stage Transition Triggers. The transition from stage s1s-1 to stage $s$ is governed by one of two conditions, balancing learning progress against a fixed training budget: (1) Progress-based Trigger: A new stage is initiated when the number of newly solved tasks since the beginning of the current stage reaches a predefined quota $Q_{s}$. (2) Budget-based Trigger: To prevent training from stalling on exceptionally difficult tasks, a budget-based trigger forces a transition if the number of iterations in the current stage exceeds a pre-allocated limit, $\lceil(T_{\max}-T_{0})/S\rceil$.
Parameter Composition and Optimization. In any given stage $s$, the effective computation for a weight matrix $W_{0}\in\theta_{B}$ is expressed as $W^{(s)}(x)=W_{0}x+\sum_{j=1}^{s}\alpha_{j}(B_{j}A_{j})x$. Crucially, upon entering a new stage $s$, we freeze both the backbone parameters $\theta_{B}$ and all previously introduced LoRA modules $\{\Delta\theta_{j}\}_{j=1}^{s-1}$. The optimization gradient is then applied only to the newly added module $\Delta\theta_{s}$ and the full set of scaling factors $\{\alpha_{j}\}_{j=1}^{s}$. This strategy offers a compelling dual advantage. First, it is computationally efficient by adaptively allocating resources according to task difficulty; new parameters are only introduced as needed for unsolved tasks, avoiding redundant computation on those already mastered. Second, the dynamic injection of new LoRA modules provides targeted plasticity for more challenging tasks without introducing negative transfer. By isolating updates within new modules while preserving the stable knowledge in the frozen backbone, our approach ensures that learning to solve difficult tasks does not disrupt previously acquired capabilities. In summary, DPS enables an intelligent allocation of model capacity, achieving a superior trade-off between efficiency and final performance. Pseudocode and implementation details are provided in Appendix A.3.
4 Experiments and Analysis
To comprehensively evaluate the effectiveness of our proposed ScaleZero model and the dynamic parameter scaling strategy, we designed a series of experiments. Section 4.1 aims to validate the performance of ScaleZero on several challenging heterogeneous multitask benchmarks (Atari 100k benchmark (Bellemare et al., 2013a), DeepMind Control Suite (Tunyasuvunakool et al., 2020), and Jericho(Hausknecht et al., 2020)), comparing it against specialized models trained independently for each task. Section 4.2 then focuses on quantitatively analyzing the specific gains in sample and computational efficiency afforded by our dynamic parameter scaling strategy (the combined method we called ScaleZero-DPS). All experiments are conducted in a purely online RL setting, without relying on any expert data, to test the model’s capabilities in a realistic learning scenario. Section 4.3 complements our benchmark results with a focused analysis of MoE, combining empirical evidence with theoretical justification to explain why MoE serves as an effective backbone for multitask learning. Implementation details are provided in Appendix A.
4.1 Multitask Learning Benchmarks
4.1.1 Performance Evaluation on Atari 100k benchmark
Setup. We evaluate ScaleZero on the large-scale Atari 100k benchmark, which comprises 26 distinct games. This benchmark is renowned for its significant heterogeneity across tasks, featuring diverse observation spaces (pixel inputs), action spaces (discrete instruction sets), and widely varying game difficulties and mechanics. We compare the multitask (MT) ScaleZero model against a single-task (ST) UniZero baseline, which is individually trained and optimized for each game. The evaluation metrics are the mean and median of the Human-Normalized Score (HNS), a standard in the field, to provide a comprehensive assessment of performance across tasks of varying difficulty.
Results. As shown in Table 1, by training a single, unified model, ScaleZero achieves a mean normalized score that surpasses the average performance of 26 independently trained single-task UniZero models. This indicates that ScaleZero not only successfully learns a diverse set of tasks but also achieves positive knowledge transfer across them. Notably, the performance gains are most pronounced on several notoriously challenging tasks, such as Seaquest. We hypothesize that this is because ScaleZero’s unified world model can learn general visual and dynamic priors from simpler tasks and effectively apply them to more difficult ones that demand complex planning and exploration strategies. Although the median score is slightly lower than the single-task baseline, which is primarily influenced by a few tasks with extremely hard exploration challenges, the superior mean score robustly demonstrates the overall advantage of our approach.
Table 1: Comparison of Human-Normalized Scores on 26 Atari games for ScaleZero (MT) vs. UniZero (ST). Detailed curves are available in Appendix B Figure 9.
Algorithm Normalized Mean Normalized Median
UniZero (ST) 0.38 0.21
ScaleZero (MT) 0.39 0.16
4.1.2 Performance Evaluation on the DeepMind Control Suite
Setup. To test the generalization capability of ScaleZero in the continuous control domain, we conducted evaluations on the DMC benchmark, which includes 18 tasks based on the MuJoCo physics engine. These tasks feature vector state space and continuous action spaces, posing higher demands on the model’s fine-grained control and dynamics modeling abilities. ScaleZero also leverages the principles of Sampled Policy Iteration (Hubert et al., 2021) same as UniZero. Note that because observation space of DMC is vector, we use the MLP encoder. The experimental setup mirrors that of the Atari experiments, comparing the multitask ScaleZero against 18 independent single-task UniZero models.
Results. The results in Table 2 show that ScaleZero achieves returns comparable to or even higher than the single-task baselines on the vast majority of DMC tasks. The learning curves, as illustrated in Figure 10, clearly reveal that ScaleZero can effectively handle the diverse physical dynamics and control requirements of different tasks within a single, unified model. This success validates the versatility of the ScaleZero architecture, proving its competitiveness in complex control problems involving high-dimensional continuous states and actions.
Task UniZero (ST) ScaleZero (MT)
acrobot-swingup 400.3 501.0
cartpole-balance 952.2 990.5
cartpole-balance_sparse 1000.0 1000.0
cartpole-swingup 801.3 769.0
cartpole-swingup_sparse 752.5 708.1
cheetah-run 517.6 510.9
ball_in_cup-catch 961.6 954.2
finger-spin 810.7 574.2
finger-turn_easy 1000.0 1000.0
finger-turn_hard 884.5 982.0
hopper-hop 120.5 138.0
hopper-stand 602.6 583.1
pendulum-swingup 865.6 866.0
reacher-easy 993.3 943.1
reacher-hard 988.8 943.5
walker-run 587.9 562.7
walker-stand 976.4 919.9
walker-walk 954.6 908.7
Mean 787.2 769.7
Median 875.1 887.3
Table 2: Performance comparison between ScaleZero (MT) and UniZero (ST) across various tasks in DMControl. For reference, L2M (MT) Schmied et al. (2023) obtained a mean score of 840 under an offline supervised learning from expert data setting. The best performance for each task is indicated in bold. ScaleZero’s higher human-normalized median scores highlight its strong performance across continuous action spaces. Detailed curves are available in Appendix C  Figure 10.
4.1.3 Performance Evaluation on the Jericho Text-Based Adventure Games
Setup. To further evaluate ScaleZero’s text-based reasoning and planning, we assess it on the Jericho (Hausknecht et al., 2020) benchmark of interactive fiction, where both observations and actions are expressed in natural language. Jericho features a combinatorial action space and sparse, delayed rewards, and it demands strong long-horizon contextual reasoning. We use four representative games—Acorncourt, Omniquest, Detective, and Zork1—covering diverse narrative structures and interaction complexity. For parity, both the single-task UniZero and the multitask ScaleZero employ the same BGE text encoder (Xiao et al., 2023). In addition, we compare against CALM+OC, a recent LM-in-the-Loop method that adaptively updates the action generator using state-feature–based categorization of gameplay transitions (Sudhakar et al., 2023). The primary evaluation metric is the average return measured after training convergence.
Results. The results in Table 3 show that ScaleZero achieves returns comparable to the single-task UniZero baseline and generally outperforms the CAML+OC method across most games. The learning curves in Figure 13 further indicate that a single, unified model can efficiently handle Jericho’s combinatorial language action space and sparse, delayed rewards. These findings validate the applicability of the ScaleZero architecture for text-based RL, demonstrating competitive performance on tasks that require language understanding and long-horizon planning.
Table 3: Average returns of ScaleZero and other baselines on four Jericho tasks. CALM+OC(Sudhakar et al., 2023) denotes the variant of CALM with online categorization based on state features. Detailed learning curves for ScaleZero and UniZero are provided in Appendix D  Figure 13.
Algorithm Acorncourt Omniquest Detective Zork1
CALM+OC (ST) 7.8 288.5 38.0
UniZero (ST) 10 10 295 44
ScaleZero (MT) 10 10 280 44
4.2 Efficiency Evaluation of Dynamic Parameter Scaling
This section aims to evaluate the effectiveness of our LoRA-based online dynamic parameter scaling strategy integrated into ScaleZero, termed ScaleZero-DPS, in improving sample and computational efficiency. To this end, we conducted a comparative analysis between the standard ScaleZero and ScaleZero-DPS on the full DMC18 benchmark. The core evaluation metric is defined as the number of environment interaction steps required to reach a performance level comparable to the single-task baseline (i.e., 100% of its performance). As illustrated in Figure 5, ScaleZero-DPS achieves the target performance using only about 80% of the interaction steps required by the standard ScaleZero. This result highlights the significant efficiency gains afforded by our dynamic scaling strategy. Furthermore, the detailed performance curves, presented in Appendix Figure 11, reveal that ScaleZero-DPS exhibits a steeper learning curve. This indicates a faster learning rate, leading to higher sample and computational efficiency throughout the training process.
Figure 5: Comparison of interaction cost for ScaleZero vs. ScaleZero-DPS on DMC18. ScaleZero-DPS reaches the target performance with approximately 80% of the interaction steps. For detailed performance curves, please refer to Appendix C Figure 11.
4.3 Understanding MoE Efficacy: An Empirical and Theoretical Analysis
The demonstrated success of Mixture-of-Experts (MoE) in multi-task scenarios motivates a deeper investigation into its underlying mechanisms. In the following, we explore these mechanisms from both experimental and theoretical perspectives. We begin by summarizing our empirical findings, which indicate that replacing standard MLP layers with MoE counterparts significantly reduces gradient conflicts. Notably, this reduction is not confined to the MoE layers themselves but also extends to adjacent network components, such as their input (detailed experimental settings and results are provided in Section˜E.1).
To complement our empirical analysis, we provide a theoretical justification for how MoE mitigates gradient conflict. Our theory builds upon prior work (Chen et al., ) that characterizes MoE training by two distinct phases: an initial exploration stage, where inputs are routed nearly uniformly across experts, followed by a router learning stage, where the router gradually specializes in directing inputs. Grounded in this two-stage framework, we propose an informal theorem asserting that, regardless of the training stage, the upper bound on gradient conflict for an MoE layer is strictly lower than that of its MLP counterpart, and that this reduction is primarily governed by the router’s behavior.
Theorem 4.1 (Upper Bound on Gradient Conflict in MoE Layers (informal)).
In a Mixture-of-Experts (MoE) layer, consider two tasks $t_{1}$ and $t_{2}$ with routing weights $\lambda^{t_{1}}_{m},\lambda^{t_{2}}_{m}$ over $M$ experts, and per-task gradients on expert $m$, denoted as $g_{t_{1}}^{(m)}$ and $g_{t_{2}}^{(m)}$. Let $G$ be the maximum gradient conflict on any single expert, and $\cos(u,v)$ denote the cosine similarity between vectors $u$ and $v$. Then the following results hold:
(1). (Generally) Define the $m$-th element of vectors $u,v$ as $u_{m}=\lambda^{t_{1}}_{m}\|g_{t_{1}}^{(m)}\|,\;v_{m}=\lambda^{t_{2}}_{m}\|g_{t_{2}}^{(m)}\|,\;m=1,\dots,M.$ The full-layer gradient conflict admits a general upper bound: $\mathrm{conflict}(G_{t_{1}},G_{t_{2}})\leq G\cdot\cos(u,v)$.
(2). (Sparse Router) In exploration stage, where tasks uniformly select experts, the expected gradient conflict of the MoE layer is approximately bounded by $G\cdot(1-M!/((M-K)!\,M^{K})$, where $K$ is the task number.
(3). (Sparse Router) In the router learning stage, where the expert sets selected by the router gradually stabilize, the gradient conflict of the MoE layer is roughly upper bounded by $\mathbb{E}[\text{conflict}_{i,j}]\approx G\cdot\frac{U}{ab}$, where $a=|S_{i}|$ and $b=|S_{j}|$ denote the sizes of the expert sets selected by tasks $i$ and $j$, and $U=|S_{i}\cap S_{j}|$ denotes the size of their intersection.
In Section˜E.2, we provide detailed explanations of the the three parts of this theorem and their related corollaries, along with complete proofs.
5 Related Work
MCTS-based Planning with Learned World Models. The integration of MCTS with deep neural networks has become a cornerstone of modern reinforcement learning, particularly in domains requiring sophisticated planning. The seminal AlphaZero series (Silver et al., 2016; 2017) established a powerful paradigm by coupling MCTS with learned policy and value networks. This approach was further advanced by MuZero (Schrittwieser et al., 2019), which obviated the need for an explicit environment simulator by learning a latent world model to jointly predict future rewards, policies, and values. This architecture enables planning entirely within a learned representation space. Subsequent research has extended the MuZero framework, focusing on enhancing model capabilities (Hubert et al., 2021; Antonoglou et al., 2021), improving sample efficiency (Danihelka et al., 2022; Ye et al., 2021), and refining training stability (Schrittwieser et al., 2021). Notably, the advent of Transformer-based world models (Micheli et al., 2022; Robine et al., 2023), as exemplified by UniZero (Pu et al., 2024), has significantly augmented representational capacity, enabling stronger long-horizon planning and generalization. However, the monolithic, dense representations that enable single-task excellence become a liability in multi-task settings, causing representational interference and destructive interference. Simply scaling these models exacerbates this architectural bottleneck. Our work try to confront this fundamental limitation.
Multi-Task Reinforcement Learning. To address the challenge of generalization across diverse tasks, MTRL aims to leverage shared knowledge to improve data efficiency and performance(Vithayathil Varghese & Mahmoud, 2020; D’Eramo et al., 2024). A prevalent architectural motif is the shared-backbone with task-specific heads (Kumar et al., 2022; Hansen et al., 2023; 2022; Lee et al., 2022a; Gallouédec et al., 2024a). While promoting knowledge transfer, this design is insufficient for world models where the core challenge lies in disentangling the latent dynamics predictions within the backbone itself, not just the final policy or value outputs. Recent research in MTRL has pursued several avenues to mitigate these issues. From a representational standpoint, context-based methods(Rakelly et al., 2019; Sodhani et al., 2021; Seo et al., 2022) condition the model on task metadata to enable selective knowledge sharing, while others introduce learnable modulation modules to preserve task-specific knowledge (Schmied et al., 2023). Architecturally, modular designs have emerged as a promising direction(Lan et al., 2023; Yang et al., 2020; Sun et al., 2022). These include soft modularization with task-aware routing (Yang et al., 2020; He et al., 2023; Hendawy et al., 2023) and constructing policy subspaces via parameter composition (Sun et al., 2022; 2023). On the optimization front, various strategies(Fernando et al., 2023; Lin et al., 2024; Ma et al., 2023) have been developed to manage conflicting gradients, a critical impediment in MTRL. These include projecting conflicting gradients into a non-conflicting space (Fernando et al., 2023) and harmonizing training through dynamic loss and gradient normalization (Lin et al., 2024; Ma et al., 2023; Chen et al., 2018). While effective at mitigating conflicts, these methods primarily address the symptoms (conflicting gradients) rather than the root cause: the architectural bottleneck of a model lacking sufficient specialized capacity to represent diverse task dynamics. These diverse approaches underscore the central challenge in MTRL: balancing knowledge sharing with task specialization. Our work seeks to achieve this balance through a novel architectural design.
Sparse and Parameter-Efficient Architectures. Scaling deep RL models is computationally expensive (Espeholt et al., 2018). Sparse activation models, like the Mixture-of-Experts (MoE) (Dai et al., 2024), address this by increasing representational power at a near-constant computational cost via conditional computation. Their success in single-task RL (Obando-Ceron et al., 2024) highlights their potential as a native architectural prior for multi-task specialization. Concurrently, parameter-efficient fine-tuning (PEFT) methods (Dou et al., 2024; Dettmers et al., 2023; Hayou et al., 2024; Liu et al., 2024c) like Low-Rank Adaptation (LoRA) (Hu et al., 2022) enable lightweight, task-specific adaptations by tuning a small subset of parameters, avoiding the cost of full fine-tuning in MTRL (Wang et al., 2023; Yang et al., 2025; Agiza et al., 2024; Zhang et al., 2025). We identify a critical synergy between these two approaches. MoE enables coarse-grained, input-dependent specialization by routing computations to expert sub-networks, tackling representational entanglement at an architectural level. In parallel, LoRA facilitates fine-grained, task-specific adaptation in a parameter-efficient manner, injecting nuanced adjustments without catastrophic interference. The synergistic integration of MoE’s scalable representation with LoRA’s efficient adaptation is a critical frontier, yet one that has seen little exploration(Dou et al., 2023). Our work bridges this gap by proposing a novel architecture that unifies MoE and LoRA within a Transformer-based world model.
6 Conclusion and Future work
In this paper, we examine the diminishing plasticity phenomenon in unified world models under multi-task scenarios and quantify its impact on model performance. By exploring the expansive design space defined by the input–model–optimization axes, we introduce ScaleZero—a principled method that carefully balances performance with sample efficiency. Our extensive experiments on pure online RL benchmarks—which include 26 Atari, 18 DMControl, and 4 Jericho environments—demonstrate that ScaleZero achieves average returns comparable to or even surpassing those of specialized single-task baselines. To further balance computational costs, we introduced a LoRA-based dynamic parameter scaling (DPS) strategy. This strategy has currently only been prototyped on the DMC benchmark, where it achieved competitive performance with 80% of the interaction steps, showcasing its potential for improving sample efficiency. Despite these promising results, several avenues remain open for future investigation: A comprehensive evaluation of the DPS strategy’s generalization, scalability, and overhead is required on broader task sets like Atari. Investigating deeper synergistic mechanisms between MoE and LoRA to further enhance model expressivity and adaptability. Combining our online framework with offline pre-training of large-scale models to potentially boost both efficiency and final performance. We believe the continued exploration of these directions will advance multi-task reinforcement learning toward the goal of general intelligence.
Appendix A Implementation Details
Appendix
This appendix details the implementation of our ScaleZero model, explorations of key design spaces, the Dynamic Parameter Scaling strategy, hyperparameter configurations, and the computational resources used for our experiments. Unless otherwise specified, the training and inference details remain consistent with those in the paper Pu et al. (2024).
A.1 Core Modifications of ScaleZero over UniZero
Our core model, ScaleZero, is built upon the UniZero baseline, with key components upgraded to enhance learning efficiency and final performance in complex multi-task environments.
Visual Encoder: For image-based inputs (e.g., Atari), we employ a Vision Transformer (ViT) as the primary encoder. The input image (64x64x3) is divided into patches (8x8), which are then converted into token embeddings via a linear projection layer. Learnable positional embeddings are added, and the sequence is fed into a standard Transformer encoder (ViT-Base architecture) to produce a 768-dimensional latent state.
Text Encoder: In text-based environments like Jericho, we use a pre-trained BGE (BAAI General Embedding) model (Xiao et al., 2023) to encode the textual observations from the game into fixed-dimensional vectors.
Unified World Model:
  • Backbone: The core of the model is a Transformer network. It processes state tokens from the encoder, historical action tokens, and optional task-embedding tokens.
  • UniZero (Baseline): The backbone consists of $N$ standard Transformer blocks. We use $N=8$ for the Atari26 benchmark and $N=4$ for other benchmarks.
  • ScaleZero (Our Model): To maintain a comparable parameter count, the backbone is composed of $N/2$ Mixture-of-Experts (MoE) Transformer blocks. In each of these blocks, the standard feed-forward network (FFN) layer is replaced by an MoE layer. This layer comprises 8 expert networks and one shared expert, governed by a gating network that performs sparse routing.
Latent State Normalization: In ScaleZero, we apply standard LayerNorm to normalize the latent state.
A.2 Explorations on Key Design Spaces
We conducted several exploratory extensions to key designs of UniZero to validate their effectiveness in multi-task settings.
Task Information Encoding: To mitigate representational conflicts in multi-task learning, we explored explicit encoding of task information. By introducing a learnable task embedding matrix, task_embed = nn.Embedding(task_id), and concatenating it with the state representation, new_latent_state = concat(latent_state, task_embed), we aimed for the model to better distinguish between different tasks. In our experiments, the latent_state dimension was 672, and the task_embed dimension was 96.
Vision Transformer (ViT): Our ViT implementation is adapted from the lucidrains/vit-pytorch [^1] https://github.com/lucidrains/vit-pytorch library. The core hyperparameters, corresponding to a ViT-Base level, are summarized in Table 4.
Table 4: Hyperparameters for the Vision Transformer (ViT) encoder.
Hyperparameter Value
image_size 64x64
patch_size 8x8
dim (Embedding Dimension) 768
depth (Transformer Layers) 6
heads (Attention Heads) 6
mlp_dim (MLP Hidden Dimension) 2048
dropout 0.1
emb_dropout (Embedding Dropout) 0.1
Latent Normalization Strategy (SimNorm): As part of our ablation studies, we explored the SimNorm strategy, inspired by the TD-MPC2 paper. SimNorm partitions the latent state $z$ into $L$ groups, where each group is a $V$-dimensional simplex $g$. The transformation is given by:
$z_{\text{sim\_norm}}=[g_{1},\dots,g_{i},\dots,g_{L}],\quad g_{i}=\frac{\exp(z_{i:i+V}/\tau)}{\sum_{j=1}^{V}\exp(z_{i:i+V_{j}}/\tau)}$ (2)
In our experiments, we set the simplex dimension $V=8$.
Mixture-of-Experts (MoE): Our MoE layer implementation is inspired by the design in the mistral-inference [^2] https://github.com/mistralai/mistral-inference/blob/main/src/mistral_inference/moe.py library. Its mechanism operates as follows: for each token in the input sequence, a gating network computes affinity scores to select the Top-1 expert. The selected tokens are then processed by their corresponding expert networks (standard feed-forward networks). The output of each expert is multiplied by its corresponding gating weight (after Softmax normalization) and then summed. Crucially, to ensure knowledge sharing and generalization, all tokens are also processed by a shared expert network, whose output is added to the final result of the MoE layer.
Multi-task Gradient Correction (MoCo): We adopted the MoCo algorithm from the LibMTL library [^3] https://github.com/median-research-group/LibMTL to handle gradient conflicts. Its core idea is to first maintain a momentum term $y_{t}$ to smoothly track each task’s gradient direction. Then, task weights $\lambda$ are dynamically updated by solving a regularized optimization problem. Finally, the gradient applied to shared parameters is the weighted sum of the momentum gradients. Key hyperparameters are listed in Table 5.
Table 5: Hyperparameters for the MoCo algorithm.
Hyperparameter Value
MoCo_beta ($\beta$) 0.99
MoCo_gamma ($\gamma$) 10.0
MoCo_rho 0.0
A.3 Dynamic Parameter Scaling (DPS)
To address the issue of non-uniform learning progress across tasks in multi-task learning, we propose the Dynamic Parameter Scaling (DPS) strategy. The core idea is to dynamically align the model’s effective capacity with the set of currently unsolved tasks. This is achieved through a curriculum-based expansion of the model’s parameters. The detailed pseudocode is provided in Algorithm 1.
Staged Training and Active Set Management. The training protocol is divided into $S+1$ stages. We define a task $\tau_{i}$ as "solved" once its performance metric surpasses a predefined threshold $\varepsilon_{i}$. The set of all currently unsolved tasks forms the active set $\mathcal{U}_{t}$. To maximize efficiency, both data collection and training updates are performed exclusively on tasks within this active set.
  • Stage 0 (Warm-up): During an initial period of $T_{0}$ iterations, only the shared backbone network parameters $\theta_{B}$ are trained. This allows the model to learn a foundational, general-purpose representation across all tasks.
  • Stage $s\geq 1$ (Expansion): Each subsequent stage introduces new, dedicated parameters in the form of an independent Low-Rank Adaptation (LoRA) module, $\Delta\theta_{s}=B_{s}A_{s}$, and a corresponding learnable scalar scaling factor $\alpha_{s}$. These new parameters are intended to provide additional capacity for tackling the remaining difficult tasks in $\mathcal{U}_{t}$.
Adaptive Stage Transition Triggers. The transition from stage s1s-1 to stage $s$ is governed by a dual-trigger mechanism, ensuring a balance between progress and resource allocation:
  1. 1.
    Progress-Driven Trigger: A new stage is initiated if the number of newly solved tasks since the beginning of the current stage reaches a predefined quota $Q_{s}$. This allows the model to expand only when it has demonstrated mastery over a subset of the current challenges.
  2. 2.
    Budget-Driven Trigger: To prevent stagnation on particularly hard tasks, a transition is forced if the number of iterations within the current stage exceeds a maximum budget, calculated as $\lceil(T_{\max}-T_{0})/S\rceil$. This ensures that the model continues to gain new capacity even if progress stalls.
Hierarchical Parameter Composition and Optimization. In any given stage $s$, the forward pass for a weight matrix $W_{0}\in\theta_{B}$ is dynamically composed as:
$W^{(s)}(x)=W_{0}x+\sum_{j=1}^{s}\alpha_{j}(B_{j}A_{j})x$ (3)
Upon entering a new stage $s$, we implement a crucial optimization strategy: we freeze the backbone parameters $\theta_{B}$ and all previously introduced LoRA modules $\{\Delta\theta_{j}\}_{j=1}^{s-1}$. The optimization gradient is then applied only to the newly introduced LoRA module $\Delta\theta_{s}=\{A_{s},B_{s}\}$ and to all existing scaling factors $\{\alpha_{j}\}_{j=1}^{s}$. This design provides two key benefits: it mitigates negative transfer by isolating new parameter updates within the fresh LoRA module, and it retains plasticity by allowing the model to dynamically re-weigh the contribution of knowledge from all previous stages via the trainable scalars $\{\alpha_{j}\}$.
The specific DPS settings used in our ScaleZero-Balanced experiments are detailed in Table 6.
Table 6: Hyperparameters for Dynamic Parameter Scaling (DPS).
Hyperparameter Value Description
curriculum_stage_num 5 Total number of expansion stages ($S$).
lora_r 64 The rank of the LoRA matrices ($r$).
lora_alpha 1 Initial scaling factor for LoRA layers.
min_stage0_iters 10,000 Duration of the warm-up stage ($T_{0}$).
max_stage_iters 5,000 Per-stage iteration budget for the budget-driven trigger.
Algorithm 1 Dynamic Parameter Scaling (DPS)
1:Set of tasks $\mathcal{T}=\{\tau_{i}\}_{i=1}^{N}$; Performance thresholds $\{\varepsilon_{i}\}_{i=1}^{N}$; Backbone params $\theta_{B}$;
2:Hyperparameters: Total stages $S$, Warm-up iterations $T_{0}$, Max total iterations $T_{\max}$, Stage quotas $\{Q_{s}\}_{s=1}^{S}$.
3:Trained parameters: $\theta_{B}$, $\{\Delta\theta_{s},\alpha_{s}\}_{s=1}^{S}$.
4:Initialize active set of unsolved tasks: $\mathcal{U}\leftarrow\mathcal{T}$.
5:for stage $s=0$ to $S$ do
6:  if $s==0$ then $\triangleright$ Stage 0: Warm-up Phase
7:  Set trainable parameters $\theta_{\text{train}}\leftarrow\theta_{B}$.
8:  Set stage iteration limit $L_{s}\leftarrow T_{0}$.
9:  for $t_{\text{stage}}=1$ to $L_{s}$ do
10:   Sample a batch exclusively from tasks in active set $\mathcal{U}$.
11:   Perform forward pass with backbone: $W(x)=W_{0}x$.
12:   Compute loss $\mathcal{L}$ and update $\theta_{\text{train}}$.
13:   Update active set: $\mathcal{U}\leftarrow\{\tau_{i}\in\mathcal{T}\mid\text{Metric}(\tau_{i})<\varepsilon_{i}\}$.
14:  end for
[…]
16:  Initialize new LoRA module $\Delta\theta_{s}=B_{s}A_{s}$ and scalar $\alpha_{s}$.
17:  Freeze $\theta_{B}$ and all previous modules $\{\Delta\theta_{j}\}_{j=1}^{s-1}$.
18:  Set trainable parameters $\theta_{\text{train}}\leftarrow\{\Delta\theta_{s}\}\cup\{\alpha_{j}\}_{j=1}^{s}$.
19:  Record solved tasks at stage start: $\mathcal{S}_{\text{start}}\leftarrow\mathcal{T}\setminus\mathcal{U}$.
20:  Set budget for current stage: $T_{\text{budget}}\leftarrow\lceil(T_{\max}-T_{0})/S\rceil$.
21:  $t_{\text{stage}}\leftarrow 0$.
22:  while $t_{\text{stage}}<T_{\text{budget}}$ do $\triangleright$ Budget-based Trigger
23:   $t_{\text{stage}}\leftarrow t_{\text{stage}}+1$.
24:   Sample a batch exclusively from tasks in active set $\mathcal{U}$.
25:   Perform forward pass: $W^{(s)}(x)=W_{0}x+\sum_{j=1}^{s}\alpha_{j}(B_{j}A_{j})x$.
26:   Compute loss $\mathcal{L}$ and update $\theta_{\text{train}}$.
27:   Update active set: $\mathcal{U}\leftarrow\{\tau_{i}\in\mathcal{T}\mid\text{Metric}(\tau_{i})<\varepsilon_{i}\}$.
28:   if $|\mathcal{T}\setminus\mathcal{U}|-|\mathcal{S}_{\text{start}}|\geq Q_{s}$ then $\triangleright$ Progress-based Trigger
29:    break $\triangleright$ Quota of newly solved tasks met, transition to next stage.
30:   end if
31:  end while
32:  end if
33:end for
A.4 Common Settings and Hyperparameters
To ensure a fair comparison, most hyperparameters are kept identical to the original UniZero paper. We use the AdamW optimizer. Key shared hyperparameters are summarized in Table 7.
Table 7: Common hyperparameters for all experiments.
Hyperparameter Value
General Training
Learning Rate 1e-4
Weight Decay 0.0001
Batch Size 256
Discount Factor ($\gamma$) 0.997
Training Context Length ($H$) 10
World Model
Num. Transformer Layers 4 (Atari-8, DMC, Jericho), 8 (Atari-26)
Embedding Dimension 768
Unroll Steps 5
Num. Experts (per MoE layer) 8 + 1 (shared)
MCTS
Num. Simulations 50
Inference Context Length ($H_{\text{infer}}$) 4
Replay Buffer
Replay Buffer Size 100,000
Loss Function Weights. The weights for the loss function are configured based on the action space type of the environment, as detailed in Table 8.
Table 8: Loss function weights for different environment types.
Loss Term Discrete (Atari, Jericho) Continuous (DMC)
Observation Loss 10.0 10.0
Value Loss 0.5 0.1
Reward Loss 1.0 0.1
Policy Loss 1.0 0.1
A.5 Training Framework and Computational Cost
All experiments were implemented using the LightZero [^4] https://github.com/opendilab/LightZero framework. All multi-task learning experiments were conducted on a single node equipped with 8x NVIDIA A100 (80G) GPUs. The approximate training times are shown in Table 9.
Table 9: Approximate training times for multi-task learning experiments.
Task Approximate Training Time
Atari-8 $\sim$2 days
Atari-26 $\sim$5 days
DMC-18 $\sim$1.5 days
Jericho-4 $\sim$1 day
Appendix B Atari Experiment Details
B.1 Benchmark Setup
We use the Arcade Learning Environment (ALE) (Bellemare et al., 2013b) as our evaluation platform. Our experiments adopt two distinct training paradigms: multitask learning (MT), where a single model is trained to master all tasks simultaneously, and single-task learning (ST), where a dedicated model is trained independently for each task.
For our primary evaluation, we compare our multitask model, ScaleZero, with a robust single-task baseline using individually trained UniZero models. We use ScaleZero to denote the multitask (MT) setting and UniZero to denote the single-task (ST) setting. The MT/ST notation is sometimes omitted for brevity where the context is unambiguous. This comparison strategy is chosen because the multitask version of UniZero performs poorly, and there is currently no widely accepted standard baseline for multitask learning on the Atari benchmark. Consequently, UniZero ST serves as a more challenging and meaningful point of reference.
Our experiments are conducted in two stages: first, we perform model ablation studies on the Atari8 multitask benchmark (comprising Alien, Boxing, ChopperCommand, Hero, MsPacman, Pong, RoadRunner, and Seaquest); subsequently, we evaluate the final model’s performance on the full Atari26 multitask benchmark.
All game observations are preprocessed into 64x64 grayscale images, with 4-frame stacking used for state input. We employ standard environment wrappers, including sticky actions (p=0.25) and a frame skip of 4(Mnih et al., 2013). Performance is measured by the Human-Normalized Score (HNS), and we report both the mean and median scores. To ensure a fair comparison, all environmental settings are aligned with those in the UniZero paper (Pu et al., 2024).
B.2 Supplemental Plasticity Analysis
B.2.1 Training Dynamics and Representational Collapse in UniZero
While Figure 1 in the main text shows performance degradation and worsening plasticity metrics for UniZero on complex tasks, Figure 7 illustrates that on four relatively simpler tasks from the Atari8 set, UniZero’s training process and corresponding metrics remain stable. This suggests a correlation between task difficulty and plasticity degradation.
To further probe the underlying cause of this degradation, we introduce the concept of effective rank (Dohare et al., 2024) to quantify the dimensionality of the model’s representation space. A low effective rank suggests representational collapse, where the model fails to maintain diverse features across tasks. Formally, consider a matrix $\bm{\Phi}\in\mathbb{R}^{n\times m}$ with singular values $\sigma_{k}$ for $k=1,2,\ldots,q$, where $q=\min(n,m)$. Let $p_{k}=\sigma_{k}/\sum_{i=1}^{q}\sigma_{i}$ be the normalized singular values, which form a probability distribution. The effective rank of $\bm{\Phi}$ is then defined as the exponential of the entropy of this distribution:
$\operatorname{erank}(\bm{\Phi})\doteq\exp\left(-\sum_{k=1}^{q}p_{k}\log(p_{k})\right)$ (4)
The effective rank is a continuous measure ranging from 1 to the conventional rank of the matrix, providing a more nuanced view of the dimensionality of the space spanned by the matrix’s columns or rows. In the context of a neural network, the effective rank of a hidden layer’s activation matrix measures the effective number of dimensions in its representation space. A low effective rank indicates that the layer’s output can be generated by a small number of units, implying that many neurons are not contributing useful, independent information. In our experiments, we estimate the effective rank from a minibatch of 256 sequence samples before training on each task.
Figure 6 illustrates this concept, providing a diagnostic view of the representation’s dimensionality. The observation that a standard baseline model can exhibit a low effective rank provides quantitative evidence for the representational collapse hypothesis.
Figure 6: Supplementary analysis of return and representation effective rank. This figure provides additional diagnostics for the baseline model shown in Figure 1, illustrating the relationship between game return (Left) and representation effective rank (Right). The performance degradation, observed as a drop in return, is largely accompanied by a corresponding decline in the representation effective rank. This provides further evidence that the performance collapse is linked to a degeneration of the model’s representation space.
Figure 7: Training dynamics of UniZero on four other tasks from the Atari8 benchmark. This figure complements Figure 1 of the main text, showing the performance curves, dormant neuron ratio of the Transformer backbone, and latent state norm for Boxing, MsPacman, RoadRunner, and Alien.
B.2.2 Complete Metrics for Design Space Ablation Study
Figure 8 provides the complete data corresponding to Figure 3 in the main text, detailing the plasticity metrics for different model variants across all Atari8 tasks. Compared to the Baseline and different variants, the MoE architecture more effectively suppresses the growth of the dormant neuron ratio and the inflation of the latent state norm during training. This suggests that the MoE architecture is better at maintaining representational diversity and plasticity.
Figure 8: Complete plasticity metric analysis for the ScaleZero design space exploration. This figure details the evolution of performance (Return), Dormant Ratio, and Latent Norm for different model variants across all 8 tasks in the Atari8 benchmark.
Intrinsic Connections and Performance Implications
The observed correlations highlight a critical challenge in multi-task learning for standard Transformer architectures: representational collapse. The mechanism of this phenomenon is as follows: the model struggles to maintain distinct representations across tasks, causing its feature space to degenerate into a "one-size-fits-all" low-dimensional space, which is suboptimal for any individual task. This collapse is quantitatively indicated by a low effective rank, a problem diagnosed in baseline models (Section 1). The contraction of the representation space renders many neurons redundant, leading to a rising dormant neuron ratio. To compensate for this impoverished representational capacity, the model is forced to scale the magnitude of remaining active representations, which in turn causes the inflation of the latent norm. Ultimately, this collapsed and impoverished representation space lacks the capacity to model the diversity of multiple tasks, resulting in poor overall performance.
In contrast, the MoE variant is designed to mitigate this very problem. By using its expert-based routing to preserve specialized representational subspaces, it addresses the symptoms of representational collapse. As shown in Figure 8, the MoE architecture effectively promotes neuron activity (suppressing the dormant ratio) and stabilizes representation magnitudes (controlling latent norm inflation), leading to superior multitask performance.
B.3 Performance on the Atari26 Multitask Benchmark
As shown in Table 1 of the main text, the mean HNS of the single multitask ScaleZero (MT) model on the Atari26 benchmark surpasses the average HNS of 26 independently trained UniZero single-task (ST) models. This result suggests that the multitask model achieves positive knowledge transfer across tasks. We hypothesize that this is due to the MoE architecture maintaining the network’s learning plasticity, enabling the transfer of general priors (e.g., object motion, collision detection) from simpler tasks to more complex ones.
However, the multitask model’s median HNS is lower than the single-task baseline. This is primarily influenced by a few tasks with high exploration difficulty or unique mechanics (e.g., PrivateEye, Pitfall), indicating that negative transfer between heterogeneous tasks is not fully resolved in the current architecture. Figure 9 provides a comparison of the learning curves for all tasks.
Figure 9: Performance comparison between ScaleZero (MT) and UniZero (ST) on the Atari26 benchmark. The figure shows the learning curves (mean and 95% confidence interval) for the multitask ScaleZero and the single-task UniZero baselines across 26 Atari games.
Appendix C DMC Experiment Details
C.1 Benchmark Setup
We use the DeepMind Control Suite (DMC) Tunyasuvunakool et al. (2020) as the evaluation benchmark for continuous control tasks. The experiments were conducted on a suite of 18 tasks (e.g., walker_walk, cheetah_run).
Following a similar methodology to our Atari experiments, we compare a single multitask ScaleZero (MT) model against a baseline of 18 single-task UniZero (ST) models, each trained independently on one task. All ScaleZero variants discussed in this section are multitask (MT) models.
The model input consists of low-dimensional state vectors provided by the environment, and the action space is a continuous vector normalized to the [-1, 1] range. The evaluation metric is the Average Return over 3 random seeds.
C.2 Performance Comparison: ScaleZero vs. UniZero
As shown in Figure 10, the multitask ScaleZero model achieves returns comparable or superior to the single-task UniZero baseline on the majority of DMC tasks. This result demonstrates that the ScaleZero architecture is applicable to control problems involving high-dimensional continuous state and action spaces.
We hypothesize that this performance stems from the model learning shared physics priors and control primitives across different robotic morphologies. The expert modules within the MoE architecture may specialize in different types of dynamics (e.g., balancing, locomotion), which are then combined by the gating network based on the current task and state.
Figure 10: Learning curves of ScaleZero (MT) vs. UniZero (ST) on the DMC18 benchmark. This figure compares the performance of the multitask ScaleZero and single-task UniZero models on 18 DMControl tasks. Solid lines represent the mean performance over 3 random seeds, and the shaded area indicates the 95% confidence interval.
C.3 Efficiency Evaluation of ScaleZero-DPS
Figure 11 compares the per-task learning curves of ScaleZero-DPS and the standard ScaleZero model to illustrate the difference in sample efficiency. The Figure shows that ScaleZero-DPS (orange line) exhibits a faster learning rate on most tasks, achieving equivalent performance levels with fewer environment steps. The asterisks (*) in the figure mark tasks that were identified as "solved" by the DPS policy and for which training was terminated early, showing the dynamic resource allocation mechanism in effect.
Figure 11: Performance and efficiency comparison of ScaleZero-DPS vs. standard ScaleZero on DMC18. This figure shows the learning curves (return vs. environment steps) for both models across 18 tasks. An asterisk (*) indicates that the task’s training was terminated early by the DPS policy.
Appendix D Jericho Experiment Details
D.1 Benchmark Setup
Environment Overview: Jericho (Hausknecht et al., 2020) is a reinforcement-learning benchmark built on classic text-adventure games, played entirely through natural-language interaction. Unlike Atari or DMC, it demands robust language understanding of free-form scene, item, and state descriptions while contending with a combinatorial, effectively unbounded action space—at each step the environment surfaces only a small candidate set, yet many legal commands outside it still execute—and coping with sparse, delayed rewards that require long-horizon exploration and planning, making Jericho a stringent testbed for text-based RL.
(a) Zork1
(b) Detective
Figure 12: First five steps of expert interaction trajectories for Zork1 and Detective. At each step, the environment supplies a set of valid candidate actions (used in our experiments); however, additional legal actions outside this set may also be executed, so the action space is, in principle, unbounded.
We conduct experiments on four representative tasks from the Jericho benchmark, namely Detective, Acorncourt, Omniquest, and Zork1. Examples of expert interaction trajectories for Zork1 and Detective are illustrated in Figure 12. Following the protocol of the previous two benchmarks, we compare our multitask ScaleZero (MT) with the single-task UniZero (ST) baseline. ScaleZero uses the same text encoder (bge-base-en-v1.5), the same inference context length of 4, and the same per-step max sequence length of 512 tokens as UniZero. The max action num and max steps for the four tasks are given in Table 10. Results are reported as the mean return over two random seeds. If, at any step, the valid action set is smaller than max action num, we pad to max action num and mark the padded entries with an action mask so they are excluded from sampling.
Figure 13: Learning curves of ScaleZero (MT) vs. UniZero (ST) on the Jericho benchmark. This figure compares the performance of the multitask ScaleZero and single-task UniZero models on 4 Jericho tasks. Solid lines represent the mean performance over 2 random seeds, and the shaded area indicates the 95% confidence interval.
Game
maximum valid actions
(in 200 episodes)
maximum steps per episode
(in 200 episodes)
max action num max steps
Acorncourt 34 17 45 100
Omniquest 24 78 25 100
Zork1 53 396 55 500
Detective 11 51 12 100
Table 10: Key game statistics for the four Jericho tasks. Maximum valid actions denotes the largest number of valid action candidates exposed by the environment at any step, observed over 200 evaluation episodes; this statistic informs the max action num parameter used in our experiments. Maximum steps per episode is the largest episode length observed over the same 200 episodes, while max steps parameter is the episode-length cap adopted during experiments.
D.2 PERFORMANCE COMPARISON: SCALEZERO VS. UNIZERO (ST)
As shown in Figure 13, the multitask ScaleZero attains returns comparable to the single-task UniZero baseline on most Jericho games, with learning curves exhibiting an “easy-first, hard-later” progression. Specifically, ScaleZero converges faster on Acorncourt, Omniquest, and Detective; on the most challenging Zork1, it initially lags but accelerates after  150k environment interactions to reach parity with UniZero. We hypothesize that this stems from multi-task training inducing transferable language priors and interaction routines, which help filter feasible commands, compress the open-ended action space, and reduce fruitless exploration.
Appendix E Gradient Analysis in MoE
E.1 Experimental Analysis
Given the strong performance of Mixture-of-Experts (MoE) in multitask settings, a natural question arises: Why MoE? We aim to investigate, both experimentally and theoretically, the underlying mechanisms behind MoE’s superior performance in multitask reinforcement learning. Specifically, our experimental setup is as follows.
E.1.1 Experiment 1: Analyzing Gradient Conflicts in MoE-based Transformers
We conduct our experiments on Atari-8. Concretely, our network architecture consists of an encoder (ViT), a backbone (Transformer), and corresponding heads. We compare two baseline methods with different backbones:
(1). Naive Transformer: The backbone consists of four standard Transformer blocks.
(2). MoE-based Transformer: The backbone also consists of four Transformer blocks, but the MLP layer in each block is replaced with an MoE layer, which comprises one shared expert and eight non-shared experts(Liu et al., 2024b). Shared experts provide cross-task general representations, enhancing generalization ability and ensuring training stability, while non-shared experts learn task-specific representations to strengthen the model’s discriminability and task adaptability. During the forward pass, all non-shared experts are selectively activated by a sparse gating network, which determines which specific expert to use for each input.
For both baseline models, we investigate gradient conflicts at different components: (1)Input before the MoE layer. (2)Output of the encoder. (3)Parameters of the MoE, including the shared expert, non-shared experts, and the entire MoE layer. We measure gradient conflict between tasks using the maximum negative cosine similarity, defined as follows:
$\text{Max Gradient Conflict}=\max_{i,j}\left(-\frac{\nabla\theta_{i}\cdot\nabla\theta_{j}}{\|\nabla\theta_{i}\|\|\nabla\theta_{j}\|}\right)$ (5)
where $\nabla\theta_{i}$ and $\nabla\theta_{j}$ denote the gradients of the $i$-th and $j$-th tasks, respectively. A higher value of the Max Gradient Conflict represents a greater degree of gradient conflict. We choose the maximum pairwise cosine similarity because it directly identifies the most severe gradient conflict between any two tasks. Unlike averaging operation, which can hide critical issues, the maximum value pinpoints the ’bottleneck pair’ that most significantly impedes stable multi-task training and overall convergence, even if other tasks cooperate well.
In MoE, when computing the gradient of an entire layer, if task $A$ selects expert $i$ while task $B$ selects expert $j$ with $i\neq j$, then the gradient of task $B$ on expert $i$ is filled with zero. The reason is that expert $i$ is not involved in the forward propagation of task $B$, and thus makes no contribution to its loss; consequently, its gradient during backpropagation should naturally be zero.
(a) Conflict comparison of MoE layer gradients.
(b) Conflict comparison of encoder output gradients.
(c) Conflict comparison of MoE input gradients.
Figure 14: Comparison of gradient conflicts between MoE and MLP baselines across different components. MoE-based Transformer exhibits fewer gradient conflicts in MoE input and Moe layer.
Observation 1: As shown in Figure˜14(a), the MoE-based Transformer exhibits fewer gradient conflicts than the MLP counterpart. (b) further shows that introducing MoE reduces gradient conflicts at the layer input, implying that MoE alleviates conflicts in other components to some extent. In the encoder, however, the conflict levels of MoE and MLP remain comparable. We argue that the main reason is that the encoder extracts general representations, which inherently exhibit fewer gradient conflicts.
E.1.2 Experiment 2: Investigating MoE Gating Mechanisms
We are particularly interested in the internal gating mechanism of MoE. Previous studiesChen et al. have shown that, under supervised learning, MoE can implicitly uncover latent cluster structures within input space. However, when faced with non-stationary data distributions generated through agent-environment interactions, it remains unclear whether MoE experts still differentiate effectively. To answer the question, we analyze the entropy of expert selection distributions, which quantifies the uncertainty of the choice of which expert to use for a task, with low entropy indicating high specialization and decisive selection, while high entropy implies uncertainty or an average utilization across multiple experts. We also recorded the Wasserstein distance Rüschendorf (1985) between the expert selection distributions of different tasks, where smaller values indicate a greater proximity to the expert selection of two tasks. We aim to quantify this relationship to find the underlying connections in expert selection among various tasks. The specific experimental steps are as follows:
(STEP1) During a forward pass at a given training step, we record the expert choices in the final MoE-based Transformer block.
(STEP2) For a specific training step $s_{t}$, we collect expert selection data over a past window of size $S$ and compute the frequency of each expert’s activation to form a probability distribution. For a given window size $S$, task $i$ and a specific task $j$, we denote the task selection distribution as $P^{S}_{i}$ and a specific probability for task $j$ as $P^{S}_{i,j}$. Different window sizes reflect attention to data over different temporal scales in non-stationary learning.
(STEP3) Based on this probability distribution, we calculate the expert selection entropy for each task $E_{task_{i}}$:
$E_{task_{i}}=-\sum_{j=1}^{N}P^{S}_{i,j}\log_{2}(P^{S}_{i,j})$
and Wasserstein Distance between task $i$ and task $i^{\prime}$ ($W_{i,i^{\prime}}$)
We consider two sizes $S$: immediate = 100, short = 1,000.555Note that at the very beginning of training, when the window cannot be fully populated, we use the available data within the window, which may lead to larger fluctuations at early steps. We present the entropy results in Figure˜15. Due to space limitations, we show in Figure˜18 the Wasserstein distances between expert selections of different tasks at multiple training stages.
Figure 15: Line plot showing the evolution of expert selection entropy in a multi-task learning setting with eight tasks. The dashed lines correspond to the entropy values of individual tasks, and the solid red line represents the aggregated entropy across all tasks. Higher entropy reflects more uniform and uncertain expert utilization, while lower entropy reflects more concentrated and specialized expert selection.
Observation 2: MoE differentiation and expert specialization. As shown in Figure˜15, as tasks progress, the entropy of the expert distribution gradually decreases, indicating a reduction in distributional uncertainty. This suggests that expert selection becomes concentrated on a few outcomes, reflecting lower uncertainty. Furthermore, in Figure˜18, we visualize the expert selection distributions of different tasks at multiple training stages, offering additional evidence of the progressive specialization in expert utilization.
E.1.3 Experiment 3: Analyzing Gradient Conflicts between shared expert and non-shared expert
To further investigate the relationship between expert selection and gradient dynamics within MoE, we employ a MoE-based Transformer to analyze gradient conflicts across different MoE components. Our experimental results are shown below.
Figure 16: Gradient conflicts across experts in MoE training. The plot shows the evolution of gradient conflicts for the shared expert and 8 non-shared experts during training interations. The logarithmic y-axis reveals that different experts experience varying levels of gradient conflicts, with the shared expert and individual experts (expert0–expert7) demonstrating distinct conflict patterns, indicating effective expert specialization in the multi-task learning framework.
Observation 3: As shown in Figure˜16, the shared expert exhibits significantly higher gradient conflicts compared to the task-specific experts. Among the non-shared experts, the level of conflicts does not show substantial differences. Overall, the shared expert accumulates even more conflicts than the entire MoE layer, indicating that most gradient conflicts within the MoE layer are concentrated on the shared expert. The introduction of several task-specific experts effectively reduces the overall gradient conflicts of the MoE layer, which is consistent with our theoretical analysis in Corollary˜E.4.
We further investigated the conflicts among different experts within MoE. Results is shown in Figure˜16 We found that shared experts bear most of the gradient conflicts, whereas individual (non-shared) experts experience almost no conflict. A plausible explanation is that in a standard gating MoE, different samples are routed to different experts, so each expert primarily receives gradient updates from a specific type of sample. This effectively performs an implicit task partitioning, leading to generally consistent gradient directions within the same expert and minimal conflicts. In contrast, in shared-expert MoE, all samples pass through the shared experts, whose parameters must adapt simultaneously to multiple tasks, diverse semantics, and even different domain distributions. This can result in highly inconsistent gradient directions, significantly increasing conflicts and making it difficult for shared experts to balance diverse task requirements during optimization.This observation naturally raises a simple question: is the alleviation of gradient conflicts mainly due to the sparse selection by the gating network? We provide a theoretical justification from the perspective of gating.
E.2 Theoretical Analysis
We observe that the magnitude of the gradient conflict in an MoE model is primarily influenced by the routing coefficient, $\rho_{i}$. In a sparse MoE, this coefficient acts as a binary variable determined by the router network. For simlicity, let us consider a top-1 sparse MoE, where only the expert with the highest routing score is selected for a given input. An effective router would learn to assign tasks that are prone to high gradient conflict to different experts. In this case, the routing coefficient for any conflicting task pair would be zero, thereby nullifying their contribution to the overall conflict. Conversely, a poorly performing router might allocate conflicting tasks to the same expert, in which case the conflict remains unmitigated.
Theorem E.1 (Upper Bound of Full-layer MoE Gradient Conflict with Sparse/Soft Routing).
Consider a MoE layer with $M$ experts. Let each expert’s gradient norm satisfy $\|g_{t}^{(m)}\|\leq R,\quad m=1,\dots,M,$ and let routing weights for a task be $\lambda_{1},\dots,\lambda_{M}\geq 0$. The full-layer gradient is $G_{t}=\operatorname{concat}(\lambda^{t}_{1}g_{t}^{(1)},\dots,\lambda^{t}_{M}g_{t}^{(M)})$, Assume that for any task pair $(t_{1},t_{2})$, the gradient conflict on a single expert is bounded by $G$. Then the full-layer MoE gradient conflict satisfies the explicit upper bound
$\mathrm{conflict}(G_{t_{1}},G_{t_{2}})\leq G\cdot\frac{\displaystyle\sum_{m=1}^{M}\lambda^{t_{1}}_{m}\lambda^{t_{2}}_{m}\,\|g_{t_{1}}^{(m)}\|\,\|g_{t_{2}}^{(m)}\|}{\displaystyle\sqrt{\sum_{m=1}^{M}(\lambda^{t_{1}}_{m})^{2}\|g_{t_{1}}^{(m)}\|^{2}}\;\sqrt{\sum_{m=1}^{M}(\lambda^{t_{2}}_{m})^{2}\|g_{t_{2}}^{(m)}\|^{2}}}\leq G$
Proof.
From the block-level conflict bound we have
$\mathrm{conflict}(G_{t_{1}},G_{t_{2}})\leq G\cdot\frac{\sum_{m=1}^{M}\lambda^{t_{1}}_{m}\lambda^{t_{2}}_{m}\,\|g_{t_{1}}^{(m)}\|\,\|g_{t_{2}}^{(m)}\|}{\|G_{t_{1}}\|\,\|G_{t_{2}}\|},$
with
$\|G_{t_{1}}\|=\sqrt{\sum_{m=1}^{M}(\lambda^{t_{1}}_{m})^{2}\|g_{t_{1}}^{(m)}\|^{2}},\qquad\|G_{t_{2}}\|=\sqrt{\sum_{m=1}^{M}(\lambda^{t_{2}}_{m})^{2}\|g_{t_{2}}^{(m)}\|^{2}}.$
Define vectors $u,v\in\mathbb{R}^{M}$ as
$u_{m}=\lambda^{t_{1}}_{m}\|g_{t_{1}}^{(m)}\|,\qquad v_{m}=\lambda^{t_{2}}_{m}\|g_{t_{2}}^{(m)}\|.$
Then the numerator equals $\langle u,v\rangle$, and the denominator equals $\|u\|\,\|v\|$. By the Cauchy–Schwarz inequality,
$\langle u,v\rangle\leq\|u\|\,\|v\|,$
which implies that the fraction is at most $1$. Substituting this bound back, we obtain
$\mathrm{conflict}(G_{t_{1}},G_{t_{2}})\leq G.$
Equality holds if and only if $u$ and $v$ are collinear (i.e., there exists $c>0$ such that u=cvu=c\,v), and each block-level conflict achieves its upper bound simultaneously. ∎
Remark 1 (Effect of Routing Strategies on Full-layer Gradient Conflict).
The upper bound of full-layer MoE gradient conflict depends strongly on the routing strategy. We summarize two representative cases:
a. Dense Routing (Soft Gating, $\sum_{m=1}^{M}\lambda_{m}=1$)
  • All experts contribute to the full-layer gradient.
  • The fraction in Theorem 1 can be interpreted as the cosine similarity between the weighted vectors $u=\lambda^{t_{1}}\odot\|g_{t_{1}}\|$ and $v=\lambda^{t_{2}}\odot\|g_{t_{2}}\|$:
    $\frac{\sum_{m=1}^{M}\lambda^{t_{1}}_{m}\lambda^{t_{2}}_{m}\|g_{t_{1}}^{(m)}\|\|g_{t_{2}}^{(m)}\|}{\|G_{t_{1}}\|\,\|G_{t_{2}}\|}=\cos(u,v)\leq 1.$
  • If the expert norms are roughly equal ($\|g_{t_{1}}^{(m)}\|\approx\|g_{t_{2}}^{(m)}\|$), the conflict upper bound is largely determined by $\cos(\lambda^{t_{1}},\lambda^{t_{2}})$.
  • Uniform weights ($\lambda_{m}=1/M$) do not automatically reduce conflict; alignment of weighted gradients matters more.
  • Overlap or alignment of routing vectors increases conflict, while orthogonal or disjoint routing vectors reduce it.
b. Sparse Routing (Top-1 gating, one $\lambda_{m}=1$, others 0)
  • Non-overlapping case (different tasks select different experts):Each task’s full-layer gradient contains only its chosen expert. Inner products between task gradients are near zero because tasks lie in different expert subspaces. Full-layer conflict is strictly less than $G$, possibly near 0.
  • Collapsed case (all tasks select the same expert): Full-layer gradient reduces to the chosen expert’s gradient. Conflict equals the single-expert bound $G$, since all gradients reside in the same subspace.
Therefore, the key to mitigating gradient conflict lies in the router’s ability to foster specialization among experts, ensuring that tasks with potentially conflicting objectives are handled by distinct, differentiated experts.
This naturally raises a critical question: Can a Mixture-of-Experts architecture indeed achieve this, and if so, what are the underlying mechanisms that enable this capability? In the following, we answer this question affirmatively. By building upon existing foundational theories of Mixture-of-ExpertsChen et al. , we present a formal derivation that explains precisely when and how this is achieved.
E.2.1 Analysis
Previews workChen et al. theoretically demonstrated that, in a single-task supervised learning setting, when the input distribution exhibits distinguishable cluster structures, the sparse gating Mixture-of-Experts (MoE) router can automatically learn the cluster center features and route samples to the most suitable expert, thereby significantly improving performance. We further hypothesize that, in multi-task learning, the input spaces of different tasks naturally correspond to distinct clusters and the signal for the cluster center can be task ID or the latent pattern on the state space. Under this assumption, the conclusions from the original work can be directly extended to the multi-task setting.
Next, we adapt Chen et al. theoretical analysis to the reinforcement learning context using a tabular example. Specifically, we treat each task as an independent Markov Decision Process (MDP) and construct a multi-task MDP set with a shared state space but task-specific transitions. In this setting, we derive the gradient conflicts that arise in the MoE layer.
We also discuss the case where no clustering signal exists between tasks and point out that the introduction of task embeddings can significantly enhance the clustering structure between tasks, thereby promoting the isolation of expert selection across different tasks in MoE.
E.2.2 Problem setting: multi-task MDP with task-specific latent clusters
We consider a Markov decision process (MDP) defined by the tuple $(\mathcal{S},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $P(s^{\prime}\!\mid\!s,a)$ the transition kernel, $R(s,a)$ the reward function, and $\gamma\in[0,1)$ the discount factor. There are $K$ tasks $T_{1},\dots,T_{K}$. Each task $T_{t}$ corresponds to a disjoint state subspace $\mathcal{S}^{(t)}$, so that
$\mathcal{S}=\bigcup_{t=1}^{K}\mathcal{S}^{(t)},\qquad\mathcal{S}^{(t)}\cap\mathcal{S}^{(t^{\prime})}=\varnothing\quad\text{for }t\neq t^{\prime}.$
Patch-based state representation.
Each state $s\in\mathcal{S}^{(t)}_{k}$ (i.e., belonging to task $t$ and cluster $k$) is represented as an unordered collection of $P$ patches in $\mathbb{R}^{d}$. The patches are randomly permuted before being presented to the model. Every state contains the following patch types:
  1. 1.
    Action-signal patch: exactly one patch equals $\alpha\,v^{(t)}_{k}$, where $v^{(t)}_{k}\in\mathbb{R}^{d}$ encodes the optimal-action feature.
  2. 2.
    Cluster-center patch: exactly one patch equals $\beta\,c^{(t)}_{k}$, indicating the (task-specific) cluster identity; a router (e.g., in a Mixture-of-Experts model) is expected to detect this signal to decide routing.
  3. 3.
    Feature-noise (confounder) patch: exactly one patch equals $\epsilon\gamma\,v^{(t)}_{k^{\prime}}$, where $k^{\prime}$ is typically a different cluster index in the same task $t$, modeling an intra-task confounding feature.
  4. 4.
    Random-noise patches: the remaining P3P-3 patches are i.i.d. draws from an isotropic Gaussian, $\mathcal{N}\!\Big{(}0,\tfrac{\sigma_{p}^{2}}{d}I_{d}\Big{)}.$
Thus the encoder receives an unordered set $\{x_{1},\dots,x_{P}\}\subset\mathbb{R}^{d}$ containing exactly one action signal, one cluster-center signal (task-specific), one intra-task confounder, and Gaussian noise patches.
MoE Model with Expert Specialization
We now analyze how a Mixture-of-Experts (MoE) model can leverage expert specialization to address the aforementioned issues.
Experts. We consider $M$ linear experts. The value function of expert $m$ is defined as
$V_{m}(s;w_{m})=\langle w_{m},\phi(s)\rangle.$ (6)
Router. The gating network assigns a score to each expert $m$:
$h_{m}(s;\theta_{m})=\langle\theta_{m},\phi(s)\rangle.$ (7)
MoE Output. The overall value function is given by
$V_{\text{MoE}}(s;W,\Theta)=\sum_{m=1}^{M}\pi_{m}(s)V_{m}(s;w_{m}),$ (8)
where the softmax gating weights are
$\pi_{m}(s)=\frac{\exp(h_{m}(s))}{\sum_{j=1}^{M}\exp(h_{j}(s))}.$ (9)
Loss Function. We minimize the mean squared Bellman error (MSBE) with a stop-gradient:
$L(W,\Theta)=\frac{1}{2}\sum_{s\in\mathcal{S}}\Big{(}V_{\mathrm{MoE}}(s)-\operatorname{stop\_grad}\big{(}(\mathcal{T}^{*}V_{\mathrm{MoE}})(s)\big{)}\Big{)}^{2}=\frac{1}{2}\sum_{s\in\mathcal{S}}\delta(s)^{2},$ (10)
where $\delta(s)$ denotes the temporal-difference (TD) error.
E.2.3 Learning Dynamics Analysis
We follow an analysis path similar to prior workChen et al. , dividing the learning process into two stages: an early expert exploration stage and a later router learning stage.
Initialization.
The expert weights $w_{m}^{(0)}$ are randomly initialized from a small zero-mean Gaussian distribution:
$w_{m}^{(0)}\sim\mathcal{N}(0,\sigma_{w}^{2}I),$ (11)
while the router weights $\theta_{m}^{(0)}$ are initialized as zero vectors:
$\theta_{m}^{(0)}=\mathbf{0}.$ (12)
Expert Exploration Stage.
According to [Chen et al. , Lemma E.3], the zero initialization of the router implies that at the beginning of training ($t=0$),
$\max_{m\in[M]}\left|P(m_{i,t}=m)-\frac{1}{M}\right|=\tilde{O}(\sigma_{0}^{1.5})\quad\text{for all }i\in[n],\,m\in[M].$
[…]
$\pi_{m}(s)\approx\frac{1}{M}.$ (13)
This ensures that each expert has an equal opportunity to learn from the data.
During the early stage of training, due to random initialization, each expert $m$ has a weight vector $w_{m}^{(0)}$ that exhibits a slightly larger inner product with some value basis vector $v_{k}$. We define the initial preference cluster of expert $m$ as
$k_{m}^{*}=\arg\max_{k\in[K]}\left|\langle w_{m}^{(0)},v_{k}\rangle\right|.$ (14)
Under gradient descent updates, the weight vector $w_{m}$ of expert $m$ will predominantly grow along the direction of its preferred basis vector $v_{k_{m}^{*}}$.[Chen et al. , Lemma E.5] In other words, during training, an expert progressively specializes in modeling a specific cluster $c_{k}$. Building upon this, we derive an upper bound on the gradient conflicts during the Expert Exploration Stage.
Theorem E.2 (Upper Bound on Single-Expert and Full-Layer MoE Gradient Conflict with Uniform Sparse Routing).
Consider a Mixture-of-Experts (MoE) layer with $M$ experts and $K$ independent tasks. Each task independently selects one expert uniformly at random: $P=\frac{1}{M}.$
Then:
1. For any single expert $m$, the expected gradient conflict is upper bounded by
$\mathbb{E}[\text{conflict}_{t_{1},t_{2}}^{(m)}]\leq G\,q_{\text{single}},\quad q_{\text{single}}=1-\Big{(}1-\frac{1}{M}\Big{)}^{K}-K\frac{1}{M}\Big{(}1-\frac{1}{M}\Big{)}^{K-1}.$
2. For the full-layer MoE with sparse router, the expected gradient conflict is upper bounded by
$\mathbb{E}[\text{conflict}_{t_{1},t_{2}}^{\text{MoE}}]\leq G\,q_{\text{layer}},\quad q_{\text{layer}}=\begin{cases}1-\dfrac{M!}{(M-K)!\,M^{K}},&K\leq M\\[10.00002pt] 1,&K>M\end{cases}.$
Proof.
Consider an arbitrary expert $m$ and the $K$ tasks. Each task independently selects expert $m$ with probability $1/M$. Let $X$ denote the number of tasks that select expert $m$. Then
$X\sim\mathrm{Bin}(K,1/M).$
A single-expert gradient conflict occurs if and only if $X\geq 2$, and by assumption, the maximum conflict for a single expert is $G$. Therefore, the probability that expert $m$ experiences a conflict is
$\Pr(X\geq 2)=1-\Big{(}1-\frac{1}{M}\Big{)}^{K}-K\frac{1}{M}\Big{(}1-\frac{1}{M}\Big{)}^{K-1}.$
By linearity of expectation, the expected gradient conflict for a single expert is upper bounded by
$\mathbb{E}[\text{conflict}_{t_{1},t_{2}}^{(m)}]\leq G\,q_{\text{single}}.$
The full-layer gradient $G_{t}$ is constructed by concatenating all expert gradients. A full-layer conflict occurs if at least one expert has a conflict (i.e., at least two tasks select that expert).
- For $K\leq M$, the probability that all $K$ tasks select distinct experts (no collision) is given combinatorially by
$\frac{M!}{(M-K)!\,M^{K}},$
so the probability of at least one collision (full-layer conflict) is
$q_{\text{layer}}=1-\frac{M!}{(M-K)!\,M^{K}}.$
- For $K>M$, by the pigeonhole principle, at least one expert must be selected by two or more tasks, so
$q_{\text{layer}}=1.$
Each conflict in an expert can contribute at most $G$ to the full-layer gradient conflict. By worst-case analysis and linearity of expectation, the expected full-layer MoE gradient conflict is therefore upper bounded by
$\mathbb{E}[\text{conflict}_{t_{1},t_{2}}^{\text{MoE}}]\leq G\,q_{\text{layer}}.$
- The single-expert bound $G\,q_{\text{single}}$ gives a local (per-expert) expected conflict. - The full-layer bound $G\,q_{\text{layer}}$ accounts for collisions across all experts in the concatenated MoE layer. - Together, they describe the expected gradient conflict behavior of a MoE layer with $K$ tasks and $M$ experts under uniform sparse routing. ∎
Router Learning Stage
During the router-learning stage, the sparse MoE layer exhibits two decisive properties that justify its use in a multi-task setting.
First, the gating network rapidly identifies the latent task structure: after only
$T_{2}=\left\lfloor\eta^{-1}M^{-2}\right\rfloor$
iterations, its weight vectors $\bm{\theta}_{m}$ become strongly aligned with the cluster-center signal $\mathbf{c}_{k}$ of every task $k$ , while simultaneously suppressing spurious correlations with label signals and noise.[Chen et al. -lemma E.14] Consequently, any input drawn from task $k$ is routed to the corresponding subset of experts $\mathcal{M}_{k}$ with probability $1-o(\frac{1}{d})$, even though the model is never given explicit task labels. [Chen et al. -Lemma E.18]
Second, each expert in $\mathcal{M}_{k}$ remains tightly specialised to its assigned task, but only chance-level performance on all other tasks. [Chen et al. -Lemma E.12,Lemma 5.2].
Taken together, these two properties guarantee automatic task separation and expert specialisation without external supervision, making the MoE layer an effective backbone for multi-task problems in which tasks form separable clusters in the input space.
Theorem E.3 (Expected Gradient Conflict on Task-specific Expert Sets).
Consider a Mixture-of-Experts (MoE) model with $K$ tasks and $M$ experts. Let tasks $i$ and $j$ have expert sets $S_{i}$ and $S_{j}$ with sizes $|S_{i}|=a$, $|S_{j}|=b$, and intersection $U=|S_{i}\cap S_{j}|$. Assume that each task selects an expert from its own expert set with probability 1O(1/d)1-O(1/d) uniformly, and from the complement set with probability $O(1/d)$ uniformly. Let $G$ denote the maximum gradient conflict if two tasks select the same expert. Then the expected gradient conflict between tasks $i$ and $j$ satisfies
$\mathbb{E}[\text{conflict}_{i,j}]\leq G\cdot\Big{(}P^{(1)}+P^{(2)}+P^{(3)}\Big{)},$ (15)
where
$P^{(1)}$ $=U\cdot\left(\frac{1-O(1/d)}{a}+\frac{O(1/d)}{M-a}\right)\left(\frac{1-O(1/d)}{b}+\frac{O(1/d)}{M-b}\right),$ (16)
$P^{(2)}$ $=(a-U+b-U)\cdot\left(\frac{1-O(1/d)}{a}+\frac{O(1/d)}{M-a}\right)\frac{O(1/d)}{M-b},$ (17)
$P^{(3)}$ $=(M-(a+b-U))\cdot\frac{O(1/d)}{M-a}\cdot\frac{O(1/d)}{M-b}.$ (18)
In particular, the dominant contribution comes from the shared experts (first class), leading to
$\mathbb{E}[\text{conflict}_{i,j}]\approx G\cdot\frac{U}{ab}+O(G/d).$ (19)
Moreover, for the full-layer MoE gradient formed by concatenating all expert parameters, the expected layer-level gradient conflict satisfies the same upper bound:
$\mathbb{E}[\text{conflict}]_{\text{layer}}\leq G\cdot\Big{(}P^{(1)}+P^{(2)}+P^{(3)}\Big{)}\approx G\cdot\frac{U}{ab}+O(G/d).$ (20)
Proof.
Consider two tasks $i$ and $j$ with expert sets $S_{i}$ and $S_{j}$. Partition the set of all experts into three classes: (1) shared experts $S_{i}\cap S_{j}$, (2) task-specific experts $S_{i}\setminus S_{j}$ and $S_{j}\setminus S_{i}$, and (3) unrelated experts outside $S_{i}\cup S_{j}$. For each class, we compute the probability that both tasks select the same expert.
For a shared expert $e\in S_{i}\cap S_{j}$, task $i$ selects it either from its own set with probability (1O(1/d))/a(1-O(1/d))/a or from outside with probability O(1/d)/(Ma)O(1/d)/(M-a). Similarly for task $j$. Therefore, the probability that both tasks select $e$ is
$p^{(1)}=\left(\frac{1-O(1/d)}{a}+\frac{O(1/d)}{M-a}\right)\left(\frac{1-O(1/d)}{b}+\frac{O(1/d)}{M-b}\right),$
and there are $U$ such experts, giving total contribution $P^{(1)}=U\cdot p^{(1)}$.
For task-specific experts $e\in(S_{i}\setminus S_{j})\cup(S_{j}\setminus S_{i})$, one task can select it from its own set while the other selects from outside. This yields
$p^{(2)}=\left(\frac{1-O(1/d)}{a}+\frac{O(1/d)}{M-a}\right)\frac{O(1/d)}{M-b},$
with a total of (aU+bU)(a-U+b-U) experts, giving $P^{(2)}=(a-U+b-U)\cdot p^{(2)}$.
For unrelated experts $e\notin S_{i}\cup S_{j}$, both tasks must select from outside, giving
$p^{(3)}=\frac{O(1/d)}{M-a}\cdot\frac{O(1/d)}{M-b},$
and there are M(a+bU)M-(a+b-U) such experts, yielding $P^{(3)}=(M-(a+b-U))\cdot p^{(3)}$.
Summing over all classes and multiplying by $G$, we obtain the expected single-expert gradient conflict:
$\mathbb{E}[\text{conflict}_{i,j}]\leq G\cdot\big{(}P^{(1)}+P^{(2)}+P^{(3)}\big{)}.$
Since the layer-level gradient is formed by concatenating all expert gradients, the layer-level conflict is the sum over all expert contributions, hence
$\mathbb{E}[\text{conflict}]_{\text{layer}}\leq G\cdot\big{(}P^{(1)}+P^{(2)}+P^{(3)}\big{)}.$
Finally, the dominant term is from the shared experts, giving the simplified approximation
$\mathbb{E}[\text{conflict}]_{\text{layer}}\approx G\cdot\frac{U}{ab}+O(G/d),$
where $O(G/d)$ absorbs all higher-order small contributions from selecting non-preferred experts. ∎
Corollary E.4 (Expected Full-layer MoE Gradient Conflict for $K$ Tasks).
Consider a MoE layer with $M$ experts and $K$ tasks. Each task $t$ has its own expert set $S_{t}$ with size $|S_{t}|=a_{t}$. Each expert’s gradient norm satisfies $\|g_{t}^{(m)}\|\leq R$, and if two tasks select the same expert, the maximum gradient conflict is $G$. Each task chooses an expert according to the following rule: with probability 1O(1/d)1-O(1/d) it selects uniformly from its own set, and with probability $O(1/d)$ it selects uniformly from the remaining experts.
For any subset of tasks $\mathcal{T}\subseteq\{1,2,\dots,K\}$, let the number of shared experts be
$U_{\mathcal{T}}=\left|\bigcap_{t\in\mathcal{T}}S_{t}\right|.$
Then the expected full-layer gradient conflict is upper bounded by
$\mathbb{E}[\text{conflict}]_{\text{layer}}\leq G\sum_{\mathcal{T}\subseteq[K],|\mathcal{T}|\geq 2}U_{\mathcal{T}}\prod_{t\in\mathcal{T}}\left(\frac{1-O(1/d)}{a_{t}}+\frac{O(1/d)}{M-a_{t}}\right),$ (21)
and its leading term can be approximated as
$\mathbb{E}[\text{conflict}]_{\text{layer}}\approx G\sum_{\mathcal{T}\subseteq[K],|\mathcal{T}|\geq 2}\frac{U_{\mathcal{T}}}{\prod_{t\in\mathcal{T}}a_{t}}+O(G/d).$ (22)
Proof.
For each task subset $\mathcal{T}\subseteq\{1,2,\dots,K\}$ with $|\mathcal{T}|\geq 2$, consider any expert $e\in\bigcap_{t\in\mathcal{T}}S_{t}$. The probability that all tasks in $\mathcal{T}$ simultaneously select this expert is given by the product of individual selection probabilities. For each task $t\in\mathcal{T}$, the probability of choosing $e$ is
$\frac{1-O(1/d)}{a_{t}}+\frac{O(1/d)}{M-a_{t}},$
where the first term accounts for selecting $e$ from the task’s own expert set and the second term accounts for selecting $e$ from outside the set. Since the tasks are independent, the joint probability that all tasks in $\mathcal{T}$ select the same shared expert $e$ is
$\prod_{t\in\mathcal{T}}\left(\frac{1-O(1/d)}{a_{t}}+\frac{O(1/d)}{M-a_{t}}\right).$
Multiplying by the number of shared experts $U_{\mathcal{T}}$ gives the total expected conflict contribution from this subset:
$P_{\mathcal{T}}=U_{\mathcal{T}}\prod_{t\in\mathcal{T}}\left(\frac{1-O(1/d)}{a_{t}}+\frac{O(1/d)}{M-a_{t}}\right).$
Summing over all task subsets with size at least 2, and multiplying by $G$, yields the full-layer expected gradient conflict:
$\mathbb{E}[\text{conflict}]_{\text{layer}}\leq G\sum_{\mathcal{T}\subseteq[K],|\mathcal{T}|\geq 2}P_{\mathcal{T}}.$
In the leading order, we can neglect the $O(1/d)$ small-probability contributions from selecting experts outside each task’s own set, which gives
$\mathbb{E}[\text{conflict}]_{\text{layer}}\approx G\sum_{\mathcal{T}\subseteq[K],|\mathcal{T}|\geq 2}\frac{U_{\mathcal{T}}}{\prod_{t\in\mathcal{T}}a_{t}}+O(G/d),$
where the dominant contribution comes from shared experts that are simultaneously selected by multiple tasks, and the higher-order terms are absorbed in $O(G/d)$. ∎
Figure 17: Nine heatmaps illustrating the evolution of expert selection distributions in the MoE-based Transformer across different tasks and training interations. Each row corresponds to a specific task, where the color intensity reflects the frequency of expert selections within a sliding window of $S=1000$. This visualization reveals how task-specific expert utilization patterns emerge and evolve during training.
E.2.4 Task Embedding for Improved Task Separation in Multi-Task Learning
In multi-task learning, tasks often share overlapping input spaces and underlying features (e.g., sentiment classification tasks using the same text encoder), making it difficult for the routing network to distinguish tasks. This overlap leads to routing confusion, where data from one task may be misrouted to experts specialized in other tasks, reducing performance. To address this, task embedding introduces a learnable vector $e_{\tau}$ for each task, which can be combined with input features. Since $e_{\tau}$ is orthogonal across tasks (or constrained via a regularization term), it helps the routing network more clearly distinguish between tasks, even when inputs are similar. Task embedding also acts as a regularization term, reducing expert load imbalance and improving the stability and performance of multi-task MoE training.
Our theoretical analysis suggests that introducing distinct task embeddings provides crucial clustering signals, enabling the router to more effectively differentiate between tasks. This, in turn, mitigates gradient conflicts and enhances overall performance. However, our empirical results, as presented in Figure˜3, show that the inclusion of naive task embeddings did not yield a significant improvement. We hypothesize that this is due to the absence of explicit constraints to maintain separation between the embeddings during optimization, which could lead to their collapse in the later stages of training. For future work, we plan to investigate more effective methods for explicitly injecting task-specific information. The goal is to empower the router to autonomously leverage the unique priors of each task, thereby further advancing multi-task learning performance.
Figure 18: Nine heatmaps illustrating the Wasserstein distances between expert selection distributions of different tasks at successive training interations with $S_{immediate}=100$. Each heatmap corresponds to a specific training interation, revealing how inter-task expert selection similarity changes over time . Diagonal elements are removed to exclude self-comparisons.
Figure 19: T-SNE Visualization. Each point denotes a 2D projection of a 768-dimensional latent embedding extracted from the world model. Points are colored by environment ID, with different colors representing different Atari games. The visualization reveals the learned representational structure: similar game states cluster together in the latent space, while distinct environments form separate groups. This indicates that the model captures task-specific features while preserving meaningful separation across different games. The figure contains 8,000 data points collected from 8 distributed training processes, where each point corresponds to a single time step from the observation sequence (400 samples × 20 time steps per sample).

### v2

@@
 Abstract
-In heterogeneous multi-task learning, tasks not only exhibit diverse observation and action spaces but also vary substantially in intrinsic difficulty.
-While conventional multi-task world models like UniZero excel in single-task settings, we find that when handling large-scale heterogeneous environments, gradient conflicts and the loss of model plasticity often constrain their sample and computational efficiency.
-In this work, we address these challenges from two perspectives: the single learning iteration and the overall learning process.
-First, we investigate the impact of key design spaces on extending UniZero to multi-task planning.
-We find that a Mixture-of-Experts (MoE) architecture provides the most substantial performance gains by mitigating gradient conflicts, leading to our proposed model, ScaleZero.
-Second, to dynamically balance the computational load across the learning process, we introduce an online, LoRA-based dynamic parameter scaling (DPS) strategy.
+In heterogeneous multitask decision-making, tasks not only exhibit diverse observation and action spaces but also vary substantially in their underlying complexities.
+While conventional multitask world models like UniZero excel in single-task settings, we find that when handling a broad and diverse suite of tasks, gradient conflicts and the loss of model plasticity often constrain their sample efficiency.
+In this work, we address these challenges from two complementary perspectives: the single learning iteration and the overall learning process.
+First, to mitigate the gradient conflicts, we systematically investigate key architectural designs for extending UniZero.
+Our investigation identifies a Mixture-of-Experts (MoE) architecture as the most effective approach.
+We demonstrate, both theoretically and empirically, that this architecture alleviates gradient conflicts by routing task-specific representations to specialized sub-networks.
+This finding leads to our proposed model, ScaleZero.
+Second, to dynamically allocate model capacity throughout the learning process, we introduce an online Dynamic Parameter Scaling (DPS) strategy.
 This strategy progressively integrates LoRA adapters in response to task-specific progress, enabling adaptive knowledge retention and parameter expansion.
-Empirical evaluations on standard benchmarks such as Atari, DMControl (DMC), and Jericho demonstrate that ScaleZero, relying exclusively on online reinforcement learning with one model, attains performance on par with specialized single-task baselines.
-Furthermore, when augmented with our dynamic parameter scaling strategy, our method achieves competitive performance while requiring only 80% of the single-task environment interaction steps.
-These findings underscore the potential of ScaleZero for effective large-scale multi-task learning.
-Our code is available at https://github.com/opendilab/LightZero.
+Evaluations on a diverse set of standard benchmarks (Atari, DMC, Jericho) demonstrate that ScaleZero, utilizing solely online reinforcement learning with one model, performs on par with specialized single-task agents.
+With the DPS strategy, it remains competitive while using just 71.5% of the environment interactions.
+These findings underscore the potential of ScaleZero for effective multitask planning.
 1 Introduction
-Unified world models (Reed et al., 2022; Gallouédec et al., 2024b; Lee et al., 2022b) represent a significant step towards generalist agents, offering a single, cohesive framework for multi-modal perception, long-horizon prediction, and decision-making by learning a shared latent space representation.
-The true power of these models is unlocked when combined with search-based planning algorithms like Monte Carlo Tree Search (MCTS) (Silver et al., 2017; Schrittwieser et al., 2019), which can perform efficient lookahead within this compact latent space.
-MCTS excels at dynamically balancing exploration and exploitation in complex decision spaces, a combination that has achieved well-established success in homogeneous task domains like board games (Ye et al., 2021; Silver et al., 2017; Schrittwieser et al., 2019).
-The natural next frontier is to extend this success to heterogeneous multi-task reinforcement learning (MTRL) (Yu et al., 2024), aiming to build a single, generalist agent that can master a wide array of diverse environments.
-However, this ambition is impeded by a formidable obstacle: training a single shared model on diverse tasks with potentially conflicting dynamics and objectives is notoriously difficult.
-While prior work (Fernando et al., 2023; Shi et al., 2023) has attempted to mitigate inter-task interference, these methods have been predominantly studied in supervised learning settings.
-The challenges and dynamics of multi-task planning within a shared world model remain largely unexplored.
-Our investigation into this specific domain of multi-task planning identifies and addresses two critical, intertwined obstacles that are not fully resolved by existing approaches: (1) Representational bottlenecks and plasticity collapse: In diverse multi-task settings, a shared model is susceptible to gradient dominance from simpler, faster-converging tasks (Cho et al., 2024).
+Unified world models (Reed et al., 2022; Lee et al., 2022; Gallouédec et al., 2024) represent a significant step towards generalist agents, offering a cohesive framework for multi-modal perception  (Zhang et al., 2022), long-horizon prediction (Ni et al., 2024), and decision-making by learning a shared latent space representation.
+By encoding the environment into a compact latent space, these learned representations enable effective planning.
+Algorithms like Monte Carlo Tree Search (MCTS) (Świechowski et al., 2023) are well-suited for this task, as they can perform lookahead efficiently within this abstract representation (Schrittwieser et al., 2019; 2021).
+MCTS excels at dynamically balancing exploration and exploitation, a combination that has achieved well-established success in homogeneous task domains like board games (Schrittwieser et al., 2019; Ye et al., 2021).
+Extending this paradigm to heterogeneous multitask reinforcement learning (MTRL) (Yu et al., 2024) is a key objective for creating generalist agents.
+However, this ambition is impeded by a formidable obstacle: training a single shared model on diverse tasks with potentially conflicting dynamics and objectives is notoriously difficult (Cho et al., 2024).
+While prior work (Sun et al., 2022; Fernando et al., 2023; Shi et al., 2023) has attempted to mitigate inter-task interference, these methods have been predominantly studied in supervised learning and model-free RL settings.
+The challenges and dynamics of multitask planning within a shared world model remain largely unexplored.
+Within this specific domain of multitask planning, our work identifies and addresses two critical, intertwined obstacles not fully resolved by existing methods: (1) Representational bottlenecks and plasticity collapse: In diverse multitask settings, a shared model is susceptible to gradient dominance from simpler, faster-converging tasks (Cho et al., 2024).
 This imbalance leads to representation interference (Yu et al., 2020; Bejnordi et al., 2024), where the learning signals for more complex tasks are suppressed.
-The consequence is a progressive loss of network plasticity (Dohare et al., 2024; Todorov et al., 2025)the fundamental ability of a model to adjust its parameters to learn from new data–imposing a hard ceiling on the model’s overall learning capacity and leading to performance collapse on challenging tasks.
-(2) Static resource allocation: Conventional architectures employ a uniform, one-size-fits-all forward pass, applying the same computational effort to every task irrespective of its intrinsic difficulty.
-This static strategy results in profound computational inefficiency, as resources are squandered on already-mastered tasks instead of being directed toward those that require further learning.
-To dissect these issues, we first conduct a rigorous empirical diagnosis in Figure 1.
-By deploying a baseline UniZero model across eight canonical Atari games, we observe a stark failure pattern.
-While the model rapidly masters simple tasks like Pong, it exhibits initial progress followed by a catastrophic performance collapse on complex, visually distinct games like Seaquest and ChopperCommand.
-We provide the quantitative evidence linking this failure to specific internal model dynamics: an uncontrolled inflation of the latent state norm and a corresponding spike in the dormant neuron ratio within the Transformer backbone, signaling a descent into a "frozen," non-adaptive state.
-Motivated by this diagnosis, we undertake a principled exploration of the UniZero design space across the axes of input, model, and optimization from the the single learning iteration perspective.
-This systematic investigation, which evaluates task embeddings, encoder architectures (ResNet (He et al., 2016) vs.
-ViT (Dosovitskiy et al., 2020)), latent normalization schemes (Hansen et al., 2023), sparse Mixture-of-Experts (MoE) backbones, and gradient rectification methods(He et al., 2020), culminates in a new, powerful model we term ScaleZero.
-Validated on a comprehensive suite of online RL benchmarks—spanning 26 Atari games, 18 DeepMind Control tasks, and 4 Jericho environments with different observation and action space—ScaleZero consistently matches and often surpasses the performance of specific single-task expert agents.
-Finally, to resolve the challenge of computational inefficiency, we introduce Dynamic Parameter Scaling (DPS) from the overall learning process perspective, a novel online mechanism that couples model capacity to learning progress.
-DPS adaptively curates the set of active tasks based on real-time return feedback and orchestrates a phased expansion of model capacity by injecting lightweight LoRA adapters.
-By strategically freezing previously trained parameters, DPS creates a curriculum of model complexity that directs computational resources where they are most needed.
-Our experiments demonstrate that when augmented with our dynamic parameter scaling strategy, our method achieves performance nearly on par with single-task agents while reducing total environment interactions by around 20%, offering a superior trade-off between final performance and computational budget.
-Our main contributions are summarized as follows: (1) We provide the quantitative diagnosis of plasticity collapse in unified world models within heterogeneous MTRL, establishing a concrete link between performance degradation and internal network dynamics.
-(2) We conduct a systematic architectural exploration that yields ScaleZero, a unified world model that demonstrates exceptional performance and generalization across distinct tasks from Atari (diverse visual challenges), DMControl (state-based continuous control), and Jericho (text-based compositional reasoning) benchmark.
-(3) We propose Dynamic Parameter Scaling (DPS), an adaptive training strategy that dynamically allocates model capacity and computational resources, reducing total environment interactions by around 20%.
+This culminates in a progressive loss of network plasticity (Dohare et al., 2024; Todorov et al., 2025)the fundamental ability of a model to adjust its parameters to learn from new data—imposing a hard ceiling on the model’s overall learning capacity and leading to performance collapse on challenging tasks.
+(2) Static resource allocation: Conventional architectures employ a uniform, one-size-fits-all forward pass, applying the same learning resources (e.g., data collection and model updates) to every task irrespective of its intrinsic difficulty.
+This static strategy results in profound computational inefficiency, as resources are squandered on converged or show diminishing returns, instead of being directed toward those that require further learning.
+To dissect these issues, we begin with a quantitative diagnosis using UniZero (Pu et al., 2024), a contemporary unified world model, as a representative testbed.
+We train a single model via online reinforcement learning across eight canonical Atari games and observe a stark failure pattern, as shown in Figure 1.
+While the model rapidly masters simple tasks like Pong, it exhibits initial progress followed by a catastrophic performance collapse on complex games with disparate dynamics and visual styles, such as Seaquest, whose complex exploration demands starkly contrast with the reactive gameplay of tasks like Pong.
+We provide quantitative evidence linking this failure to specific internal model dynamics (Lyle et al., 2022): an uncontrolled inflation of the latent state norm and a corresponding spike in the dormant neuron ratio (Sokar et al., 2023) within the transformer backbone, signaling a collapse in model plasticity, rendering the network unable to adapt to new information from challenging tasks.
+Based on this diagnosis, we tackle these challenges from two complementary perspectives.
+First, as an internal, architectural solution, we undertake a principled exploration of the design space across the core dimensions input representation, model architecture, and optimization strategy.
+This systematic investigation, which evaluates five key axes—task conditioning, encoder architectures (ResNet (He et al., 2016) vs.
+ViT (Dosovitskiy et al., 2020)), latent normalization schemes, backbone design, and multitask optimization strategies (He et al., 2020)—culminates in a new, powerful model we term ScaleZero.
+Among the evaluated design choices, the integration of a sparse Mixture-of-Experts (Shazeer et al., 2017) backbone yields the most significant gains by fundamentally addressing plasticity collapse.
+To provide a deeper understanding of this key architectural choice, we present a dedicated empirical and theoretical analysis in Section 4.3, explaining MoE’s effectiveness in multitask planning.
+Validated on a comprehensive suite of online RL benchmarks spanning diverse modalities—including 26 visually-complex Atari games (Bellemare et al., 2013), 18 state-based DeepMind Control (DMC) Suite tasks (Tunyasuvunakool et al., 2020), and 4 text-based Jericho environments (Hausknecht et al., 2020)ScaleZero robustly matches and often surpasses the performance of single-task expert agents.
+Second, as an external, procedural strategy to optimize the overall learning process, we introduce Dynamic Parameter Scaling (DPS), a novel online mechanism that couples model capacity to learning progress.
+DPS adaptively curates the set of active tasks based on real-time return feedback and orchestrates a phased expansion of model capacity by injecting lightweight LoRA adapters (Hu et al., 2022).
+By strategically freezing previously trained parameters, DPS creates a curriculum of model capacity that directs computational resources where they are most needed.
+Our experiments demonstrate that when augmented with this strategy, our method achieves performance nearly on par with single-task agents while reducing total environment interactions by around 28.5% on the DMC benchmark, offering a superior trade-off between final performance and computational budget.
+Our main contributions are summarized as follows: (1) We provide the quantitative diagnosis of plasticity collapse in unified world models within heterogeneous MTRL, establishing a concrete link between performance degradation and internal learning dynamics.
+(2) We conduct a systematic architectural exploration that yields ScaleZero, a unified world model that demonstrates exceptional performance and generalization across distinct tasks from three distinct benchmarks.
+(3) We propose Dynamic Parameter Scaling, an adaptive training strategy that dynamically allocates model capacity and computational resources, reducing total environment interactions by around 28.5%.
 2 Background
-2.1 Unified World Models in MCTS-Based Reinforcement Learning
-Model-based reinforcement learning, particularly methods employing MCTS for planning within a learned latent space, has set the standard in complex sequential decision-making domains (Schrittwieser et al., 2019; Antonoglou et al., 2021; Ye et al., 2021).
-These methods learn a world model composed of three core components: (i) a representation model, $z_{t}=h_{\theta}(o_{t-H+1:t},a_{t-H+1:t-1})$, which encodes an observation history into a latent state, where $H$ is the training sequence length; (ii) a dynamics model, $(\hat{z}_{t+1},\hat{r}_{t})=g_{\theta}(z_{t},a_{t})$, predicting the subsequent latent state and reward; and (iii) a prediction model, $(p_{t},v_{t})=f_{\theta}(z_{t})$, which estimates a policy and value to guide the MCTS planner.
-Recent architectures like UniZero (Pu et al., 2024) push this paradigm further by unifying these three components into a single Transformer-based architecture.
-This monolithic design allows for end-to-end optimization via a composite loss function that aligns the model’s predictions with MCTS-derived targets for policy ($\pi_{t}$) and value ($\hat{v}_{t}$), alongside objectives for reward and state reconstruction:
-$\mathcal{L}_{\text{UniZero}}=\sum_{t=0}^{H-1}\left(\mathcal{L}_{\text{value}}(v_{t},\hat{v}_{t})+\mathcal{L}_{\text{policy}}(p_{t},\pi_{t})+\mathcal{L}_{\text{reward}}(\hat{r}_{t},r_{t})+\mathcal{L}_{\text{dynamics}}(\hat{z}_{t+1},\operatorname{sg}(z_{t+1}))\right),$ (1)
-where $\operatorname{sg}(\cdot)$ denotes the stop-gradient operator.
-While parameter-efficient, this unified training scheme creates a critical vulnerability in heterogeneous settings: the shared parameters are subject to a single, aggregated gradient signal, which can be dominated by easier tasks (Ma et al., 2023), thereby impeding learning on more complex ones and leading to plasticity collapse.
-2.2 Gradient Conflict in Multi-Task Reinforcement Learning
-In MTRL, we consider a distribution of $K$ distinct tasks, $\mathcal{T}=\{\tau_{1},\dots,\tau_{K}\}$.
-The objective is to learn a single set of shared parameters $\theta$ for a task-conditioned policy $\pi_{\theta}(a|s,k)$ that maximizes expected returns across all tasks, where $k\in\{1,\dots,K\}$ is the task identifier.
-Given a batch of data from all tasks, the total loss is the sum of the individual task losses: $\mathcal{L}_{\text{MTRL}}(\theta)=\sum_{k=1}^{K}\mathcal{L}^{k}_{\text{UniZero}},$ where $\mathcal{L}^{k}_{\text{UniZero}}$ is the loss computed on data from task $k$.
-The learning dynamics are driven by the total gradient $G_{\text{total}}=\nabla_{\theta}\mathcal{L}_{\text{MTRL}}$, which is the sum of per-task gradients: $G_{\text{total}}=\sum_{k=1}^{K}g_{k}$, where $g_{k}=\nabla_{\theta}\mathcal{L}^{k}_{\text{UniZero}}$.
-The core problem, known as gradient conflict, arises when these per-task gradients are misaligned.
-For any two tasks $i$ and $j$, the conflict can be quantified by the cosine similarity of their gradients: $\cos(g_{i},g_{j})=\frac{g_{i}\cdot g_{j}}{\|g_{i}\|\|g_{j}\|}$.
-A negative value indicates that an update improving performance on task $i$ will degrade performance on task $j$.
-In a monolithic architecture where all parameters $\theta$ are shared, frequent negative similarities lead to destructive interference (Sodhani et al., 2021).
-This forces the model into a suboptimal compromise, creating a representational bottleneck and impeding mastery of tasks with conflicting or complex objectives (Hansen et al., 2023).
-2.3 Sparse and Parameter-Efficient Architectures
-A significant challenge in developing unified models for heterogeneous environments is overcoming issues such as representational bottlenecks, plasticity collapse, and the rigidity imposed by static resource allocation.
-To address these challenges, the literature has explored several sparse and parameter-efficient architectures that operate at different perspectives.
-One prominent approach involves replacing the standard feed-forward network in a Transformer block (Vaswani et al., 2017) with a sparse Mixture-of-Experts architecture (Shazeer et al., 2017; Liu et al., 2024a).
-This model comprises $M$ expert networks and a trainable sparse gating function, $G(x)$, which selects a small top-k subset of experts for each input.
-The forward pass is defined as: $\mathrm{MoE}(x)=\sum_{i=1}^{M}G_{i}(x)\cdot\mathrm{Expert}_{i}(x).$ This conditional computation creates distinct representational pathways for different inputs and significantly improve model performance and efficiency with fewer resources.
-For parameter-efficient adaptation of large models, Low-Rank Adaptation (LoRA) (Hu et al., 2022) has become a widely adopted technique.
-LoRA adapts a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times m}$ by learning a low-rank update, $\Delta W=BA$, where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times m}$, with the rank $r\ll\min(d,m)$.
-The effective weight matrix then operates on an input $x$ as: $W(x)=\bigl{(}W_{0}+\alpha\,BA\bigr{)}x,$ where a learnable scalar $\alpha$ controls the magnitude of the update.
-By freezing the original weights $W_{0}$ and training only the compact low-rank factors $(A,B)$, LoRA facilitates efficient adaptation and isolates task-specific modifications.
-In this context, our work fuses these ideas from two perspectives.
-At the single learning iteration level, a gating function sparsely selects LoRA modules, treating them as experts.
-At the overall learning process level, we incrementally add new LoRA modules for new tasks—freezing prior ones—to efficiently scale model capacity.
+2.1 MCTS with Learned World Models
+Monte Carlo Tree Search is a powerful planning algorithm that has demonstrated remarkable success in domains with known rules (e.g.
+board games).
+Methods employing MCTS for planning within a learned latent space represent the state-of-the-art in complex sequential decision-making (Schrittwieser et al., 2019; Ye et al., 2021).
+These approaches learn a world model comprising three components: (i) a representation model $h_{\theta}$ that encodes observation-action history into a latent state $z_{t}$; (ii) a dynamics model $g_{\theta}$ that predicts the next latent state and reward $(\hat{z}_{t+1},\hat{r}_{t})=g_{\theta}(z_{t},a_{t})$; and (iii) a prediction model $f_{\theta}$ that outputs a policy and value $(p_{t},v_{t})$ to guide the MCTS planner.
+Recent architectures, exemplified by UniZero (Pu et al., 2024), unify these components into a single, monolithic Transformer.
+This design enables end-to-end optimization via a composite loss that aligns the model’s predictions with targets derived from the MCTS planner: a policy target ($\pi_{t}$) based on the root node’s visit counts, and a bootstrapped TD value target ($\hat{v}^{targ}_{t}$)  (Schrittwieser et al., 2019).
+These are combined with objectives for predicting the reward ($r_{t}$) and the next latent state ($z_{t+1}$):
+$\mathcal{L}_{\text{Unified}}=\sum_{t=0}^{H-1}\left(\mathcal{L}_{\text{value}}(v_{t},\hat{v}^{targ}_{t})+\mathcal{L}_{\text{policy}}(p_{t},\pi_{t})+\mathcal{L}_{\text{reward}}(\hat{r}_{t},r_{t})+\mathcal{L}_{\text{dynamics}}(\hat{z}_{t+1},\operatorname{sg}(z_{t+1}))\right),$ (1)
+where $H$ is the context length, $\operatorname{sg}(\cdot)$ means stop-gradient.
+While sample-efficient and hyperparameter-insensitive for single-task training, this unified scheme exposes a critical vulnerability: in multitask settings, the shared transformer backbone is updated by an aggregated gradient.
+This shared update mechanism becomes the nexus of interference, impeding a single model’s ability to learn multiple complex task and leading directly to the plasticity collapse we identified in our diagnosis.
+2.2 Gradient Conflict in Unified Architectures
+The challenge of MTRL is to train a single, task-conditioned policy $\pi_{\theta}(a|s,k)$ on a distribution of $K$ tasks, $\{\tau_{1},\dots,\tau_{K}\}$.
+Even within a single task, the composite losses of a world model (e.g., for dynamics, reward, and policy) can conflict, creating a challenging multi-objective optimization problem (Ma et al., 2023).
+In MTRL, this challenge is compounded by inter-task gradient conflicts.
+In a unified architecture with shared parameters $\theta$, the total loss is the sum of individual task losses, $\mathcal{L}_{\text{MTRL}}(\theta)=\sum_{k=1}^{K}\mathcal{L}^{k}$, where $\mathcal{L}^{k}$ is the loss from Eq.
+1 for task $k$.
+Consequently, the learning dynamics are driven by the sum of per-task gradients: $G_{\text{total}}=\sum_{k=1}^{K}g_{k}$, where $g_{k}=\nabla_{\theta}\mathcal{L}^{k}$.
+The core problem, known as gradient conflict (Liu et al., 2021), arises when these per-task gradients are misaligned, a condition quantified by their cosine similarity: $\cos(g_{i},g_{j})<0$.
+A negative similarity indicates that an update improving performance on task $i$ actively degrades performance on task $j$.
+In a monolithic architecture where all core parameters are shared, frequent gradient conflicts lead to destructive interference (Sodhani et al., 2021).
+This dynamic forces the model into a suboptimal compromise, creating the representational bottlenecks and catastrophic performance drops characteristic of plasticity collapse (Dohare et al., 2024).
+2.3 Architectural Paradigms for Mitigating Interference
+To overcome the limitations of monolithic designs, architectural solutions have been proposed to enable parameter specialization and mitigate interference.
+Two prominent paradigms are particularly relevant.
+Sparse Mixture-of-Experts replaces the dense feed-forward network in a Transformer block with a set of $N$ parallel "expert" networks (Cai et al., 2025).
+For each input, a trainable gating function $G(x)$ sparsely selects a small subset of experts (e.g., top-k) to process the token: $\mathrm{MoE}(x)=\sum_{i=1}^{N}G_{i}(x)\cdot\mathrm{Expert}_{i}(x).$ This conditional computation creates specialized pathways within the model, allowing different inputs (e.g., from different tasks) to be processed by distinct subsets of parameters, thereby reducing interference and increasing model capacity without a proportional rise in computational cost.
+Recent designs also explore hybrid approaches, such as including a shared expert alongside specialized ones, to balance generalization and specialization (Dai et al., 2024).
+Low-Rank Adaptation offers a parameter-efficient method for adapting large pre-trained models (Hu et al., 2022).
+Instead of fine-tuning the entire weight matrix $W_{0}$, LoRA freezes $\theta_{B}$ and injects a trainable, low-rank "update" matrix, $\Delta\theta=BA$.
+The modified forward pass becomes $(\theta_{B}+\alpha BA)x$.
+By training only the small low-rank factors $(A,B)$, LoRA can efficiently learn task-specific modifications while preserving the general knowledge in the pre-trained weights.
+This approach has proven effective for creating extensive multitask capabilities on top of a single base model, where different LoRA modules can be composed to handle diverse tasks (Huang et al., 2023).
 3 Method
-This section systematically explores the core challenges that unified world models face in heterogeneous multi-task learning.
-We begin in Section 3.1 by quantitatively diagnosing the phenomenon of plasticity collapse in shared-parameter models under a multi-task paradigm, uncovering its underlying mechanisms.
-Subsequently, in Section 3.2, we systematically dissect the design space of the UniZero model.
-Through a series of ablation studies, we converge on a new architecture, termed ScaleZero, which effectively mitigates plasticity degradation.
-Finally, to address the inefficiency of static resource allocation, we introduce Dynamic Parameter Scaling (DPS) in Section 3.3.
-This strategy adaptively expands the model’s capacity in response to task-specific learning progress, thereby achieving a superior trade-off between computational efficiency and final performance.
-3.1 Diagnosing Plasticity Collapse in Multi-Task Learning
-Figure 1: Diagnosing plasticity collapse in the baseline UniZero model on a multi-task Atari benchmark.
+We begin in Sec.
+3.1 by quantitatively diagnosing plasticity collapse in multitask training.
+To address this, Sec.
+3.2 presents a systematic design space exploration across five axes—which culminates in our proposed model, ScaleZero.
+Finally, to resolve the static allocation issue, we introduce Dynamic Parameter Scaling in Sect.
+3.3, a LoRA-based strategy that adaptively allocates model capacity.
+3.1 Diagnosing Plasticity Collapse in Multitask Planning
+Figure 1: Plasticity collapse in the baseline (UniZero) on a multitask Atari benchmark.
 While simple tasks like Pong and Hero show stable learning, complex tasks such as Seaquest and ChopperCommand suffer a catastrophic performance collapse in later training (Top).
-This performance failure is precisely correlated with a sharp spike in the dormant neuron ratio of the Transformer backbone (Bottom Left) and an uncontrolled inflation of the latent state norm (Bottom Right), empirically validating the link between external performance and internal network dynamics.
-Metrics.
-In MTRL settings, shared-parameter models are susceptible to a freezing phenomenon, where parts of the network become saturated and lose their ability to adapt to new information.
-This plasticity collapse stems from the inherent heterogeneity across tasks in terms of their features, difficulty, and informational content, ultimately constraining the model’s overall learning potential.
-To quantitatively measure this degradation in network adaptability, we utilize two key metrics: (1) Dormant Neuron Ratio (Sokar et al., 2023): This metric measures the proportion of neurons in a given layer whose activation magnitude falls below a predefined threshold $\epsilon$ (set to 0.025 in this work) over a time series.
-A high dormant ratio indicates that a significant portion of the network’s pathways are inactive and no longer contribute to information processing.
-$\text{DormantRatio}(l)=\frac{1}{N_{l}}\sum_{i=1}^{N_{l}}\mathbf{1}\!\left(|h_{i}^{l}|\leq\epsilon\right),$ where $h_{i}^{l}$ is the activation of the $i$-th neuron in layer $l$, and $N_{l}$ is the total number of neurons in that layer.
-(2) Latent State Norm: This metric computes the average L2 norm of the latent state vectors $z_{t}$ produced by the encoder over a time series.
-An abnormally inflated latent norm is often correlated with training instability and slower learning.
-As our model processes sequential data in mini-batches, we average this norm across both the training sequence length $T$ and the mini-batch size $B$.
-Analysis.
-Our experiments on a multi-task suite of 8 Atari games (see Appendix A) reveal a clear failure pattern: while simple tasks (e.g., Pong) converge stably, complex tasks (e.g., Seaquest) suffer a catastrophic performance collapse late in training (Figure 1).
-This collapse precisely corresponds to a sharp increase in the dormant neuron ratio and an uncontrolled inflation of the latent state norm as shown in bottom of Figure 1.
-In addition, we have added the trend of the feature effective rank Dohare et al.
-(2024) and training return in Appendix Figure 6, which also shows a marked decline in the performance degradation range.
-Thus, we attribute this plasticity loss to two intertwined factors: (1) Gradient Competition, where simpler, faster-converging tasks generate dominant gradients that suppress signals from more complex tasks, causing critical neural pathways to "go dormant"; and (2) Representation Interference, where a single, shared model struggles to encode diverse task dynamics, creating a representational bottleneck that constricts the solution space.
-Motivated by this diagnosis, we propose to mitigate these issues by introducing task-differentiated mechanisms from two distinct perspectives.
-First, from the perspective of a single learning iteration, we investigate modifications that resolve interference within each forward and backward pass, leading to the design of ScaleZero (Section 3.2).
-Second, from the perspective of the overall learning process, we introduce a novel training strategy that dynamically allocates model capacity over time, we named as Dynamic Parameter Scaling strategy (Section 3.3).
-3.2 Exploring the Design Space of UniZero
+This failure is precisely correlated with a sharp spike in the dormant neuron ratio of the transformer (Bottom Left) and an uncontrolled inflation of the latent state norm (Bottom Right), empirically validating the link between external performance and internal learning dynamics.
+To empirically substantiate the challenges of representational bottlenecks and plasticity collapse, we conduct a quantitative analysis of a unified world model’s internal dynamics during multitask training.
+We track two key metrics indicative of network plasticity degradation: (1) $\textit{DormantRatio}(l)=\frac{1}{N_{l}}\sum_{i=1}^{N_{l}}\mathbf{1}!\left(|h_{i}^{l}|\leq\epsilon\right):$ This metric quantifies the proportion of neurons whose activation magnitude $|h_{i}^{l}|$ falls below a threshold $\epsilon$.
+A high ratio signifies a loss of active pathways and reduced network plasticity.
+(2) Latent State Norm: We compute the average L2 norm of the latent state ($|z_{t}|_{2}$).
+An uncontrolled inflation of this norm is a well-known indicator of training instability (Team et al., 2025).
+Diagnostic experiments on UniZero, conducted across a multitask Atari8 suite (Appendix B), reveal a critical failure mode in complex learning scenarios.
+While simpler tasks like Pong achieve stable convergence, more complex ones such as Seaquest exhibit initial learning followed by a catastrophic performance collapse (Figure 1).
+This collapse coincides with a sharp increase in the dormant neuron ratio and an inflated latent state norm.
+The representational nature of this failure is corroborated by a marked decline in the feature effective rank (Dohare et al., 2024), indicating a degradation of the learned feature space (Figure 5).
+We attribute this failure mode to two intertwined factors: (1) Gradient Competition, where gradients from simpler tasks overwhelm those from complex ones, and (2) Representation Interference, where the shared network fails to maintain diverse features, leading to a representational bottleneck.
+Based on this diagnosis, our method tackles plasticity collapse from two complementary perspectives: (1) an internal, architectural solution, ScaleZero (Section 3.2), designed to mitigate interference within a single learning iteration.
+(2) an external, procedural strategy, Dynamic Parameter Scaling (Section 3.3), which optimizes resource allocation across the entire learning process.
+3.2 Architecture Design of ScaleZero via Principled Exploration
 (a) Design Space of UniZero for Multitask learning
 (b) Dynamic Parameter Scaling
-Figure 2: (a) A systematic exploration of the UniZero design space across five axes: task conditioning, encoder architecture, latent normalization, backbone design, and multitask optimization.
+Figure 2: (a) A systematic exploration of the UniZero design space across five axes: task conditioning, encoder architecture, latent normalization, backbone design, and optimization.
 This investigation informs the design of our proposed ScaleZero model.
-(b) A conceptual diagram of Dynamic Parameter Scaling (DPS), our adaptive training strategy.
-DPS progressively expands model capacity by injecting lightweight LoRA adapters in stages, triggered by learning progress, to direct computational resources toward unsolved tasks while preserving existing knowledge.
-Figure 3: Performance impact of architectural modifications on the Atari8 multi-task benchmark.
-This ablation study across the UniZero design space reveals that replacing the standard Transformer backbone with a Mixture-of-Experts architecture yields the most significant and consistent performance gains.
+(b) A conceptual diagram of Dynamic Parameter Scaling (DPS).
+DPS progressively expands model capacity by injecting LoRA adapters in stages, triggered by learning progress.
+This creates a curriculum of model, directing resources toward unsolved tasks while preserving existing knowledge.
+Figure 3: Performance impact of architectural modifications on the Atari8 multitask benchmark.
+This ablation across the UniZero design space reveals that replacing the Transformer backbone with a Mixture-of-Experts architecture yields the most significant and consistent performance gains.
 In contrast, other interventions, with the partial exception of SimNorm, provide marginal or inconsistent benefits.
 These results underscore the centrality of the MoE’s conditional computation in overcoming the limitations of a shared, dense backbone.
-Figure 4: Detailed analysis correlating performance with plasticity metrics across model variants in Figure 3.
-On two representative hard tasks (ChopperCommand and Seaquest), the performance collapse of baseline models directly correlates with catastrophic plasticity loss—evidenced by a sharp rise in the dormant neuron ratio and latent norm.
-In stark contrast, the MoE-based model sustains high returns by keeping the dormant ratio near-zero and the latent norm stable.
-This confirms that the MoE’s superior performance stems directly from its success in mitigating the diagnosed plasticity issues.
-Analyses for the remaining six environments, which show consistent trends, are deferred in Figure 8.
-Based on the diagnosis above, we conducted a systematic decomposition and evaluation of the UniZero model’s design space to identify architectural components that can effectively counteract plasticity loss.
-We explored five key dimensions, as illustrated in Figure 2 (left): task information, encoder, latent normalization, backbone, and optimization strategy.
-The experimental details is showed in Appendix B.
-Figure 3 provides a overall comparison of the performance contributions from each design dimension.
-A key observation is that the MoE backbone yields the most significant performance gains.
-In contrast, other modifications, with the partial exception of SimNorm, fail to deliver stable performance improvements.
-To investigate the mechanisms behind these results, Figure 4 illustrates the relationship between task returns and plasticity-related metrics (i.e., transformer dormant ratio and latent norm) for two representative tasks, ChopperCommand and Seaquest.
-Analyses for the remaining six environments are provided in Appendix B.
-The key findings from this exploration are detailed below:
+Motivated by the above diagnosis of plasticity collapse, we conduct a principled exploration of the UniZero design space to forge a robust solution on Atari8 multitask benchmark (see Appendix B for experimental details).
+We systematically evaluate five key design axes (Figure 2(a)) drawn from three core dimensions (input representation, model architecture, and optimization strategy): task conditioning, encoder architecture, latent normalization, backbone design, and optimization strategy.
+Specific implementation details are provided in Appendix A.3.
+The comprehensive results, summarized in Figure 3, reveal that the integration of a MoE backbone provides a uniquely significant and robust performance uplift, directly counteracting the failure modes identified in our initial diagnosis.
+Below, we detail the findings of each axis.
 Explicit Task Conditioning.
-We experimented with concatenating a learnable task embedding indexed by task ID to the state representation to provide an explicit task identity.
-While this accelerated initial convergence on simpler tasks, its long-term benefit for complex tasks was marginal, and it failed to significantly improve internal plasticity metrics (Figure 4).
-This suggests that shallow task identifiers are insufficient to resolve deep-seated representational conflicts in online RL setting.
+A shared model processing raw observations may face ambiguity in discerning the underlying task, especially when initial states are similar across environments (Rakelly et al., 2019; Sodhani et al., 2021).
+To test if providing an explicit task identity $k$ could resolve ambiguity, we concatenate a learnable task embedding $e_{k}$ to the latent state $z$.
+While this approach accelerates initial convergence on simpler tasks, it offers marginal long-term benefits for complex ones and fails to mitigate plasticity collapse, as evidenced by minimal improvement in metrics like the dormant neuron ratio (Figure 6).
+This indicates that such input-level task conditioning are insufficient to resolve deep-seated representational conflicts in online multitask RL.
 Encoder Architecture (ResNet vs.
 ViT).
-We compared a standard ResNet encoder against a ViT to assess the impact of encoder architecture.
-In our experiments on the Atari8 benchmark, the choice of encoder did not yield a significant performance advantage.
-We hypothesize that this is due to the limited number of tasks and the relatively low visual heterogeneity across the environments in this specific benchmark.
-However, we anticipate that the ViT’s superior capacity and scalability will become more critical in larger-scale, more heterogeneous multi-task settings.
+We contrast a standard ResNet-like encoder in  Pu et al.
+(2024) with a Vision Transformer to assess the encoder’s impact on visual representation quality.
+On the Atari8 benchmark, the choice of encoder does not yield a significant performance difference.
+We attribute this to the limited visual heterogeneity in this specific task suite, but posit that the scalability and powerful feature extraction of ViT are critical for broader, more diverse multitask domains.
 Latent Normalization (LayerNorm vs.
 SimNorm).
-To suppress the inflation of the latent norm, we compared a standard LayerNorm (Ba et al., 2016) against a SimNorm (Hansen et al., 2023) to assess the impact of latent normalization.
-SimNorm (Hansen et al., 2023) normalize the latent representation by projecting $z$ into $L$ fixed-dimensional simplices using a softmax operation, which clamps the L2 norm of latent vectors within a predefined threshold.
-Although SimNorm effectively stabilized training and prevented performance collapse, its hard constraint appeared to curtail the expressive power of the representations, highlighting a trade-off between stability and representational capacity.
+To directly combat the observed latent norm inflation, we evaluate SimNorm (Hansen et al., 2023) against the standard LayerNorm (Ba et al., 2016).
+SimNorm, which projects latents onto a simplex to enforce a norm constraint, effectively stabilized training and prevented performance collapse, consistent with its reported success in less heterogeneous settings (Hansen et al., 2023).
+However, in our highly diverse multitask setup, this hard constraint appears to curtail representational expressiveness, leading to suboptimal final performance.
+This highlights a critical trade-off between stability and capacity for online heterogeneous RL training.
 Mixture-of-Experts Backbone.
-To fundamentally address plasticity loss in the Transformer backbone, we replaced its standard feed-forward networks (FFNs) with MoE layers.
-In an MoE, a gating network sparsely routes each input token to a subset of "expert" sub-networks.
-This conditional computation mechanism decouples the dense computational pathways into multiple, parallel, specialized ones.
-This design: (1) reduces parameter collision, as different tasks can leverage different combinations of experts, mitigating representation interference; and (2) mitigates gradient conflict, as gradient updates for different tasks are largely directed to the distinct sets of activated experts.
-As shown in Figures 4 and 4, the MoE backbone delivered the most significant performance improvements on complex tasks like Seaquest and maintained a relatively low dormant ratio throughout training, effectively controlling the latent norm.
-Multi-Task Gradient Correction.
-We implemented a dynamic gradient re-weighting scheme inspired by MoCo (Fernando et al., 2023) to directly address gradient interference.
-Despite yielding gains on some tasks, its significant computational overhead and inconsistent efficacy across the heterogeneous task suite led us to explore a more efficient gradient correction in future.
-ScaleZero.
-Based on this exploration, we integrated the most impactful components to form the ScaleZero architecture.
-Its core design combines: (1) a ViT Encoder for powerful feature extraction; (2) an MoE Backbone to fundamentally address plasticity collapse through sparse, conditional computation; and (3) Standard Layer Normalization on the latent states to ensure stability without sacrificing expressive power.
-As validated in Section 4, ScaleZero provides a novel and effective blueprint for building generalist world models.
-3.3 Dynamic Parameter Scaling for Balanced Multi-Task Learning
-In heterogeneous MTRL, static, one-size-fits-all model architectures are suboptimal.
-They tend to waste computational resources on tasks that have already been mastered and often lack the necessary plasticity to effectively tackle the remaining, more difficult tasks.
-This can lead to inefficient training and risks negative transfer.
-To address this, we propose Dynamic Parameter Scaling (DPS), a strategy illustrated in Figure 2 (right).
-The core insight of DPS is that a model’s effective capacity should dynamically align with the current set of unsolved tasks.
-This approach instantiates a "curriculum of model complexity," adaptively expanding the model with new parametric resources directed precisely at the most challenging tasks still in progress.
-Formalism and Mechanism.
-Let the set of tasks be $\mathcal{T}=\{\tau_{i}\}_{i=1}^{N}$ and the parameters of the shared backbone network be $\theta_{B}$.
-We define a task $\tau_{i}$ as "solved" when its performance metric, $\text{Metric}(\tau_{i})$, meets or exceeds a predefined threshold $\varepsilon_{i}$.
-At any time $t$, the set of all unsolved tasks is denoted as the active set $\mathcal{U}_{t}\subseteq\mathcal{T}$.
-To conserve computational resources, both data collection and model training are exclusively performed on the tasks within this active set.
-Once a task is considered solved, it is removed from $\mathcal{U}_{t}$, and we cease all related training activities for it.
-The training protocol is divided into $S+1$ stages ($s=0,1,\dots,S$).
-It begins with Stage 0 (Warm-up), an initial period of $T_{0}$ iterations where only the backbone parameters $\theta_{B}$ are trained.
-Each subsequent Stage $s\geq 1$ (Expansion) introduces an independent Low-Rank Adaptation (LoRA) module, $\Delta\theta_{s}$.
-This module consists of a trainable matrix pair $\Delta\theta_{s}=B_{s}A_{s}$ (where $B_{s}\in\mathbb{R}^{d\times r},A_{s}\in\mathbb{R}^{r\times k}$ with rank $r\ll d,k$) and a learnable scalar scaling factor $\alpha_{s}\in\mathbb{R}$.
+The most impactful modification is replacing the dense feed-forward networks with sparse MoE layers in UniZero’s transformer.
+This conditional computation paradigm directly addresses the core challenges of multitask planning by mitigating the intertwined issues of representational interference and gradient conflict, as it routes task-specific computations and their corresponding gradient updates through distinct, specialized expert sub-networks.
+As validated in Figures 3 and 6, the MoE backbone yields substantial performance gains.
+We attribute this success to its inherent ability to mitigate gradient conflict, a mechanism we analyze in depth in Section 4.3.
+There, we provide both empirical validation and a theoretical analysis demonstrating that the upper bound on gradient conflict for an MoE layer is strictly lower than that of its dense counterpart.
+Multitask Gradient Correction.
+We also explored a dynamic gradient re-weighting scheme inspired by MoCo (Fernando et al., 2023) to directly mitigate gradient interference between tasks.
+While this approach showed promise on some tasks, it introduced significant computational overhead by increasing per-step training time by approximately 40%, and demonstrated inconsistent efficacy across our heterogeneous task suite.
+Given this unfavorable performance-cost trade-off, we defer the investigation of more efficient gradient correction methods to future work.
+The ScaleZero Architecture.
+The result of our systematic investigation is ScaleZero, a novel architecture designed as a highly efficient and scalable world model.
+It is defined by the integration of three core components: (1) a ViT-like Encoder for powerful and scalable feature extraction; (2) an MoE Backbone that uses sparse, conditional computation to increase model capacity while mitigating task interference; and (3) an explicit Layer Normalization applied to the encoded latent state to ensure a robust balance between training stability and representational expressiveness.
+This core is complemented by a modular design that decouples task-specific encoders and heads, enabling flexible handling of diverse input and output modalities.
+As validated in Section 4, ScaleZero provides an effective blueprint for building generalist world models.
+Key differences from UniZero are detailed in Appendix A.2.
+3.3 Dynamic Parameter Scaling for Efficient Multitask Learning
+To address the second challenge in multitask planning—static resource allocation—we propose Dynamic Parameter Scaling (DPS).
+This strategy, illustrated in Figure 2(b), abandons the one-size-fits-all paradigm by dynamically coupling model capacity with learning progress.
+DPS operates on two synergistic principles: (1) adaptive task curation, which focuses computational effort exclusively on unsolved tasks, and (2) progressive capacity expansion, which injects new parameters only when needed.
+This combination creates a "curriculum of model complexity," directing resources precisely where they are most required.
+DPS employs dynamic, staged training to manage tasks and parameters.
+  • Task Curation: Let $\mathcal{T}=\{\tau_{i}\}_{i=1}^{N}$ be the full set of tasks.
+A task $\tau_{i}$ is deemed "solved" once its performance metric, $\text{Metric}(\tau_{i})$, surpasses a threshold $\varepsilon_{i}$.
+At any time $t$, we maintain an active set of unsolved tasks, $\mathcal{U}_{t}\subseteq\mathcal{T}$.
+To maximize efficiency, both data collection and gradient updates are performed only for tasks within $\mathcal{U}_{t}$.
+Once solved, a task is removed from this set, ceasing all computational overhead associated with it.
+  • Staged Capacity Expansion: The training proceeds in $S+1$ stages.
+Stage 0 (Warm-up) trains the base model $\theta_{B}$ for an initial $T_{0}$ iterations.
+Triggered by learning progress, each subsequent Stage $s\geq 1$ (Expansion) injects a new, independent LoRA module, $\Delta\theta_{s}$.
+This module adapts a target weight matrix of dimensions $d\times k$ via a low-rank decomposition $\Delta\theta_{s}=B_{s}A_{s}$, where $B_{s}\in\mathbb{R}^{d\times r}$, $A_{s}\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$.
+The module’s update is scaled by a learnable scalar $\alpha_{s}$.
 Adaptive Stage Transition Triggers.
-The transition from stage s1s-1 to stage $s$ is governed by one of two conditions, balancing learning progress against a fixed training budget: (1) Progress-based Trigger: A new stage is initiated when the number of newly solved tasks since the beginning of the current stage reaches a predefined quota $Q_{s}$.
-(2) Budget-based Trigger: To prevent training from stalling on exceptionally difficult tasks, a budget-based trigger forces a transition if the number of iterations in the current stage exceeds a pre-allocated limit, $\lceil(T_{\max}-T_{0})/S\rceil$.
-Parameter Composition and Optimization.
+The transition from stage s1s-1 to $s$ is initiated when either of two conditions is met, balancing progress against a fixed budget: (1) Progress-based Trigger: A new stage begins if the number of newly solved tasks in the current stage reaches a quota $Q_{s}$.
+(2) Budget-based Trigger: To prevent stagnation on particularly arduous tasks, a transition is forced if the iteration count in the current stage exceeds a pre-allocated limit, $\lceil(T_{\max}-T_{0})/S\rceil$.
+Optimization and Parameter Isolation.
 In any given stage $s$, the effective computation for a weight matrix $W_{0}\in\theta_{B}$ is expressed as $W^{(s)}(x)=W_{0}x+\sum_{j=1}^{s}\alpha_{j}(B_{j}A_{j})x$.
-Crucially, upon entering a new stage $s$, we freeze both the backbone parameters $\theta_{B}$ and all previously introduced LoRA modules $\{\Delta\theta_{j}\}_{j=1}^{s-1}$.
-The optimization gradient is then applied only to the newly added module $\Delta\theta_{s}$ and the full set of scaling factors $\{\alpha_{j}\}_{j=1}^{s}$.
-This strategy offers a compelling dual advantage.
-First, it is computationally efficient by adaptively allocating resources according to task difficulty; new parameters are only introduced as needed for unsolved tasks, avoiding redundant computation on those already mastered.
-Second, the dynamic injection of new LoRA modules provides targeted plasticity for more challenging tasks without introducing negative transfer.
-By isolating updates within new modules while preserving the stable knowledge in the frozen backbone, our approach ensures that learning to solve difficult tasks does not disrupt previously acquired capabilities.
-In summary, DPS enables an intelligent allocation of model capacity, achieving a superior trade-off between efficiency and final performance.
-Pseudocode and implementation details are provided in Appendix A.3.
+Crucially, upon entering stage $s$, all prior parameters—the backbone $\theta_{B}$ and adapters $\{\Delta\theta_{j}\}_{j=1}^{s-1}$—are frozen.
+Optimization is thereby focused on the newly added LoRA module $\Delta\theta_{s}$ and the full set of scaling factors $\{\alpha_{j}\}_{j=1}^{s}$.
+This strategy provides two decisive advantages.
+First, it implements a resource-efficient training curriculum, as new parameters are only introduced when required by the remaining unsolved tasks.
+Second, by isolating new learning within dedicated adapters, DPS provides targeted plasticity for difficult tasks while preventing catastrophic forgetting or negative transfer to previously mastered skills, whose knowledge is preserved in the frozen backbone.
+DPS thus achieves a superior trade-off between computational efficiency and final performance.
+(See Appendix A.4 for pseudocode).
 4 Experiments and Analysis
-To comprehensively evaluate the effectiveness of our proposed ScaleZero model and the dynamic parameter scaling strategy, we designed a series of experiments.
-Section 4.1 aims to validate the performance of ScaleZero on several challenging heterogeneous multitask benchmarks (Atari 100k benchmark (Bellemare et al., 2013a), DeepMind Control Suite (Tunyasuvunakool et al., 2020), and Jericho(Hausknecht et al., 2020)), comparing it against specialized models trained independently for each task.
-Section 4.2 then focuses on quantitatively analyzing the specific gains in sample and computational efficiency afforded by our dynamic parameter scaling strategy (the combined method we called ScaleZero-DPS).
-All experiments are conducted in a purely online RL setting, without relying on any expert data, to test the model’s capabilities in a realistic learning scenario.
-Section 4.3 complements our benchmark results with a focused analysis of MoE, combining empirical evidence with theoretical justification to explain why MoE serves as an effective backbone for multitask learning.
+Our experiments are designed to rigorously validate the two central claims of this work.
+First, in Section 4.1, we demonstrate that our proposed ScaleZero effectively overcomes the representational bottlenecks and plasticity collapse issue.
+We test its ability to achieve state-of-the-art performance as a single, generalist agent across a suite of heterogeneous multi-task benchmarks spanning visual, state-based, and text-based domains (Atari (Bellemare et al., 2013), DMControl (Tunyasuvunakool et al., 2020), and Jericho (Hausknecht et al., 2020)).
+Second, in Section 4.2, we quantify the efficiency gains of our DPS strategy, showing it successfully addresses the problem of static resource allocation.
+Finally, in Section 4.3, we provide a dedicated empirical and theoretical analysis of the Mixture-of-Experts mechanism, substantiating its critical role as the architectural cornerstone of ScaleZero’s success.
+All experiments are conducted in a purely online RL setting, without reliance on expert data, to assess true learning capability.
 Implementation details are provided in Appendix A.
 4.1 Multitask Learning Benchmarks
-4.1.1 Performance Evaluation on Atari 100k benchmark
-Setup.
-We evaluate ScaleZero on the large-scale Atari 100k benchmark, which comprises 26 distinct games.
-This benchmark is renowned for its significant heterogeneity across tasks, featuring diverse observation spaces (pixel inputs), action spaces (discrete instruction sets), and widely varying game difficulties and mechanics.
-We compare the multitask (MT) ScaleZero model against a single-task (ST) UniZero baseline, which is individually trained and optimized for each game.
-The evaluation metrics are the mean and median of the Human-Normalized Score (HNS), a standard in the field, to provide a comprehensive assessment of performance across tasks of varying difficulty.
-Results.
-As shown in Table 1, by training a single, unified model, ScaleZero achieves a mean normalized score that surpasses the average performance of 26 independently trained single-task UniZero models.
-This indicates that ScaleZero not only successfully learns a diverse set of tasks but also achieves positive knowledge transfer across them.
-Notably, the performance gains are most pronounced on several notoriously challenging tasks, such as Seaquest.
-We hypothesize that this is because ScaleZero’s unified world model can learn general visual and dynamic priors from simpler tasks and effectively apply them to more difficult ones that demand complex planning and exploration strategies.
-Although the median score is slightly lower than the single-task baseline, which is primarily influenced by a few tasks with extremely hard exploration challenges, the superior mean score robustly demonstrates the overall advantage of our approach.
-Table 1: Comparison of Human-Normalized Scores on 26 Atari games for ScaleZero (MT) vs. UniZero (ST). Detailed curves are available in Appendix B Figure 9.
-Algorithm Normalized Mean Normalized Median
+4.1.1 Atari Multitask Benchmark
+Table 1: Performance comparison of our multi-task model, ScaleZero (MT), against the single-task UniZero (ST) baseline across discrete (Atari) and continuous (DMControl) control benchmarks.
+Algorithm Norm. Mean Norm. Median
 UniZero (ST) 0.38 0.21
 ScaleZero (MT) 0.39 0.16
-4.1.2 Performance Evaluation on the DeepMind Control Suite
-Setup.
-To test the generalization capability of ScaleZero in the continuous control domain, we conducted evaluations on the DMC benchmark, which includes 18 tasks based on the MuJoCo physics engine.
-These tasks feature vector state space and continuous action spaces, posing higher demands on the model’s fine-grained control and dynamics modeling abilities.
-ScaleZero also leverages the principles of Sampled Policy Iteration (Hubert et al., 2021) same as UniZero.
-Note that because observation space of DMC is vector, we use the MLP encoder.
-The experimental setup mirrors that of the Atari experiments, comparing the multitask ScaleZero against 18 independent single-task UniZero models.
-Results.
-The results in Table 2 show that ScaleZero achieves returns comparable to or even higher than the single-task baselines on the vast majority of DMC tasks.
-The learning curves, as illustrated in Figure 10, clearly reveal that ScaleZero can effectively handle the diverse physical dynamics and control requirements of different tasks within a single, unified model.
-This success validates the versatility of the ScaleZero architecture, proving its competitiveness in complex control problems involving high-dimensional continuous states and actions.
+(a) Human-normalized scores on 26 Atari games.
+ScaleZero achieves a higher mean score.
+Full curves are in Appendix B.
+We omit other MT baselines as we are unaware of prior work tackling all 26 games with a single online RL agent.
 Task UniZero (ST) ScaleZero (MT)
@@
 Median 875.1 887.3
-Table 2: Performance comparison between ScaleZero (MT) and UniZero (ST) across various tasks in DMControl.
+(b) Raw scores on 18 DMControl tasks.
+ScaleZero achieves a superior median score, indicating robust generalist performance.
+Full curves are in Appendix C.
 For reference, L2M (MT) Schmied et al.
-(2023) obtained a mean score of 840 under an offline supervised learning from expert data setting.
-The best performance for each task is indicated in bold.
-ScaleZero’s higher human-normalized median scores highlight its strong performance across continuous action spaces.
-Detailed curves are available in Appendix C  Figure 10.
-4.1.3 Performance Evaluation on the Jericho Text-Based Adventure Games
+(2023) reported a mean score of 840 on a 10-task subset of this benchmark, trained via offline supervised learning from expert data.
 Setup.
-To further evaluate ScaleZero’s text-based reasoning and planning, we assess it on the Jericho (Hausknecht et al., 2020) benchmark of interactive fiction, where both observations and actions are expressed in natural language.
-Jericho features a combinatorial action space and sparse, delayed rewards, and it demands strong long-horizon contextual reasoning.
-We use four representative games—Acorncourt, Omniquest, Detective, and Zork1—covering diverse narrative structures and interaction complexity.
-For parity, both the single-task UniZero and the multitask ScaleZero employ the same BGE text encoder (Xiao et al., 2023).
-In addition, we compare against CALM+OC, a recent LM-in-the-Loop method that adaptively updates the action generator using state-feature–based categorization of gameplay transitions (Sudhakar et al., 2023).
-The primary evaluation metric is the average return measured after training convergence.
+We evaluate ScaleZero on the 26 games of the Atari 100k benchmark, a standard testbed for heterogeneous MTRL due to its diverse visual styles, action spaces, and game mechanics.
+We compare the single multi-task (MT) ScaleZero agent against a strong single-task (ST) UniZero baseline, where 26 separate models are trained for each game.
+Performance is measured using the standard Human-Normalized Score (HNS) (Agarwal et al., 2021), reporting both mean and median to capture overall capability.
+Experimental details are provided in Appendix B.
 Results.
-The results in Table 3 show that ScaleZero achieves returns comparable to the single-task UniZero baseline and generally outperforms the CAML+OC method across most games.
-The learning curves in Figure 13 further indicate that a single, unified model can efficiently handle Jericho’s combinatorial language action space and sparse, delayed rewards.
-These findings validate the applicability of the ScaleZero architecture for text-based RL, demonstrating competitive performance on tasks that require language understanding and long-horizon planning.
-Table 3: Average returns of ScaleZero and other baselines on four Jericho tasks.
-CALM+OC(Sudhakar et al., 2023) denotes the variant of CALM with online categorization based on state features.
-Detailed learning curves for ScaleZero and UniZero are provided in Appendix D  Figure 13.
+As shown in Table 1, ScaleZero, as a single generalist agent, achieves a mean normalized score that surpasses the performance of 26 independently trained single-task specialists.
+This result demonstrates that ScaleZero effectively mitigates the negative interference observed in heterogeneous MTRL, enabling robust knowledge transfer.
+Notably, performance gains are particularly pronounced on notoriously difficult exploration tasks like Seaquest—a game we identified as a primary failure case for baseline models in our initial diagnosis (Figure 1) While the median score is slightly lower, influenced by a few hard-exploration games, the higher mean score confirms the overall advantage and generalization of our approach.
+4.1.2 DeepMind Control Suite
+Setup.
+To test ScaleZero’s generalization to continuous control, we evaluate it on 18 tasks from the DeepMind Control (DMC) Suite.
+These tasks, based on the MuJoCo physics engine (Todorov et al., 2012), feature low-dimensional state inputs and continuous action spaces, demanding precise dynamics modeling.
+Since observations are vector-based, we use an MLP encoder.
+The experimental protocol remains consistent, comparing the single MT ScaleZero agent against 18 individually trained ST UniZero models.
+Experimental details are provided in Appendix C.
+Results.
+1(b) demonstrate that ScaleZero achieves competitive performance, surpassing the baseline on several tasks and achieving a superior median score.
+This metric is particularly indicative of a strong generalist agent, as it shows robust performance across the majority of tasks rather than excelling on a few easy ones.
+This success validates the versatility of the ScaleZero, proving its effectiveness extends beyond visual domains to distinct and complex continuous control problems.
+4.1.3 Jericho Text-Based Adventure Games
+Setup.
+To assess performance on tasks requiring language understanding and long-horizon planning, we use the Jericho benchmark.
+These games feature combinatorial action spaces and sparse rewards, posing a reasoning challenge.
+We evaluate on four representative games and compare against the ST UniZero baseline and CALM+OC (Sudhakar et al., 2023), a recent language-model-based method.
+Both ScaleZero and UniZero use the same BGE text encoder (Xiao et al., 2023) for fair comparison.
+Results.
+As reported in Table 2, ScaleZero achieves performance on par with specialized single-task agents and is competitive with the strong CALM+OC baseline.
+This shows that a unified ScaleZero agent can effectively learn policies in environments where both states and actions are linguistic.
+These findings confirm that the design of ScaleZero are modality-agnostic, successfully extending to the domain of text-based reasoning and planning.
+Experimental details are provided in Appendix D.
+Table 2: Average returns comparison on four Jericho tasks. Detailed learning curves are in Appendix D.
 Algorithm Acorncourt Omniquest Detective Zork1
-CALM+OC (ST) 7.8 288.5 38.0
+CALM+OC (ST) (Sudhakar et al., 2023) 7.8 288.5 38.0
 UniZero (ST) 10 10 295 44
@@
 4.2 Efficiency Evaluation of Dynamic Parameter Scaling
-This section aims to evaluate the effectiveness of our LoRA-based online dynamic parameter scaling strategy integrated into ScaleZero, termed ScaleZero-DPS, in improving sample and computational efficiency.
-To this end, we conducted a comparative analysis between the standard ScaleZero and ScaleZero-DPS on the full DMC18 benchmark.
-The core evaluation metric is defined as the number of environment interaction steps required to reach a performance level comparable to the single-task baseline (i.e., 100% of its performance).
-As illustrated in Figure 5, ScaleZero-DPS achieves the target performance using only about 80% of the interaction steps required by the standard ScaleZero.
-This result highlights the significant efficiency gains afforded by our dynamic scaling strategy.
-Furthermore, the detailed performance curves, presented in Appendix Figure 11, reveal that ScaleZero-DPS exhibits a steeper learning curve.
-This indicates a faster learning rate, leading to higher sample and computational efficiency throughout the training process.
+Having established the performance of ScaleZero, we now evaluate the efficiency gains from our Dynamic Parameter Scaling strategy.
+We compare the standard ScaleZero against ScaleZero-DPS on the DMC18 benchmark.
+The primary metric is the number of environment interactions required to match the final performance of the single-task baselines.
+As shown in Figure 4, ScaleZero-DPS achieves this performance target using only 71.5% of the interactions needed by the standard ScaleZero.
+This corresponds to a substantial 28.5% reduction in sample complexity, directly validating the effectiveness in resolving the static resource allocation problem.
+The steeper learning curves shown in Figure 11 further confirm that this gain is consistent throughout the training process.
+A mechanistic analysis in Appendix C.4 reveals the agent’s efficiency stems from an emergent hierarchical strategy.
+The model learns to preserve foundational knowledge in early layers while applying targeted plasticity in later layers to adapt to new tasks, providing a clear mechanistic explanation for its performance.
 Figure 4: Comparison of interaction cost for ScaleZero vs.
 ScaleZero-DPS on DMC18.
-ScaleZero-DPS reaches the target performance with approximately 80% of the interaction steps.
-For detailed performance curves, please refer to Appendix C Figure 11.
+ScaleZero-DPS reaches the target performance with a 28.5% reduction in environment steps.
+Detailed curves are in Appendix C.
 4.3 Understanding MoE Efficacy: An Empirical and Theoretical Analysis
-The demonstrated success of Mixture-of-Experts (MoE) in multi-task scenarios motivates a deeper investigation into its underlying mechanisms.
-In the following, we explore these mechanisms from both experimental and theoretical perspectives.
-We begin by summarizing our empirical findings, which indicate that replacing standard MLP layers with MoE counterparts significantly reduces gradient conflicts.
-Notably, this reduction is not confined to the MoE layers themselves but also extends to adjacent network components, such as their input (detailed experimental settings and results are provided in Section˜E.1).
-To complement our empirical analysis, we provide a theoretical justification for how MoE mitigates gradient conflict.
-Our theory builds upon prior work (Chen et al., ) that characterizes MoE training by two distinct phases: an initial exploration stage, where inputs are routed nearly uniformly across experts, followed by a router learning stage, where the router gradually specializes in directing inputs.
-Grounded in this two-stage framework, we propose an informal theorem asserting that, regardless of the training stage, the upper bound on gradient conflict for an MoE layer is strictly lower than that of its MLP counterpart, and that this reduction is primarily governed by the router’s behavior.
+The superior performance of ScaleZero motivates a deeper investigation into its core architectural component: the Mixture-of-Experts backbone.
+Here, we analyze the mechanisms of MoE from both empirical and theoretical perspectives.
+Our empirical findings (detailed in Section˜E.1) reveal that replacing dense MLP layers with MoE counterparts not only reduces gradient conflict within the MoE layers themselves but also in adjacent network components.
+To explain this phenomenon, we provide a theoretical justification grounded in the two-phase training characterization of MoE (Chen et al., 2022).
+We propose the following informal theorem, which asserts that an MoE layer maintains a strictly lower upper bound on gradient conflict than a dense equivalent, governed primarily by the router’s specialization.
+Detailed explanations, proofs and corollaries are provided in Section˜E.2.
 Theorem 4.1 (Upper Bound on Gradient Conflict in MoE Layers (informal)).
@@
 Let $G$ be the maximum gradient conflict on any single expert, and $\cos(u,v)$ denote the cosine similarity between vectors $u$ and $v$.
-Then the following results hold:
-(1).
+Then the following results hold: (1).
 (Generally) Define the $m$-th element of vectors $u,v$ as $u_{m}=\lambda^{t_{1}}_{m}\|g_{t_{1}}^{(m)}\|,\;v_{m}=\lambda^{t_{2}}_{m}\|g_{t_{2}}^{(m)}\|,\;m=1,\dots,M.$ The full-layer gradient conflict admits a general upper bound: $\mathrm{conflict}(G_{t_{1}},G_{t_{2}})\leq G\cdot\cos(u,v)$.
 (2).
-(Sparse Router) In exploration stage, where tasks uniformly select experts, the expected gradient conflict of the MoE layer is approximately bounded by $G\cdot(1-M!/((M-K)!\,M^{K})$, where $K$ is the task number.
+(Sparse Router) In the exploration stage, where tasks uniformly select experts, the expected gradient conflict of the MoE layer is approximately bounded by $G\cdot(1-M!/((M-K)!\,M^{K}))$, where $K$ is the number of tasks.
 (3).
-(Sparse Router) In the router learning stage, where the expert sets selected by the router gradually stabilize, the gradient conflict of the MoE layer is roughly upper bounded by $\mathbb{E}[\text{conflict}_{i,j}]\approx G\cdot\frac{U}{ab}$, where $a=|S_{i}|$ and $b=|S_{j}|$ denote the sizes of the expert sets selected by tasks $i$ and $j$, and $U=|S_{i}\cap S_{j}|$ denotes the size of their intersection.
-In Section˜E.2, we provide detailed explanations of the the three parts of this theorem and their related corollaries, along with complete proofs.
+(Sparse Router) In the router learning stage, where expert sets stabilize, the gradient conflict is roughly upper bounded by $\mathbb{E}[\text{conflict}_{i,j}]\approx G\cdot\frac{U}{ab}$, where $a=|S_{i}|$, $b=|S_{j}|$ are the sizes of the expert sets for tasks $i$ and $j$, and $U=|S_{i}\cap S_{j}|$ is their intersection size.
 5 Related Work
-MCTS-based Planning with Learned World Models.
-The integration of MCTS with deep neural networks has become a cornerstone of modern reinforcement learning, particularly in domains requiring sophisticated planning.
-The seminal AlphaZero series (Silver et al., 2016; 2017) established a powerful paradigm by coupling MCTS with learned policy and value networks.
-This approach was further advanced by MuZero (Schrittwieser et al., 2019), which obviated the need for an explicit environment simulator by learning a latent world model to jointly predict future rewards, policies, and values.
-This architecture enables planning entirely within a learned representation space.
-Subsequent research has extended the MuZero framework, focusing on enhancing model capabilities (Hubert et al., 2021; Antonoglou et al., 2021), improving sample efficiency (Danihelka et al., 2022; Ye et al., 2021), and refining training stability (Schrittwieser et al., 2021).
-Notably, the advent of Transformer-based world models (Micheli et al., 2022; Robine et al., 2023), as exemplified by UniZero (Pu et al., 2024), has significantly augmented representational capacity, enabling stronger long-horizon planning and generalization.
-However, the monolithic, dense representations that enable single-task excellence become a liability in multi-task settings, causing representational interference and destructive interference.
-Simply scaling these models exacerbates this architectural bottleneck.
-Our work try to confront this fundamental limitation.
-Multi-Task Reinforcement Learning.
-To address the challenge of generalization across diverse tasks, MTRL aims to leverage shared knowledge to improve data efficiency and performance(Vithayathil Varghese & Mahmoud, 2020; D’Eramo et al., 2024).
-A prevalent architectural motif is the shared-backbone with task-specific heads (Kumar et al., 2022; Hansen et al., 2023; 2022; Lee et al., 2022a; Gallouédec et al., 2024a).
-While promoting knowledge transfer, this design is insufficient for world models where the core challenge lies in disentangling the latent dynamics predictions within the backbone itself, not just the final policy or value outputs.
-Recent research in MTRL has pursued several avenues to mitigate these issues.
-From a representational standpoint, context-based methods(Rakelly et al., 2019; Sodhani et al., 2021; Seo et al., 2022) condition the model on task metadata to enable selective knowledge sharing, while others introduce learnable modulation modules to preserve task-specific knowledge (Schmied et al., 2023).
-Architecturally, modular designs have emerged as a promising direction(Lan et al., 2023; Yang et al., 2020; Sun et al., 2022).
-These include soft modularization with task-aware routing (Yang et al., 2020; He et al., 2023; Hendawy et al., 2023) and constructing policy subspaces via parameter composition (Sun et al., 2022; 2023).
-On the optimization front, various strategies(Fernando et al., 2023; Lin et al., 2024; Ma et al., 2023) have been developed to manage conflicting gradients, a critical impediment in MTRL.
-These include projecting conflicting gradients into a non-conflicting space (Fernando et al., 2023) and harmonizing training through dynamic loss and gradient normalization (Lin et al., 2024; Ma et al., 2023; Chen et al., 2018).
-While effective at mitigating conflicts, these methods primarily address the symptoms (conflicting gradients) rather than the root cause: the architectural bottleneck of a model lacking sufficient specialized capacity to represent diverse task dynamics.
-These diverse approaches underscore the central challenge in MTRL: balancing knowledge sharing with task specialization.
-Our work seeks to achieve this balance through a novel architectural design.
+MCTS with Learned World Models.
+Planning in a learned latent space, popularized by MuZero (Schrittwieser et al., 2019), is a dominant RL paradigm (Hubert et al., 2021; Antonoglou et al., 2021; Danihelka et al., 2022; Xuan et al., 2024; Niu et al., 2024).
+Recent works have integrated Transformers to enhance representational capacity (Micheli et al., 2022; Pu et al., 2024; Zhou et al., 2024), achieving state-of-the-art performance in single-task domains.
+However, their monolithic design is a critical liability in heterogeneous multitask settings, where they suffer from representational interference and plasticity collapse.
+Our work directly confronts this architectural bottleneck.
+Multitask Reinforcement Learning.
+MTRL aims to improve data efficiency by sharing knowledge across tasks (Vithayathil Varghese & Mahmoud, 2020; Kumar et al., 2022; Hansen et al., 2023; Lee et al., 2022; Gallouédec et al., 2024).
+Existing methods mitigate interference through architectural solutions like task-specific modules (Schmied et al., 2023; Sun et al., 2022) or optimization-based approaches that manage gradient conflicts (Fernando et al., 2023; Ma et al., 2023).
+These strategies, however, have been predominantly studied outside of latent-space planning, where the unique challenge of disentangling dynamics prediction remains largely unaddressed.
 Sparse and Parameter-Efficient Architectures.
-Scaling deep RL models is computationally expensive (Espeholt et al., 2018).
-Sparse activation models, like the Mixture-of-Experts (MoE) (Dai et al., 2024), address this by increasing representational power at a near-constant computational cost via conditional computation.
-Their success in single-task RL (Obando-Ceron et al., 2024) highlights their potential as a native architectural prior for multi-task specialization.
-Concurrently, parameter-efficient fine-tuning (PEFT) methods (Dou et al., 2024; Dettmers et al., 2023; Hayou et al., 2024; Liu et al., 2024c) like Low-Rank Adaptation (LoRA) (Hu et al., 2022) enable lightweight, task-specific adaptations by tuning a small subset of parameters, avoiding the cost of full fine-tuning in MTRL (Wang et al., 2023; Yang et al., 2025; Agiza et al., 2024; Zhang et al., 2025).
-We identify a critical synergy between these two approaches.
-MoE enables coarse-grained, input-dependent specialization by routing computations to expert sub-networks, tackling representational entanglement at an architectural level.
-In parallel, LoRA facilitates fine-grained, task-specific adaptation in a parameter-efficient manner, injecting nuanced adjustments without catastrophic interference.
-The synergistic integration of MoE’s scalable representation with LoRA’s efficient adaptation is a critical frontier, yet one that has seen little exploration(Dou et al., 2023).
-Our work bridges this gap by proposing a novel architecture that unifies MoE and LoRA within a Transformer-based world model.
+We draw inspiration from two complementary paradigms.
+Sparse MoE models offer a natural architectural prior for multitask specialization by enabling conditional computation (Dai et al., 2024; Obando-Ceron et al., 2024).
+Concurrently, parameter-efficient methods like LoRA (Hu et al., 2022) provide a lightweight mechanism for task-specific adaptation (Wang et al., 2023; Zhang et al., 2025).
+The synergy of integrating these paradigms within a transformer-based world model is a promising but underexplored frontier (Dou et al., 2023; Huang et al., 2024) that we bridge in this work.
+See Appendix F for more discussion.
 6 Conclusion and Future work
-In this paper, we examine the diminishing plasticity phenomenon in unified world models under multi-task scenarios and quantify its impact on model performance.
-By exploring the expansive design space defined by the input–model–optimization axes, we introduce ScaleZero—a principled method that carefully balances performance with sample efficiency.
-Our extensive experiments on pure online RL benchmarks—which include 26 Atari, 18 DMControl, and 4 Jericho environments—demonstrate that ScaleZero achieves average returns comparable to or even surpassing those of specialized single-task baselines.
-To further balance computational costs, we introduced a LoRA-based dynamic parameter scaling (DPS) strategy.
-This strategy has currently only been prototyped on the DMC benchmark, where it achieved competitive performance with 80% of the interaction steps, showcasing its potential for improving sample efficiency.
-Despite these promising results, several avenues remain open for future investigation: A comprehensive evaluation of the DPS strategy’s generalization, scalability, and overhead is required on broader task sets like Atari.
-Investigating deeper synergistic mechanisms between MoE and LoRA to further enhance model expressivity and adaptability.
-Combining our online framework with offline pre-training of large-scale models to potentially boost both efficiency and final performance.
-We believe the continued exploration of these directions will advance multi-task reinforcement learning toward the goal of general intelligence.
+In this work, we diagnose and addressed the critical issue of plasticity collapse in multitask world models.
+We introduced ScaleZero, a unified architecture incorporating a sparse MoE backbone to create specialized computational pathways.
+Our experiments across 48 distinct tasks (Atari, DMControl, Jericho) demonstrate that ScaleZero robustly matches or surpasses the performance of specialized single-task agents.
+Furthermore, we propose DPS, a LoRA-based strategy that achieves competitive performance on DMControl while using just 71.5% of the environment interactions, highlighting a promising path toward greater sample efficiency.
+Our work provides a robust architectural foundation for developing more capable and efficient generalist agents.
+See Appendix G for a full discussion.
+Ethics Statement
+We have read and adhered to the ICLR Code of Ethics.
+Our research focuses on the development of novel algorithms for multitask reinforcement learning, specifically proposing ScaleZero and the Dynamic Parameter Scaling (DPS) strategy.
+All experiments were conducted on standard, publicly available benchmarks (Atari, DMC, and Jericho), which are simulated environments.
+Our work does not involve human subjects, nor does it use any personally identifiable or sensitive data.
+The proposed methods are foundational and do not have immediate, direct real-world applications that could pose ethical risks or societal harm.
+To the best of our knowledge, our research does not create or exacerbate discrimination or bias, as the training environments do not contain demographic or social data.
+We have strived to uphold the principles of scientific integrity and transparency by providing a thorough description of our methods and experimental setup.
+We declare no conflicts of interest.
+Reproducibility Statement
+To ensure the reproducibility of our findings, we provide comprehensive details in the appendix of this paper.
+The appendix includes: (1) A detailed description of the implementation of our proposed model, ScaleZero, and the online Dynamic Parameter Scaling (DPS) strategy.
+(2) A thorough account of the experimental setup, including all hyperparameters, evaluation protocols, and the computational infrastructure used.
+(3) The complete theoretical proofs for the claims made in the paper.
+Furthermore, we commit to releasing our source code publicly upon the completion of the peer-review process to allow for full verification of our results and to facilitate future research in this area.
+Appendix Contents
+  • A.
+    Implementation Details ........................................................................................................................................................................A
+  • B.
+    Atari Experiment Details ........................................................................................................................................................................B
+  • C.
+    DMC Experiment Details ........................................................................................................................................................................C
+  • D.
+    Jericho Experiment Details ........................................................................................................................................................................D
+  • E.
+    Gradient Analysis in MoE ........................................................................................................................................................................E
+  • F.
+    Detailed Related Work ........................................................................................................................................................................F
+  • G.
+    Detailed Conclusion and Future Work ........................................................................................................................................................................G
+  • H.
+    The Use of Large Language Models ........................................................................................................................................................................H
 Appendix A Implementation Details
 Appendix
-This appendix details the implementation of our ScaleZero model, explorations of key design spaces, the Dynamic Parameter Scaling strategy, hyperparameter configurations, and the computational resources used for our experiments.
-Unless otherwise specified, the training and inference details remain consistent with those in the paper Pu et al.
-(2024).
-A.1 Core Modifications of ScaleZero over UniZero
-Our core model, ScaleZero, is built upon the UniZero baseline, with key components upgraded to enhance learning efficiency and final performance in complex multi-task environments.
-Visual Encoder: For image-based inputs (e.g., Atari), we employ a Vision Transformer (ViT) as the primary encoder.
-The input image (64x64x3) is divided into patches (8x8), which are then converted into token embeddings via a linear projection layer.
-Learnable positional embeddings are added, and the sequence is fed into a standard Transformer encoder (ViT-Base architecture) to produce a 768-dimensional latent state.
-Text Encoder: In text-based environments like Jericho, we use a pre-trained BGE (BAAI General Embedding) model (Xiao et al., 2023) to encode the textual observations from the game into fixed-dimensional vectors.
-Unified World Model:
-  • Backbone: The core of the model is a Transformer network. It processes state tokens from the encoder, historical action tokens, and optional task-embedding tokens.
-  • UniZero (Baseline): The backbone consists of $N$ standard Transformer blocks. We use $N=8$ for the Atari26 benchmark and $N=4$ for other benchmarks.
-  • ScaleZero (Our Model): To maintain a comparable parameter count, the backbone is composed of $N/2$ Mixture-of-Experts (MoE) Transformer blocks.
-In each of these blocks, the standard feed-forward network (FFN) layer is replaced by an MoE layer.
-This layer comprises 8 expert networks and one shared expert, governed by a gating network that performs sparse routing.
-Latent State Normalization: In ScaleZero, we apply standard LayerNorm to normalize the latent state.
-A.2 Explorations on Key Design Spaces
-We conducted several exploratory extensions to key designs of UniZero to validate their effectiveness in multi-task settings.
-Task Information Encoding: To mitigate representational conflicts in multi-task learning, we explored explicit encoding of task information.
+This section provides a comprehensive overview of the implementation details for ScaleZero and the associated experimental methodology.
+For clarity, we first establish a baseline by providing a concise summary of the UniZero baseline (Pu et al., 2024), including its training and MCTS procedures.
+Building upon this foundation, we then sequentially detail the core modifications that define ScaleZero, our explorations of key design spaces, the Dynamic Parameter Scaling (DPS) strategy, and finally, the full hyperparameter and computational setup to ensure reproducibility.
+Our code is available at https://github.com/opendilab/LightZero.
+A.1 UniZero
+As outlined above, we begin by detailing the foundational framework of our baseline, UniZero, to provide the necessary context for our contributions.
+The learning process revolves around two core procedures operating in a loop:
+  1. 1.
+    collect_experience: The agent interacts with the environment to gather trajectories.
+At each timestep, it uses Monte Carlo Tree Search (MCTS) within its learned latent world model to generate an improved policy, from which an action is sampled.
+The resulting experience tuples (observation, action, reward, done) and the MCTS-improved policy are stored in a replay buffer.
+  2. 2.
+    update_world_model: The world model, policy network, and value network are jointly optimized using batches of data sampled from the replay buffer.
+This step updates the model’s understanding of the environment’s dynamics and improves its decision-making capabilities.
+A.1.1 MCTS in the Learned Latent Space
+The MCTS procedure is central to UniZero’s planning capability and is used during the collect_experience phase.
+For each action selection, a search tree is built in the latent space over a fixed number of simulations.
+Each simulation consists of three phases:
+  • Selection: Starting from the root node (the latent state of the current observation), the search traverses the tree by selecting actions that maximize an upper confidence bound score.
+This score, calculated using the PUCT formula, balances exploiting high-value actions (Q-value) and exploring less-visited actions (policy prior P).
+The action $a^{*}$ is chosen according to:
+    $a^{*}=\arg\max_{a}\left[Q(\hat{z},a)+P(\hat{z},a)\frac{\sqrt{\sum_{b}N(\hat{z},b)}}{1+N(\hat{z},a)}\left(c_{1}+\log\left(\frac{\sum_{b}N(\hat{z},b)+c_{2}+1}{c_{2}}\right)\right)\right]$ (2)
+    where $N$ is the visit count, $Q$ is the mean action-value, $P$ is the policy prior, and $c_{1},c_{2}$ are exploration constants.
+  • Expansion: When a leaf node is reached, the dynamics network is called to predict the next latent state $\hat{z}^{l+1}$ and reward $\hat{r}^{l}$.
+The decision network predicts a policy $p^{l}$ and value $v^{l}$ for the new state.
+This new node is added to the tree, and its outgoing edges are initialized with the predicted policy priors.
+  • Backup: The value $v^{l}$ predicted at the new leaf node is used to update the statistics of the nodes along the simulation path.
+The Q-values and visit counts $N$ for all parent state-action pairs are updated recursively back to the root.
+After all simulations are complete, the visit counts at the root node are normalized to form a temperature-controlled distribution, which serves as the improved policy for action selection.
+A.1.2 Model and Policy Optimization
+The update_world_model step optimizes the network parameters by minimizing a joint loss function over batches of trajectories sampled from the replay buffer.
+The optimization objective in UniZero is composed of four main prediction losses:
+$\begin{split}\mathcal{L}_{\text{UniZero}}(\theta)\doteq\mathop{\mathbb{E}}_{(\dots)\sim\mathcal{B}}\Big[\sum_{t=0}^{H-1}\Big(&\beta_{z}\underbrace{\|\hat{z}_{t+1}-\operatorname{sg}(\bar{h}(o_{t+1}))\|^{2}_{2}}_{\text{next-latent prediction}}+\beta_{r}\underbrace{\operatorname{CE}(\hat{r}_{t},r_{t})}_{\text{reward prediction}}\\ +&\beta_{p}\underbrace{\operatorname{CE}({p}_{t},\pi_{t})}_{\text{policy prediction}}+\beta_{v}\underbrace{\operatorname{CE}({v}_{t},\hat{v}^{targ}_{t})}_{\text{value prediction}}\Big)\Big]\end{split}$ (3)
+where $H$ is the training context length and the loss components are defined as follows:
+  • Next-latent Prediction: The model predicts the next latent state $\hat{z}_{t+1}$.
+The target is generated by a slowly-updating target encoder $\bar{h}$ and is detached from the computation graph using a stop-gradient operator ($\operatorname{sg}$).
+  • Reward and Value Prediction: The model predicts the immediate reward $\hat{r}_{t}$ and the state value $v_{t}$.
+To handle varying scales across tasks, both predictions are framed as discrete classification problems and optimized using a cross-entropy (CE) loss.
+The value target $\hat{v}^{targ}_{t}$ is a bootstrapped $n$-step TD target computed using future rewards and a target value network.
+  • Policy Prediction: The model’s policy head $p_{t}$ is trained to mimic the MCTS-improved policy $\pi_{t}$.
+This serves as a policy distillation step, which enhances training stability compared to direct policy gradient methods.
+The terms $\beta_{z},\beta_{r},\beta_{p},\beta_{v}$ are fixed coefficients that balance the contribution of each component.
+A.2 Core Modifications of ScaleZero over UniZero
+Our core model, ScaleZero, is built upon the UniZero baseline, with key components upgraded to enhance learning efficiency and final performance in complex multitask environments.
+The primary architectural differences are summarized in Table 3, with detailed explanations provided in the subsequent sections.
+Table 3: Summary of core architectural differences between UniZero and ScaleZero.
+Component UniZero (Baseline) ScaleZero (Our Model)
+Backbone Standard Transformer blocks. Mixture-of-Experts (MoE) Transformer blocks.
+Latent State Normalization SimNorm (L1 regularization). Standard LayerNorm.
+Visual Encoder ResNet-like architecture trained from scratch. Vision Transformer (ViT) trained from scratch.
+A.2.1 Backbone: From Dense Transformers to Sparse MoE
+The most significant modification lies in the model’s backbone.
+While both models employ a Transformer architecture to process state, action, and task tokens, their internal structure differs fundamentally.
+  • UniZero (Baseline): The backbone consists of $N$ standard Transformer blocks (e.g., $N=8$ for Atari26).
+  • ScaleZero (Our Model): To maintain a comparable parameter count while increasing model capacity, the backbone is composed of $N/2$ Mixture-of-Experts (MoE) Transformer blocks.
+In each block, the standard feed-forward network (FFN) is replaced by an MoE layer.
+This layer comprises 8 specialized expert networks and one shared expert, governed by a sparse top-1 routing gate.
+This design creates conditional computation pathways, mitigating task interference.
+A.2.2 Encoders and Normalization
+Encoders
+For encoding observations, ScaleZero adopts a similar strategy to the baseline but with a clear decision justified by empirical results.
+  • Visual Encoder (e.g., Atari): Our primary encoder is a Vision Transformer (ViT) trained from scratch.
+The input image (64x64x3) is divided into 8x8 patches, projected into token embeddings, and processed by a standard ViT-Base encoder to produce a 768-dimensional latent state.
+  • Text Encoder (e.g., Jericho): For text-based environments, we use a pre-trained BGE (BAAI General Embedding) model (Xiao et al., 2023) to encode textual observations into fixed-dimensional vectors.
+Justification for From-Scratch Visual Encoder
+In preliminary experiments, we explored leveraging a powerful pre-trained visual encoder, specifically a frozen DINOv2 model (dinov2_vits14) (Oquab et al., 2023), to potentially improve feature quality.
+However, this approach did not yield significant or consistent performance gains over the from-scratch ViT.
+We hypothesize this is due to the domain gap between the natural images DINOv2 was trained on and the distinct visual characteristics of Atari frames.
+This finding justifies our final choice and highlights the importance of domain-specific visual representations.
+We leave a more in-depth investigation of pre-trained models, including fine-tuning strategies, as future work.
+Latent State Normalization
+ScaleZero replaces the L1-based SimNorm used in UniZero with standard LayerNorm for normalizing the latent state. This change was found to be effective within our modified architecture.
+A.2.3 Task-Specific Architectural Configurations
+It is important to note that certain architectural choices are adapted based on the benchmark, rather than being a modification of ScaleZero over UniZero.
+These configurations apply to our entire framework.
+  • Encoders: For the DeepMind Control (DMC) suite, which features heterogeneous vector observation spaces, we employ independent MLP encoders for each task.
+For benchmarks with a unified input modality (images for Atari, text for Jericho), a single shared encoder is used across all tasks.
+  • Prediction Heads: To accommodate the distinct action spaces of each task, independent prediction heads (for policy, value, etc.) are universally employed across all benchmarks.
+A.3 Explorations on Key Design Spaces
+We conducted several exploratory extensions to key designs of UniZero to validate their effectiveness in multitask settings.
+Task Information Encoding: To mitigate representational conflicts in multitask learning, we explored explicit encoding of task information.
 By introducing a learnable task embedding matrix, task_embed = nn.Embedding(task_id), and concatenating it with the state representation, new_latent_state = concat(latent_state, task_embed), we aimed for the model to better distinguish between different tasks.
@@
 Crucially, to ensure knowledge sharing and generalization, all tokens are also processed by a shared expert network, whose output is added to the final result of the MoE layer.
-Multi-task Gradient Correction (MoCo): We adopted the MoCo algorithm from the LibMTL library [^3] https://github.com/median-research-group/LibMTL to handle gradient conflicts.
-Its core idea is to first maintain a momentum term $y_{t}$ to smoothly track each task’s gradient direction.
-Then, task weights $\lambda$ are dynamically updated by solving a regularized optimization problem.
-Finally, the gradient applied to shared parameters is the weighted sum of the momentum gradients.
+Multitask Gradient Correction with MoCo.
+In multitask learning, gradients from different tasks often conflict, hindering overall progress.
+To mitigate this, we adapt the MoCo algorithm (Fernando et al., 2023), originally from the LibMTL library [^3] https://github.com/median-research-group/LibMTL, for our distributed training framework.
+The core idea is to correct the gradients applied to the shared model parameters by dynamically adjusting task weights based on smoothed gradient directions.
+This process involves two key steps: momentum-based gradient smoothing and dynamic weight adjustment.
+  • (1).
+Momentum-based Gradient Smoothing.
+Instead of using the noisy, single-step gradient $g_{i}(t)$ for each task $i$ at step $t$, MoCo maintains a momentum-based moving average of the gradients, denoted as $y_{i}(t)$.
+This provides a more stable and reliable estimate of each task’s long-term descent direction.
+The update rule is:
+    $y_{i}(t)=(1-\beta)y_{i}(t-1)+\beta g^{\prime}_{i}(t)$ (5)
+    where $g^{\prime}_{i}(t)$ is the raw gradient $g_{i}(t)$ normalized and then scaled by its corresponding loss value, and $\beta$ is the momentum coefficient.
+  • (2).
+Dynamic Weight Adjustment.
+The task weights, $\lambda(t)=\{\lambda_{1}(t),\dots,\lambda_{N}(t)\}$, are updated to minimize the squared norm of the aggregated gradient, i.e., $\min_{\lambda}\frac{1}{2}\|\sum_{i=1}^{N}\lambda_{i}y_{i}(t)\|^{2}$.
+Our implementation approximates a gradient descent step on this objective, followed by a softmax projection to ensure the weights form a valid probability distribution.
+The update for the weights is:
+    $\lambda(t)=\text{softmax}\left(\lambda(t-1)-\gamma(Y(t)Y(t)^{T})\lambda(t-1)\right)$ (6)
+    where $Y(t)$ is the matrix whose columns are the momentum gradients $\{y_{i}(t)\}$, and $\gamma$ is the learning rate for the weights.
+The final corrected gradient applied to the shared parameters $\theta_{s}$ is then the weighted sum $g_{s}(t)=\sum_{i=1}^{N}\lambda_{i}(t)y_{i}(t)$.
+  • Our implementation is tailored for a Distributed Data Parallel (DDP) environment.
+All gradient correction computations occur on the rank-0 process, and the resulting corrected gradient is then broadcast to all other processes to ensure synchronized updates.
+The detailed procedure is outlined in Algorithm 1.
 Key hyperparameters are listed in Table 5.
+Algorithm 1 MoCo in a Distributed Data Parallel (DDP) Setting
+1:Shared parameters $\theta_{s}$, task losses $\{L_{i}\}_{i=1}^{N}$.
+2:Momentum gradients $\{y_{i}\}_{i=1}^{N}$, task weights $\{\lambda_{i}\}_{i=1}^{N}$.
+3:Hyperparameters: Momentum coefficient $\beta$, weight learning rate $\gamma$.
+4:Updated shared parameters $\theta_{s}$.
+5:for each training step do
+[…]
+7:  Compute local raw gradients $\{g_{i}^{\text{local}}\}$ for tasks on process pp.
+8:  Gather all local gradients $\{g_{i}^{\text{local}}\}$ and losses $\{L_{i}^{\text{local}}\}$ to the rank-0 process.
+[…]
+10:  Let $\{g_{i}\}_{i=1}^{N}$ and $\{L_{i}\}_{i=1}^{N}$ be the full sets of gradients and losses.
+11:  for $i=1$ to NN do
+12:  Normalize and scale gradient: $g^{\prime}_{i}\leftarrow\frac{g_{i}}{\|g_{i}\|+\epsilon}\cdot L_{i}$.
+13:  Update momentum gradient: $y_{i}\leftarrow(1-\beta)y_{i}+\beta g^{\prime}_{i}$.
+14:  end for
+15:  Let $Y\leftarrow[y_{1},\dots,y_{N}]$.
+16:  Update task weights: $\lambda\leftarrow\text{softmax}(\lambda-\gamma(YY^{T})\lambda)$.
+17:  Compute final corrected gradient: $g_{s}\leftarrow\sum_{i=1}^{N}\lambda_{i}y_{i}$.
+18:  Broadcast the corrected gradient $g_{s}$ from rank-0 to all other processes.
+[…]
+20:  Set gradients of shared parameters $\nabla\theta_{s}\leftarrow g_{s}$.
+21:  Perform optimizer step on $\theta_{s}$.
+22:end for
 Table 5: Hyperparameters for the MoCo algorithm.
-Hyperparameter Value
-MoCo_beta ($\beta$) 0.99
-MoCo_gamma ($\gamma$) 10.0
-MoCo_rho 0.0
-A.3 Dynamic Parameter Scaling (DPS)
-To address the issue of non-uniform learning progress across tasks in multi-task learning, we propose the Dynamic Parameter Scaling (DPS) strategy.
-The core idea is to dynamically align the model’s effective capacity with the set of currently unsolved tasks.
-This is achieved through a curriculum-based expansion of the model’s parameters.
-The detailed pseudocode is provided in Algorithm 1.
-Staged Training and Active Set Management.
-The training protocol is divided into $S+1$ stages.
-We define a task $\tau_{i}$ as "solved" once its performance metric surpasses a predefined threshold $\varepsilon_{i}$.
-The set of all currently unsolved tasks forms the active set $\mathcal{U}_{t}$.
-To maximize efficiency, both data collection and training updates are performed exclusively on tasks within this active set.
-  • Stage 0 (Warm-up): During an initial period of $T_{0}$ iterations, only the shared backbone network parameters $\theta_{B}$ are trained.
-This allows the model to learn a foundational, general-purpose representation across all tasks.
-  • Stage $s\geq 1$ (Expansion): Each subsequent stage introduces new, dedicated parameters in the form of an independent Low-Rank Adaptation (LoRA) module, $\Delta\theta_{s}=B_{s}A_{s}$, and a corresponding learnable scalar scaling factor $\alpha_{s}$.
-These new parameters are intended to provide additional capacity for tackling the remaining difficult tasks in $\mathcal{U}_{t}$.
-Adaptive Stage Transition Triggers. The transition from stage s1s-1 to stage $s$ is governed by a dual-trigger mechanism, ensuring a balance between progress and resource allocation:
-  1. 1.
-    Progress-Driven Trigger: A new stage is initiated if the number of newly solved tasks since the beginning of the current stage reaches a predefined quota $Q_{s}$.
-This allows the model to expand only when it has demonstrated mastery over a subset of the current challenges.
-  2. 2.
-    Budget-Driven Trigger: To prevent stagnation on particularly hard tasks, a transition is forced if the number of iterations within the current stage exceeds a maximum budget, calculated as $\lceil(T_{\max}-T_{0})/S\rceil$.
-This ensures that the model continues to gain new capacity even if progress stalls.
-Hierarchical Parameter Composition and Optimization. In any given stage $s$, the forward pass for a weight matrix $W_{0}\in\theta_{B}$ is dynamically composed as:
-$W^{(s)}(x)=W_{0}x+\sum_{j=1}^{s}\alpha_{j}(B_{j}A_{j})x$ (3)
-Upon entering a new stage $s$, we implement a crucial optimization strategy: we freeze the backbone parameters $\theta_{B}$ and all previously introduced LoRA modules $\{\Delta\theta_{j}\}_{j=1}^{s-1}$.
-The optimization gradient is then applied only to the newly introduced LoRA module $\Delta\theta_{s}=\{A_{s},B_{s}\}$ and to all existing scaling factors $\{\alpha_{j}\}_{j=1}^{s}$.
-This design provides two key benefits: it mitigates negative transfer by isolating new parameter updates within the fresh LoRA module, and it retains plasticity by allowing the model to dynamically re-weigh the contribution of knowledge from all previous stages via the trainable scalars $\{\alpha_{j}\}$.
-The specific DPS settings used in our ScaleZero-Balanced experiments are detailed in Table 6.
-Table 6: Hyperparameters for Dynamic Parameter Scaling (DPS).
-Hyperparameter Value Description
-curriculum_stage_num 5 Total number of expansion stages ($S$).
-lora_r 64 The rank of the LoRA matrices ($r$).
-lora_alpha 1 Initial scaling factor for LoRA layers.
-min_stage0_iters 10,000 Duration of the warm-up stage ($T_{0}$).
-max_stage_iters 5,000 Per-stage iteration budget for the budget-driven trigger.
+Hyperparameter Symbol Value
+Momentum Coefficient $\beta$ 0.99
+Weight Learning Rate $\gamma$ 10.0
+Weight Regularization $\rho$ 0.0
+A.4 Dynamic Parameter Scaling (DPS)
+This section provides a detailed breakdown of the Dynamic Parameter Scaling (DPS) strategy introduced in Section 3.3.
+DPS is designed to dynamically couple model capacity with learning progress, creating a "curriculum of model complexity" that directs resources precisely where they are most required.
+Detailed pseudocode is provided in Algorithm 2.
 Algorithm 2 Dynamic Parameter Scaling (DPS)
@@
 33:end for
-A.4 Common Settings and Hyperparameters
-To ensure a fair comparison, most hyperparameters are kept identical to the original UniZero paper. We use the AdamW optimizer. Key shared hyperparameters are summarized in Table 7.
-Table 7: Common hyperparameters for all experiments.
+Task Curation and Active Set Management.
+As described in the main text, we define the full set of tasks as $\mathcal{T}=\{\tau_{i}\}_{i=1}^{N}$.
+A task $\tau_{i}$ is deemed "solved" once its performance metric, $\text{Metric}(\tau_{i})$, surpasses a predefined threshold $\varepsilon_{i}$.
+At any time $t$ during training, we maintain an active set of all unsolved tasks, $\mathcal{U}_{t}\subseteq\mathcal{T}$.
+To maximize computational efficiency, both data collection and gradient updates are performed exclusively for tasks within this active set $\mathcal{U}_{t}$.
+Once a task is solved, it is removed from the active set, thereby ceasing all computational overhead associated with it.
+Staged Capacity Expansion. The training process is partitioned into $S+1$ stages, following the protocol outlined in the main text:
+  • Stage 0 (Warm-up): During an initial period of $T_{0}$ iterations, only the shared backbone network parameters, $\theta_{B}$, are trained.
+This stage allows the model to learn a foundational, general-purpose representation across all tasks.
+  • Stage $s\geq 1$ (Expansion): Each subsequent stage introduces a new, independent LoRA module, $\Delta\theta_{s}$.
+This module comprises a trainable low-rank matrix pair $\Delta\theta_{s}=B_{s}A_{s}$ (with rank $r\ll d,k$) and a corresponding learnable scalar scaling factor $\alpha_{s}$.
+These new parameters provide additional model capacity for tackling the remaining difficult tasks in the active set.
+Adaptive Stage Transition Triggers. The transition from stage s1s-1 to stage $s$ is governed by a dual-trigger mechanism, balancing learning progress against a fixed budget:
+  1. 1.
+    Progress-based Trigger: A new stage is initiated if the number of newly solved tasks within the current stage reaches a predefined quota $Q_{s}$.
+  2. 2.
+    Budget-based Trigger: To prevent stagnation on particularly arduous tasks, a transition is forced if the iteration count within the current stage exceeds a pre-allocated limit, calculated as $\lceil(T_{\max}-T_{0})/S\rceil$.
+Optimization and Parameter Isolation. In any given stage $s$, the forward pass for a weight matrix $W_{0}\in\theta_{B}$ is dynamically composed as:
+$W^{(s)}(x)=W_{0}x+\sum_{j=1}^{s}\alpha_{j}(B_{j}A_{j})x$ (7)
+Crucially, upon entering a new stage $s$, all prior parameters—the backbone $\theta_{B}$ and the matrices of all previously introduced adapters $\{A_{j},B_{j}\}_{j=1}^{s-1}$—are frozen.
+Optimization is thereby focused exclusively on the newly added LoRA module $\Delta\theta_{s}=\{A_{s},B_{s}\}$ and on the full set of currently active scaling factors $\{\alpha_{j}\}_{j=1}^{s}$.
+In practice, this is achieved by using ‘torch.detach()‘ on the outputs of adapters from previous stages, which stops their gradients while preserving their forward contribution.
+This strategy provides two decisive advantages:
+  • It implements a resource-efficient training curriculum, as new parameters are only introduced when required by the remaining unsolved tasks.
+  • By isolating new learning within dedicated adapters, DPS provides targeted plasticity for difficult tasks while preventing catastrophic forgetting or negative transfer to previously mastered skills, whose knowledge is preserved in the frozen backbone and prior adapters.
+Stabilized Learnable Scaling Factors.
+To prevent training instability from unbounded learnable scalars, we constrain the scaling factors $\alpha_{j}$ within a narrow, stable range (e.g., $[0.8,1.2]$).
+This is achieved via a re-parameterization trick, where each $\alpha_{j}$ is computed from an unbounded underlying parameter $\hat{\alpha}_{j}$ as follows:
+$\alpha_{j}=\text{offset}+\text{range}\cdot\tanh(\hat{\alpha}_{j})$ (8)
+This design ensures that the contribution of each adapter is modulated smoothly without causing drastic shifts in the output distribution.
+The specific DPS hyperparameters used in our ScaleZero-DPS experiments are detailed in Table 6.
+Table 6: Hyperparameter configuration for Dynamic Parameter Scaling in our ScaleZero-DPS experiments.
+Hyperparameter Symbol Value Description
+curriculum_stage_num $S+1$ 5 Total number of stages (1 warm-up + 4 expansion).
+lora_r $r$ 64 The rank of the LoRA matrices.
+lora_alpha - 1 Initial value for LoRA scaling (for ‘scaling = alpha / r‘).
+lora_scale_init $\alpha_{j,\text{init}}$ 1.0 Initial value of the learnable scaling factors.
+lora_scale_range - 0.2 Range for learnable scalars, yielding $\alpha_{j}\in[0.8,1.2]$.
+min_stage0_iters $T_{0}$ 10,000 Number of iterations for the warm-up stage.
+max_stage_iters - 5,000 Per-stage iteration budget, i.e., $\lceil(T_{\max}-T_{0})/S\rceil$.
+A.5 Hyperparameter Settings
+Our hyperparameter configuration for ScaleZero largely follows the original UniZero paper to ensure a fair comparison.
+We employ the AdamW optimizer for all experiments.
+Table 7 lists the hyperparameters that are kept consistent across all benchmarks.
+Architectural configurations and loss weights are detailed subsequently, followed by benchmark-specific settings in Tables 9, 10, and 11.
+Table 7: Hyperparameters shared across all benchmarks.
 Hyperparameter Value
 General Training
+Optimizer AdamW
 Learning Rate 1e-4
-Weight Decay 0.0001
-Batch Size 256
-Discount Factor ($\gamma$) 0.997
-Training Context Length ($H$) 10
-World Model
-Num. Transformer Layers 4 (Atari-8, DMC, Jericho), 8 (Atari-26)
+Weight Decay 1e-4
+World Model (Transformer)
 Embedding Dimension 768
-Unroll Steps 5
-Num. Experts (per MoE layer) 8 + 1 (shared)
+Number of Heads 8
 MCTS
 Num. Simulations 50
-Inference Context Length ($H_{\text{infer}}$) 4
-Replay Buffer
-Replay Buffer Size 100,000
-Loss Function Weights. The weights for the loss function are configured based on the action space type of the environment, as detailed in Table 8.
+Our ScaleZero model incorporates Mixture-of-Experts layers within its Transformer backbone for the Atari and DMC benchmarks to enhance model capacity and specialization.
+In contrast, the model for the Jericho benchmark does not use MoE layers.
+When MoE is used, each MoE layer consists of 8 specialist experts and 1 shared expert, with the router selecting the top-1 expert per token.
+The final loss is a weighted sum of several components, with weights varying between discrete and continuous action space environments, as detailed in Table 8.
 Table 8: Loss function weights for different environment types.
-Loss Term Discrete (Atari, Jericho) Continuous (DMC)
-Observation Loss 10.0 10.0
+Loss Term Discrete Action Spaces (Atari, Jericho) Continuous Action Spaces (DMC)
 Value Loss 0.5 0.1
@@
 Policy Loss 1.0 0.1
-A.5 Training Framework and Computational Cost
-All experiments were implemented using the LightZero [^4] https://github.com/opendilab/LightZero framework.
-All multi-task learning experiments were conducted on a single node equipped with 8x NVIDIA A100 (80G) GPUs.
+Observation Loss 10.0 10.0
+A.5.1 Atari-Specific Settings
+For Atari, observations are $64\times 64\times 3$ images processed by a ViT encoder. Training is conducted for 400,000 total environment steps.
+Table 9: Key hyperparameters for Atari experiments.
+Hyperparameter Value
+General Training
+Effective Global Batch Size 512
+Discount Factor ($\gamma$) 0.997
+World Model
+Num. Transformer Layers 2 (Atari-8), 4 (Atari-26)
+Training Context Length ($H$) 10
+Encoder Type ViT
+MCTS
+Inference Context Length ($H_{\text{infer}}$) 4
+Replay Buffer
+Replay Buffer Size 500,000
+Replay Ratio 0.25 (4-layer model), 0.5 (8-layer model)
+A.5.2 DMC-Specific Settings
+For the DMC suite, state vectors are processed by an MLP encoder. Training runs for 400,000 total environment steps.
+Table 10: Key hyperparameters for DMC experiments.
+Hyperparameter Value
+General Training
+Effective Global Batch Size 512
+Discount Factor ($\gamma$) 0.99
+Frame Skip 4 (8 for Pendulum)
+World Model
+Num. Transformer Layers 4
+Training Context Length ($H$) 5
+Encoder Type MLP
+Num. Sampled Actions 20
+MCTS
+Inference Context Length ($H_{\text{infer}}$) 2
+Replay Buffer
+Replay Buffer Size 1,000,000
+Replay Ratio 0.25
+A.5.3 Jericho-Specific Settings
+For the text-based Jericho suite, a pre-trained language model (BAAI/bge-base-en-v1.5) is used as the text encoder. Training runs for 500,000 total environment steps.
+Table 11: Key hyperparameters for Jericho experiments.
+Hyperparameter Value
+General Training
+Effective Global Batch Size 256
+Discount Factor ($\gamma$) 0.997
+World Model
+Num. Transformer Layers 2
+MoE Layers None
+Training Context Length ($H$) 10
+Encoder Type Text (BAAI/bge-base-en-v1.5)
+MCTS
+Inference Context Length ($H_{\text{infer}}$) 4
+Replay Buffer
+Replay Buffer Size 500,000
+Replay Ratio 0.1
+A.6 Training Framework and Computational Cost
+All multitask learning experiments were conducted using one open-sourced framework on a single node equipped with 8x NVIDIA A100 (80G) GPUs.
 The approximate training times are shown in Table 12.
-Table 9: Approximate training times for multi-task learning experiments.
-Task Approximate Training Time
+We promise to open source the relevant code after the review process is completed.
+A.6.1 Multi-Task Training Implementation Details
+Our multi-GPU, multitask training framework is implemented based on PyTorch’s Distributed Data Parallel (DDP) (Li et al., 2020). The core logic is as follows:
+  • Static Task Partitioning: The entire set of tasks (e.g., 26 Atari games) is statically partitioned among the available GPUs (ranks). Each GPU is assigned an exclusive subset of tasks to manage.
+  • Per-Task Resource Allocation: On each GPU, dedicated resources, including a data collector, an evaluator, and a replay buffer, are instantiated for each assigned task.
+  • Aggregated Batch Construction: During each training iteration, every GPU constructs a single training batch by sampling data segments from all replay buffers it manages.
+This creates a heterogeneous batch containing experiences from multiple tasks.
+  • Synchronized Gradient Update: This aggregated batch is processed by the shared model to compute a single gradient.
+The DDP framework then averages the gradients across all GPUs, and the resulting synchronized gradient is used to perform a single optimization step on the shared model parameters.
+This approach enables the joint optimization of a single model using data streams from multiple environments in a distributed and efficient manner.
+Table 12: Approximate training times for multitask learning experiments in Section 4.
+Benchmark Approximate Training Time
 Atari-8 $\sim$2 days
-Atari-26 $\sim$5 days
+Atari-26 $\sim$8 days
 DMC-18 $\sim$1.5 days
@@
 B.1 Benchmark Setup
-We use the Arcade Learning Environment (ALE) (Bellemare et al., 2013b) as our evaluation platform.
-Our experiments adopt two distinct training paradigms: multitask learning (MT), where a single model is trained to master all tasks simultaneously, and single-task learning (ST), where a dedicated model is trained independently for each task.
-For our primary evaluation, we compare our multitask model, ScaleZero, with a robust single-task baseline using individually trained UniZero models.
-We use ScaleZero to denote the multitask (MT) setting and UniZero to denote the single-task (ST) setting.
-The MT/ST notation is sometimes omitted for brevity where the context is unambiguous.
-This comparison strategy is chosen because the multitask version of UniZero performs poorly, and there is currently no widely accepted standard baseline for multitask learning on the Atari benchmark.
-Consequently, UniZero ST serves as a more challenging and meaningful point of reference.
-Our experiments are conducted in two stages: first, we perform model ablation studies on the Atari8 multitask benchmark (comprising Alien, Boxing, ChopperCommand, Hero, MsPacman, Pong, RoadRunner, and Seaquest); subsequently, we evaluate the final model’s performance on the full Atari26 multitask benchmark.
-All game observations are preprocessed into 64x64 grayscale images, with 4-frame stacking used for state input.
-We employ standard environment wrappers, including sticky actions (p=0.25) and a frame skip of 4(Mnih et al., 2013).
-Performance is measured by the Human-Normalized Score (HNS), and we report both the mean and median scores.
-To ensure a fair comparison, all environmental settings are aligned with those in the UniZero paper (Pu et al., 2024).
-B.2 Supplemental Plasticity Analysis
+This section details the experimental setup for the Atari benchmarks discussed in Section 4.1. We use the Arcade Learning Environment (ALE) (Bellemare et al., 2013) as our evaluation platform.
+Our evaluation protocol compares our multitask model, ScaleZero (MT), against a strong single-task baseline comprising individually trained UniZero (ST) models.
+This comparison is chosen because the multitask version of UniZero exhibits significant performance degradation, and a standard, widely-accepted baseline for online MTRL on the full Atari suite is not available.
+The UniZero ST baseline thus provides a more challenging and meaningful point of reference for evaluating positive knowledge transfer.
+Our experiments proceed in two stages: first, we conduct model design and ablation studies on the Atari8 (multitask) benchmark (a subset of 8 games: Alien, Boxing, ChopperCommand, Hero, MsPacman, Pong, RoadRunner, and Seaquest).
+Second, we evaluate the final ScaleZero architecture on the full Atari26 (multitask) benchmark.
+For all games, observations are preprocessed into 64x64 RGB images.
+The state input to the model is a single frame, resulting in an input tensor of shape 3x64x64 without any frame stacking.
+We use standard environment wrappers, including sticky actions (p=0.25) and a frame skip of 4 (Mnih et al., 2013).
+Performance is measured by the Human-Normalized Score (HNS), with both mean and median scores reported to assess overall capability.
+All environmental settings are strictly aligned with those in the UniZero paper (Pu et al., 2024) and DI-engine (Niu et al., 2021) to ensure a fair comparison.
+B.2 Supplementary Analysis of Plasticity and Representational Collapse
+This section provides a deeper empirical analysis of the plasticity loss and representational collapse phenomena diagnosed in Section 4.
 B.2.1 Training Dynamics and Representational Collapse in UniZero
-While Figure 1 in the main text shows performance degradation and worsening plasticity metrics for UniZero on complex tasks, Figure 7 illustrates that on four relatively simpler tasks from the Atari8 set, UniZero’s training process and corresponding metrics remain stable.
-This suggests a correlation between task difficulty and plasticity degradation.
-To further probe the underlying cause of this degradation, we introduce the concept of effective rank (Dohare et al., 2024) to quantify the dimensionality of the model’s representation space.
-A low effective rank suggests representational collapse, where the model fails to maintain diverse features across tasks.
-Formally, consider a matrix $\bm{\Phi}\in\mathbb{R}^{n\times m}$ with singular values $\sigma_{k}$ for $k=1,2,\ldots,q$, where $q=\min(n,m)$.
-Let $p_{k}=\sigma_{k}/\sum_{i=1}^{q}\sigma_{i}$ be the normalized singular values, which form a probability distribution.
-The effective rank of $\bm{\Phi}$ is then defined as the exponential of the entropy of this distribution:
+While Figure 1 in the main text highlights performance degradation and plasticity loss for UniZero on complex tasks, Figure 7 shows that on four simpler tasks from the Atari8 set, UniZero’s training remains stable.
+This suggests a strong correlation between task difficulty and the onset of plasticity degradation.
+To quantitatively diagnose the cause of this degradation, we measure the effective rank (Dohare et al., 2024) of the model’s representation space.
+A low effective rank indicates representational collapse, where the model fails to maintain diverse features required for multitask learning.
+For a matrix $\bm{\Phi}\in\mathbb{R}^{n\times m}$ with normalized singular values $p_{k}$, the effective rank is the exponential of the Shannon entropy of its singular value distribution:
 $\operatorname{erank}(\bm{\Phi})\doteq\exp\left(-\sum_{k=1}^{q}p_{k}\log(p_{k})\right)$ (9)
-The effective rank is a continuous measure ranging from 1 to the conventional rank of the matrix, providing a more nuanced view of the dimensionality of the space spanned by the matrix’s columns or rows.
-In the context of a neural network, the effective rank of a hidden layer’s activation matrix measures the effective number of dimensions in its representation space.
-A low effective rank indicates that the layer’s output can be generated by a small number of units, implying that many neurons are not contributing useful, independent information.
-In our experiments, we estimate the effective rank from a minibatch of 256 sequence samples before training on each task.
-Figure 6 illustrates this concept, providing a diagnostic view of the representation’s dimensionality.
-The observation that a standard baseline model can exhibit a low effective rank provides quantitative evidence for the representational collapse hypothesis.
-Figure 6: Supplementary analysis of return and representation effective rank.
-This figure provides additional diagnostics for the baseline model shown in Figure 1, illustrating the relationship between game return (Left) and representation effective rank (Right).
-The performance degradation, observed as a drop in return, is largely accompanied by a corresponding decline in the representation effective rank.
-This provides further evidence that the performance collapse is linked to a degeneration of the model’s representation space.
-Figure 7: Training dynamics of UniZero on four other tasks from the Atari8 benchmark.
-This figure complements Figure 1 of the main text, showing the performance curves, dormant neuron ratio of the Transformer backbone, and latent state norm for Boxing, MsPacman, RoadRunner, and Alien.
-B.2.2 Complete Metrics for Design Space Ablation Study
-Figure 8 provides the complete data corresponding to Figure 3 in the main text, detailing the plasticity metrics for different model variants across all Atari8 tasks.
-Compared to the Baseline and different variants, the MoE architecture more effectively suppresses the growth of the dormant neuron ratio and the inflation of the latent state norm during training.
-This suggests that the MoE architecture is better at maintaining representational diversity and plasticity.
+In our context, we compute the effective rank of a hidden layer’s activation matrix from a minibatch of 256 sequence samples.
+As shown in Figure 5, the performance collapse in the baseline model is directly accompanied by a decline in the representation’s effective rank, providing quantitative evidence for the representational collapse hypothesis.
+Figure 5: Analysis of representation effective rank.
+This figure supplements the diagnosis in Figure 1, showing the relationship between game return (Left) and representation effective rank (Right) for the baseline model.
+The sharp drop in performance correlates strongly with a decline in effective rank, substantiating the claim that performance failure is linked to a collapse in the dimensionality of the model’s representation space.
+Figure 6: Correlation of performance with plasticity metrics for model variants.
+On two challenging tasks (ChopperCommand and Seaquest), the performance collapse of baseline models directly correlates with catastrophic plasticity loss, evidenced by a sharp rise in the dormant neuron ratio and latent norm.
+In stark contrast, the MoE-based model (ScaleZero) sustains high returns by maintaining a near-zero dormant ratio and a stable latent norm.
+This confirms that MoE’s superior performance stems directly from its success in mitigating the diagnosed plasticity issues.
+Full results for all 8 tasks are in Figure 8.
+Figure 7: Training dynamics of UniZero on four simpler Atari8 tasks.
+Complementing Figure 1 from the main text, this figure shows stable performance curves, dormant neuron ratios, and latent state norms for UniZero on Boxing, MsPacman, RoadRunner, and Alien, suggesting plasticity loss is task-difficulty dependent.
+B.2.2 Mechanism of Representational Collapse
+The correlations observed in our experiments reveal a critical failure mode in standard multitask architectures.
+The model, struggling to reconcile conflicting task demands, forces its feature space to degenerate into a low-dimensional, "one-size-fits-all" subspace, which is suboptimal for any individual task.
+This is the mechanism of representational collapse, quantitatively diagnosed by a low effective rank.
+This contraction of the feature space renders many neurons redundant, leading to a rising dormant neuron ratio.
+To compensate for this impoverished representational capacity, the model is forced to scale the magnitude of the few remaining active features, causing an inflation of the latent norm.
+Ultimately, this collapsed representation lacks the capacity to model task diversity, resulting in poor performance.
+In contrast, ScaleZero’s MoE architecture is designed to mitigate this exact problem.
+By routing tasks to specialized expert sub-networks, it preserves distinct representational subspaces.
+As shown in Figure 8, the MoE architecture effectively maintains neuron activity (low dormant ratio) and stabilizes representation magnitudes (controlled latent norm), thereby preventing representational collapse and enabling superior multitask performance.
 Figure 8: Complete plasticity metric analysis for the ScaleZero design space exploration.
-This figure details the evolution of performance (Return), Dormant Ratio, and Latent Norm for different model variants across all 8 tasks in the Atari8 benchmark.
-Intrinsic Connections and Performance Implications
-The observed correlations highlight a critical challenge in multi-task learning for standard Transformer architectures: representational collapse.
-The mechanism of this phenomenon is as follows: the model struggles to maintain distinct representations across tasks, causing its feature space to degenerate into a "one-size-fits-all" low-dimensional space, which is suboptimal for any individual task.
-This collapse is quantitatively indicated by a low effective rank, a problem diagnosed in baseline models (Section 1).
-The contraction of the representation space renders many neurons redundant, leading to a rising dormant neuron ratio.
-To compensate for this impoverished representational capacity, the model is forced to scale the magnitude of remaining active representations, which in turn causes the inflation of the latent norm.
-Ultimately, this collapsed and impoverished representation space lacks the capacity to model the diversity of multiple tasks, resulting in poor overall performance.
-In contrast, the MoE variant is designed to mitigate this very problem.
-By using its expert-based routing to preserve specialized representational subspaces, it addresses the symptoms of representational collapse.
-As shown in Figure 8, the MoE architecture effectively promotes neuron activity (suppressing the dormant ratio) and stabilizes representation magnitudes (controlling latent norm inflation), leading to superior multitask performance.
-B.3 Performance on the Atari26 Multitask Benchmark
-As shown in Table 1 of the main text, the mean HNS of the single multitask ScaleZero (MT) model on the Atari26 benchmark surpasses the average HNS of 26 independently trained UniZero single-task (ST) models.
-This result suggests that the multitask model achieves positive knowledge transfer across tasks.
-We hypothesize that this is due to the MoE architecture maintaining the network’s learning plasticity, enabling the transfer of general priors (e.g., object motion, collision detection) from simpler tasks to more complex ones.
-However, the multitask model’s median HNS is lower than the single-task baseline.
-This is primarily influenced by a few tasks with high exploration difficulty or unique mechanics (e.g., PrivateEye, Pitfall), indicating that negative transfer between heterogeneous tasks is not fully resolved in the current architecture.
-Figure 9 provides a comparison of the learning curves for all tasks.
-Figure 9: Performance comparison between ScaleZero (MT) and UniZero (ST) on the Atari26 benchmark.
-The figure shows the learning curves (mean and 95% confidence interval) for the multitask ScaleZero and the single-task UniZero baselines across 26 Atari games.
+This figure provides the full data for the ablation study, detailing the evolution of performance (Return), Dormant Ratio, and Latent Norm for different model variants across all 8 tasks in the Atari8 multitask benchmark.
+The MoE-based model consistently outperforms others by maintaining superior plasticity metrics.
+B.3 Full Performance on the Atari26 Benchmark
+This section provides the complete learning curves corresponding to the summary results in 1(a).
+As reported in the main text, the single multitask ScaleZero (MT) agent achieves a higher mean HNS than the average of 26 specialized UniZero (ST) agents.
+This result demonstrates that ScaleZero achieves net positive knowledge transfer, which we attribute to the MoE architecture’s ability to maintain plasticity and transfer general priors (e.g., object physics) from simple to complex tasks.
+However, the median HNS of ScaleZero is lower than the ST baseline.
+This is primarily influenced by poor performance on a few hard-exploration or mechanically unique games (e.g., PrivateEye, Pitfall), indicating that negative interference is not fully eliminated.
+The complete learning curves for all 26 games are shown in Figure 9.
+Figure 9: Learning curves for ScaleZero (MT) vs.
+UniZero (ST) on the Atari26 benchmark.
+This figure shows the full per-game learning curves (mean and 95% confidence interval) for the multitask ScaleZero agent compared against 26 single-task UniZero baselines.
 Appendix C DMC Experiment Details
 C.1 Benchmark Setup
-We use the DeepMind Control Suite (DMC) Tunyasuvunakool et al.
-(2020) as the evaluation benchmark for continuous control tasks.
-The experiments were conducted on a suite of 18 tasks (e.g., walker_walk, cheetah_run).
-Following a similar methodology to our Atari experiments, we compare a single multitask ScaleZero (MT) model against a baseline of 18 single-task UniZero (ST) models, each trained independently on one task.
-All ScaleZero variants discussed in this section are multitask (MT) models.
-The model input consists of low-dimensional state vectors provided by the environment, and the action space is a continuous vector normalized to the [-1, 1] range.
-The evaluation metric is the Average Return over 3 random seeds.
+This section details the experimental setup for the DeepMind Control (DMC) Suite (Tunyasuvunakool et al., 2020), which serves as our benchmark for continuous control as presented in Section 4.1.
+Experiments were conducted on a suite of 18 tasks (e.g., walker_walk, cheetah_run), following the experimental setup of Pu et al.
+(2024).
+Following the methodology of our Atari experiments, we compare a single ScaleZero (MT) model against a baseline of 18 individually trained UniZero (ST) models.
+Model inputs are low-dimensional state vectors from the environment, and the action space is a continuous vector normalized to [1,1][-1,1].
+Performance is measured as the average return over 8 evaluation episodes, with results averaged across 2 random seeds.
 C.2 Performance Comparison: ScaleZero vs. UniZero
-As shown in Figure 10, the multitask ScaleZero model achieves returns comparable or superior to the single-task UniZero baseline on the majority of DMC tasks.
-This result demonstrates that the ScaleZero architecture is applicable to control problems involving high-dimensional continuous state and action spaces.
-We hypothesize that this performance stems from the model learning shared physics priors and control primitives across different robotic morphologies.
-The expert modules within the MoE architecture may specialize in different types of dynamics (e.g., balancing, locomotion), which are then combined by the gating network based on the current task and state.
+As summarized in 1(b) and detailed in Figure 10, the multitask ScaleZero model achieves performance competitive with, and often superior to, the single-task UniZero baseline.
+Notably, ScaleZero achieves a higher median score, indicating robust generalist performance across the majority of tasks rather than excelling on only a few.
+This validates that the architectural principles of ScaleZero are effective for complex, state-based continuous control problems.
+We hypothesize this success stems from the MoE layers learning to specialize in shared physical priors and control primitives (e.g., balancing vs.
+locomotion), which are then composed by the router based on the task.
 Figure 10: Learning curves of ScaleZero (MT) vs.
 UniZero (ST) on the DMC18 benchmark.
-This figure compares the performance of the multitask ScaleZero and single-task UniZero models on 18 DMControl tasks.
-Solid lines represent the mean performance over 3 random seeds, and the shaded area indicates the 95% confidence interval.
-C.3 Efficiency Evaluation of ScaleZero-DPS
-Figure 11 compares the per-task learning curves of ScaleZero-DPS and the standard ScaleZero model to illustrate the difference in sample efficiency.
-The Figure shows that ScaleZero-DPS (orange line) exhibits a faster learning rate on most tasks, achieving equivalent performance levels with fewer environment steps.
-The asterisks (*) in the figure mark tasks that were identified as "solved" by the DPS policy and for which training was terminated early, showing the dynamic resource allocation mechanism in effect.
-Figure 11: Performance and efficiency comparison of ScaleZero-DPS vs.
+This figure provides the full per-task learning curves, comparing the performance of the single multitask ScaleZero agent against 18 specialized UniZero baselines.
+Solid lines represent the mean performance over 2 random seeds, and the shaded area indicates the 95% confidence interval.
+C.3 Efficiency Evaluation of Dynamic Parameter Scaling (DPS)
+This section provides the detailed learning curves that substantiate the efficiency gains of Dynamic Parameter Scaling (DPS), as reported in Section 4.2.
+Figure 11 compares the per-task sample efficiency of ScaleZero-DPS against the standard ScaleZero model.
+The curves demonstrate that ScaleZero-DPS (orange) learns significantly faster on most tasks, achieving target performance levels with fewer environment interactions.
+This validates the claim that DPS leads to a substantial reduction in sample complexity.
+The asterisks (*) denote tasks where training was terminated early by the DPS policy upon reaching a "solved" threshold, visually confirming the dynamic allocation of computational resources away from mastered tasks.
+Figure 11: Sample efficiency comparison of ScaleZero-DPS vs.
 standard ScaleZero on DMC18.
@@
 environment steps) for both models across 18 tasks.
-An asterisk (*) indicates that the task’s training was terminated early by the DPS policy.
+The faster rise of the ScaleZero-DPS curves (orange) illustrates its superior sample efficiency.
+An asterisk (*) indicates that the DPS policy halted training for that task, demonstrating the dynamic resource allocation mechanism in action.
+C.4 Mechanistic Analysis of Dynamic Parameter Scaling
+To understand the mechanisms by which Dynamic Parameter Scaling (DPS) achieves its sample efficiency, we conduct a granular analysis of the agent’s internal parameter dynamics.
+This section presents two complementary visualizations: a time-series plot showing the continuous evolution of key metrics (Figure 12), and a suite of "importance matrices" providing discrete snapshots at critical training junctures (Figure 13).
+C.4.1 Setup
+Time-Series Dynamics (Figure 12): This figure provides a macroscopic view of the training process.
+The top panel correlates high-level metrics—normalized performance scores and the number of solved tasks—with the progression of training stages.
+The bottom panel offers a view into the internal mechanics by tracking the average learnable importance (‘alpha‘ scale) of LoRA adapters.
+It specifically compares the behavior of adapters in the first Transformer layer (Layer 0, solid lines) versus the last (Layer 3, dashed lines), highlighting the emergence of different dynamic behaviors at different model depths.
+Stage-Wise Importance Matrices (Figure 13): This multi-panel figure offers a more granular, event-based perspective.
+Each matrix is an "importance matrix" where a cell at ‘(row r, column c)‘ shows the average ‘alpha‘ scale of the adapter introduced in Stage ‘c‘, as measured at the end of Stage ‘r‘’s training.
+This reveals how the model re-evaluates the contribution of all historical adapters after each phase of targeted learning.
+To provide a comprehensive view, this analysis is presented for:
+  • The overall average across all layers and adapter types (Panel a).
+  • A per-layer breakdown for Layer 0 and Layer 3 to show hierarchical differences (Panels b-c).
+  • A per-type breakdown for adapters applied to the Query, Key, Value, and output Projection linear layers to reveal functional specialization (Panels d-g).
+C.4.2 Results and Analysis
+The analysis of these figures reveals several key dynamics of the DPS-managed learning process.
+The time-series plot (Figure 12) illustrates the staged training protocol.
+In this particular experiment, the stage transitions (vertical dashed lines) are triggered by a pre-allocated iteration budget, consistent with the method’s budget-based trigger mechanism.
+These capacity expansions occur as overall performance and the number of solved tasks increase over the course of training.
+The divergence in ‘alpha‘ scales between Layer 0 and Layer 3 suggests that the model applies different update strategies to shallow versus deep layers as training progresses.
+The importance matrices in Figure 13 provide further detail on this behavior:
+  • Overall Strategy (Panel a): The data is consistent with a dual mechanism of knowledge retention and plasticity.
+The importance scale of the foundational adapter from Stage 0 (column 0) is not static; it increases from 1.0 to 1.097 by the end of training.
+This suggests that the model increases the relative contribution of its initial knowledge as it encounters more complex tasks.
+Concurrently, the diagonal values show how the model focuses on the newest adapter during each stage, which is then re-weighted in subsequent stages as its knowledge is integrated.
+  • Hierarchical Differentiation (Panels b vs.
+c): A notable difference is observed between Layer 0 and Layer 3.
+In Layer 0 (Panel b), the importance of the initial adapter (cell (4,0)) exhibits high stability with moderate growth (to 1.058), indicating its role as a stable foundation for low-level features.
+In contrast, Layer 3 (Panel c) shows more pronounced dynamic re-weighting.
+The foundational adapter’s importance grows more significantly (to 1.099), suggesting that higher-level, abstract representations rely heavily on and amplify this core knowledge.
+The greater variance in Layer 3’s alpha values indicates that later layers are more involved in adaptive, task-specific strategy modifications.
+  • Functional Differentiation by Layer Type (Panels d-g): The analysis reveals distinct behaviors based on which linear layer within the attention mechanism an adapter modifies.
+    • Query (Q) and Key (K) Adapters (Panels d, e): Adapters applied to the Q and K layers, which produce the representations for calculating attention scores, generally exhibit importance scales that increase over time.
+For instance, the Stage 0 Key adapter’s scale reaches 1.104.
+This suggests a continuous strengthening of the mechanisms that determine token-to-token relevance.
+    • Value (V) and Projection (Proj) Adapters (Panels f, g): In contrast, adapters for the V layer (which provides the content to be aggregated) and the output Projection layer (which combines attention head outputs) show more complex dynamics, including suppression (alpha < 1.0).
+For example, the Value adapter from Stage 3 is down-weighted to 0.889 after Stage 4.
+This behavior may indicate a more cautious integration of new value-based information or a re-balancing of attention head outputs to integrate new skills without disrupting existing ones.
+In summary, the DPS method facilitates a structured learning process.
+The model appears to preserve and amplify foundational knowledge, primarily in early layers, while enabling dynamic, adaptive strategy-switching in later layers.
+Furthermore, it modulates the importance of adapters based on their specific function within the self-attention mechanism (Query/Key vs.
+Value/Projection).
+This combination of knowledge retention and targeted, hierarchical plasticity offers a mechanistic explanation for the method’s observed sample efficiency.
+Figure 12: Internal dynamics of ScaleZero-DPS during training.
+(Top) Performance and task-solving progress across stages.
+(Bottom) Average ‘alpha‘ scales for adapters in the first (Layer 0, solid) and last (Layer 3, dashed) layers, showing the emergence of hierarchical specialization.
+(a) Overall Average
+(b) Layer 0 Only
+(c) Layer 3 Only
+(d) Query Adapters
+(e) Key Adapters
+(f) Value Adapters
+(g) Projection Adapters
+Figure 13: Evolution of Adapter Importance Across Training Stages and Model Components.
+Each matrix displays the learned importance (average ‘alpha‘ scale) of adapters.
+A cell at ‘(row r, col c)‘ represents the importance of the adapter from Stage ‘c‘ after Stage ‘r‘’s training is complete.
+(a) The overall average shows a clear pattern of knowledge preservation (column 0) and targeted plasticity (diagonal).
+(b-c) The layer-specific breakdown reveals that foundational knowledge is more rigidly preserved in early layers (Layer 0) while later layers (Layer 3) are more dynamic and adaptive.
+(d-g) The type-specific breakdown shows functional differentiation in how adapters are utilized over time, reflecting their distinct roles in the attention mechanism.
 Appendix D Jericho Experiment Details
@@
 (b) Detective
-Figure 12: First five steps of expert interaction trajectories for Zork1 and Detective.
+(c) Omniquest
+(d) Acorncourt
+Figure 14: First five steps of expert interaction trajectories across environments.
 At each step, the environment supplies a set of valid candidate actions (used in our experiments); however, additional legal actions outside this set may also be executed, so the action space is, in principle, unbounded.
@@
 Specifically, ScaleZero converges faster on Acorncourt, Omniquest, and Detective; on the most challenging Zork1, it initially lags but accelerates after  150k environment interactions to reach parity with UniZero.
-We hypothesize that this stems from multi-task training inducing transferable language priors and interaction routines, which help filter feasible commands, compress the open-ended action space, and reduce fruitless exploration.
+We hypothesize that this stems from multitask training inducing transferable language priors and interaction routines, which help filter feasible commands, compress the open-ended action space, and reduce fruitless exploration.
 Appendix E Gradient Analysis in MoE
 E.1 Experimental Analysis
-Given the strong performance of Mixture-of-Experts (MoE) in multitask settings, a natural question arises: Why MoE?
-We aim to investigate, both experimentally and theoretically, the underlying mechanisms behind MoE’s superior performance in multitask reinforcement learning.
-Specifically, our experimental setup is as follows.
+To understand why Mixture-of-Experts (MoE) architectures excel in multitask settings, this study investigates the core mechanisms driving their performance.
+Focusing on multitask reinforcement learning, we conduct a dual analysis combining theoretical inquiry with empirical validation.
+The following section details the experimental protocol designed for this purpose.
 E.1.1 Experiment 1: Analyzing Gradient Conflicts in MoE-based Transformers
@@
 We compare two baseline methods with different backbones:
-(1). Naive Transformer: The backbone consists of four standard Transformer blocks.
-(2).
-MoE-based Transformer: The backbone also consists of four Transformer blocks, but the MLP layer in each block is replaced with an MoE layer, which comprises one shared expert and eight non-shared experts(Liu et al., 2024b).
+(1) Naive Transformer: The backbone consists of four standard Transformer blocks.
+(2) MoE-based Transformer: The backbone also consists of four Transformer blocks, but the MLP layer in each block is replaced with an MoE layer, which comprises one shared expert and eight non-shared experts(Liu et al., 2024).
 Shared experts provide cross-task general representations, enhancing generalization ability and ensuring training stability, while non-shared experts learn task-specific representations to strengthen the model’s discriminability and task adaptability.
@@
 We choose the maximum pairwise cosine similarity because it directly identifies the most severe gradient conflict between any two tasks.
-Unlike averaging operation, which can hide critical issues, the maximum value pinpoints the ’bottleneck pair’ that most significantly impedes stable multi-task training and overall convergence, even if other tasks cooperate well.
+Unlike averaging operation, which can hide critical issues, the maximum value pinpoints the ’bottleneck pair’ that most significantly impedes stable multitask training and overall convergence, even if other tasks cooperate well.
 In MoE, when computing the gradient of an entire layer, if task $A$ selects expert $i$ while task $B$ selects expert $j$ with $i\neq j$, then the gradient of task $B$ on expert $i$ is filled with zero.
@@
 Figure 16: Comparison of gradient conflicts between MoE and MLP baselines across different components. MoE-based Transformer exhibits fewer gradient conflicts in MoE input and Moe layer.
-Observation 1: As shown in Figure˜14(a), the MoE-based Transformer exhibits fewer gradient conflicts than the MLP counterpart.
+Observation 1: As shown in Figure˜16(a), the MoE-based Transformer exhibits fewer gradient conflicts than the MLP counterpart.
 (b) further shows that introducing MoE reduces gradient conflicts at the layer input, implying that MoE alleviates conflicts in other components to some extent.
@@
 Previous studiesChen et al.
-have shown that, under supervised learning, MoE can implicitly uncover latent cluster structures within input space.
+(2022) have shown that, under supervised learning, MoE can implicitly uncover latent cluster structures within input space.
 However, when faced with non-stationary data distributions generated through agent-environment interactions, it remains unclear whether MoE experts still differentiate effectively.
-To answer the question, we analyze the entropy of expert selection distributions, which quantifies the uncertainty of the choice of which expert to use for a task, with low entropy indicating high specialization and decisive selection, while high entropy implies uncertainty or an average utilization across multiple experts.
-We also recorded the Wasserstein distance Rüschendorf (1985) between the expert selection distributions of different tasks, where smaller values indicate a greater proximity to the expert selection of two tasks.
+To answer the question, we analyze the entropy of expert selection distributions, which quantifies the uncertainty of the choice of which expert to use for a task, with low entropy indicating high specialization and decisive selection, while high entropy implies uncertainty or an average utilization across experts.
+We also record the Wasserstein distance  (Rüschendorf, 1985) between the expert selection distributions of different tasks, where smaller values indicate a greater proximity to the expert selection of two tasks.
 We aim to quantify this relationship to find the underlying connections in expert selection among various tasks.
@@
 and Wasserstein Distance between task $i$ and task $i^{\prime}$ ($W_{i,i^{\prime}}$)
-We consider two sizes $S$: immediate = 100, short = 1,000.555Note that at the very beginning of training, when the window cannot be fully populated, we use the available data within the window, which may lead to larger fluctuations at early steps.
-We present the entropy results in Figure˜15.
-Due to space limitations, we show in Figure˜18 the Wasserstein distances between expert selections of different tasks at multiple training stages.
-Figure 15: Line plot showing the evolution of expert selection entropy in a multi-task learning setting with eight tasks.
+We consider two sizes $S$: immediate = 100, short = 1,000.444Note that at the very beginning of training, when the window cannot be fully populated, we use the available data within the window, which may lead to larger fluctuations at early steps.
+We present the entropy results in Figure˜17.
+Due to space limitations, we show in Figure˜20 the Wasserstein distances between expert selections of different tasks at multiple training stages.
+Figure 17: Line plot showing the evolution of expert selection entropy in a multitask learning setting with eight tasks.
 The dashed lines correspond to the entropy values of individual tasks, and the solid red line represents the aggregated entropy across all tasks.
@@
 Observation 2: MoE differentiation and expert specialization.
-As shown in Figure˜15, as tasks progress, the entropy of the expert distribution gradually decreases, indicating a reduction in distributional uncertainty.
+As shown in Figure˜17, as tasks progress, the entropy of the expert distribution gradually decreases, indicating a reduction in distributional uncertainty.
 This suggests that expert selection becomes concentrated on a few outcomes, reflecting lower uncertainty.
-Furthermore, in Figure˜18, we visualize the expert selection distributions of different tasks at multiple training stages, offering additional evidence of the progressive specialization in expert utilization.
+Furthermore, in Figure˜20, we visualize the expert selection distributions of different tasks at multiple training stages, offering additional evidence of the progressive specialization in expert utilization.
 E.1.3 Experiment 3: Analyzing Gradient Conflicts between shared expert and non-shared expert
@@
 The plot shows the evolution of gradient conflicts for the shared expert and 8 non-shared experts during training interations.
-The logarithmic y-axis reveals that different experts experience varying levels of gradient conflicts, with the shared expert and individual experts (expert0–expert7) demonstrating distinct conflict patterns, indicating effective expert specialization in the multi-task learning framework.
-Observation 3: As shown in Figure˜16, the shared expert exhibits significantly higher gradient conflicts compared to the task-specific experts.
+The logarithmic y-axis reveals that different experts experience varying levels of gradient conflicts, with the shared expert and individual experts (expert0–expert7) demonstrating distinct conflict patterns, indicating effective expert specialization in the multitask learning framework.
+Observation 3: As shown in Figure˜18, the shared expert exhibits significantly higher gradient conflicts compared to the task-specific experts.
 Among the non-shared experts, the level of conflicts does not show substantial differences.
@@
 We further investigated the conflicts among different experts within MoE.
-Results is shown in Figure˜16 We found that shared experts bear most of the gradient conflicts, whereas individual (non-shared) experts experience almost no conflict.
+Results is shown in Figure˜18 We found that shared experts bear most of the gradient conflicts, whereas individual (non-shared) experts experience almost no conflict.
 A plausible explanation is that in a standard gating MoE, different samples are routed to different experts, so each expert primarily receives gradient updates from a specific type of sample.
@@
 By building upon existing foundational theories of Mixture-of-ExpertsChen et al.
-, we present a formal derivation that explains precisely when and how this is achieved.
+(2022), we present a formal derivation that explains precisely when and how this is achieved.
 E.2.1 Analysis
 Previews workChen et al.
-theoretically demonstrated that, in a single-task supervised learning setting, when the input distribution exhibits distinguishable cluster structures, the sparse gating Mixture-of-Experts (MoE) router can automatically learn the cluster center features and route samples to the most suitable expert, thereby significantly improving performance.
-We further hypothesize that, in multi-task learning, the input spaces of different tasks naturally correspond to distinct clusters and the signal for the cluster center can be task ID or the latent pattern on the state space.
-Under this assumption, the conclusions from the original work can be directly extended to the multi-task setting.
+(2022) theoretically demonstrated that, in a single-task supervised learning setting, when the input distribution exhibits distinguishable cluster structures, the sparse gating Mixture-of-Experts (MoE) router can automatically learn the cluster center features and route samples to the most suitable expert, thereby significantly improving performance.
+We further hypothesize that, in multitask learning, the input spaces of different tasks naturally correspond to distinct clusters and the signal for the cluster center can be task ID or the latent pattern on the state space.
+Under this assumption, the conclusions from the original work can be directly extended to the multitask setting.
 Next, we adapt Chen et al.
-theoretical analysis to the reinforcement learning context using a tabular example.
-Specifically, we treat each task as an independent Markov Decision Process (MDP) and construct a multi-task MDP set with a shared state space but task-specific transitions.
+(2022) theoretical analysis to the reinforcement learning context using a tabular example.
+Specifically, we treat each task as an independent Markov Decision Process (MDP) and construct a multitask MDP set with a shared state space but task-specific transitions.
 In this setting, we derive the gradient conflicts that arise in the MoE layer.
 We also discuss the case where no clustering signal exists between tasks and point out that the introduction of task embeddings can significantly enhance the clustering structure between tasks, thereby promoting the isolation of expert selection across different tasks in MoE.
-E.2.2 Problem setting: multi-task MDP with task-specific latent clusters
+E.2.2 Problem setting: multitask MDP with task-specific latent clusters
 We consider a Markov decision process (MDP) defined by the tuple $(\mathcal{S},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ the action space, $P(s^{\prime}\!\mid\!s,a)$ the transition kernel, $R(s,a)$ the reward function, and $\gamma\in[0,1)$ the discount factor.
@@
 E.2.3 Learning Dynamics Analysis
-We follow an analysis path similar to prior workChen et al. , dividing the learning process into two stages: an early expert exploration stage and a later router learning stage.
+We follow an analysis path similar to prior workChen et al. (2022), dividing the learning process into two stages: an early expert exploration stage and a later router learning stage.
 Initialization.
@@
 Expert Exploration Stage.
-According to [Chen et al. , Lemma E.3], the zero initialization of the router implies that at the beginning of training ($t=0$),
+According to [Chen et al. (2022), Lemma E.3], the zero initialization of the router implies that at the beginning of training ($t=0$),
 $\max_{m\in[M]}\left|P(m_{i,t}=m)-\frac{1}{M}\right|=\tilde{O}(\sigma_{0}^{1.5})\quad\text{for all }i\in[n],\,m\in[M].$
@@
 Under gradient descent updates, the weight vector $w_{m}$ of expert $m$ will predominantly grow along the direction of its preferred basis vector $v_{k_{m}^{*}}$.[Chen et al.
-, Lemma E.5] In other words, during training, an expert progressively specializes in modeling a specific cluster $c_{k}$.
+(2022), Lemma E.5] In other words, during training, an expert progressively specializes in modeling a specific cluster $c_{k}$.
 Building upon this, we derive an upper bound on the gradient conflicts during the Expert Exploration Stage.
@@
 Router Learning Stage
-During the router-learning stage, the sparse MoE layer exhibits two decisive properties that justify its use in a multi-task setting.
+During the router-learning stage, the sparse MoE layer exhibits two decisive properties that justify its use in a multitask setting.
 First, the gating network rapidly identifies the latent task structure: after only
@@
 iterations, its weight vectors $\bm{\theta}_{m}$ become strongly aligned with the cluster-center signal $\mathbf{c}_{k}$ of every task $k$ , while simultaneously suppressing spurious correlations with label signals and noise.[Chen et al.
--lemma E.14] Consequently, any input drawn from task $k$ is routed to the corresponding subset of experts $\mathcal{M}_{k}$ with probability $1-o(\frac{1}{d})$, even though the model is never given explicit task labels.
+(2022)-lemma E.14] Consequently, any input drawn from task $k$ is routed to the corresponding subset of experts $\mathcal{M}_{k}$ with probability $1-o(\frac{1}{d})$, even though the model is never given explicit task labels.
 [Chen et al.
--Lemma E.18]
-Second, each expert in $\mathcal{M}_{k}$ remains tightly specialised to its assigned task, but only chance-level performance on all other tasks. [Chen et al. -Lemma E.12,Lemma 5.2].
-Taken together, these two properties guarantee automatic task separation and expert specialisation without external supervision, making the MoE layer an effective backbone for multi-task problems in which tasks form separable clusters in the input space.
+(2022)-Lemma E.18]
+Second, each expert in $\mathcal{M}_{k}$ remains tightly specialised to its assigned task, but only chance-level performance on all other tasks. [Chen et al. (2022)-Lemma E.12,Lemma 5.2].
+Taken together, these two properties guarantee automatic task separation and expert specialisation without external supervision, making the MoE layer an effective backbone for multitask problems in which tasks form separable clusters in the input space.
 Theorem E.3 (Expected Gradient Conflict on Task-specific Expert Sets).
@@
 This visualization reveals how task-specific expert utilization patterns emerge and evolve during training.
-E.2.4 Task Embedding for Improved Task Separation in Multi-Task Learning
-In multi-task learning, tasks often share overlapping input spaces and underlying features (e.g., sentiment classification tasks using the same text encoder), making it difficult for the routing network to distinguish tasks.
+E.2.4 Task Embedding for Improved Task Separation in multitask Learning
+In multitask learning, tasks often share overlapping input spaces and underlying features (e.g., sentiment classification tasks using the same text encoder), making it difficult for the routing network to distinguish tasks.
 This overlap leads to routing confusion, where data from one task may be misrouted to experts specialized in other tasks, reducing performance.
@@
 Since $e_{\tau}$ is orthogonal across tasks (or constrained via a regularization term), it helps the routing network more clearly distinguish between tasks, even when inputs are similar.
-Task embedding also acts as a regularization term, reducing expert load imbalance and improving the stability and performance of multi-task MoE training.
+Task embedding also acts as a regularization term, reducing expert load imbalance and improving the stability and performance of multitask MoE training.
 Our theoretical analysis suggests that introducing distinct task embeddings provides crucial clustering signals, enabling the router to more effectively differentiate between tasks.
@@
 For future work, we plan to investigate more effective methods for explicitly injecting task-specific information.
-The goal is to empower the router to autonomously leverage the unique priors of each task, thereby further advancing multi-task learning performance.
+The goal is to empower the router to autonomously leverage the unique priors of each task, thereby further advancing multitask learning performance.
 Figure 20: Nine heatmaps illustrating the Wasserstein distances between expert selection distributions of different tasks at successive training interations with $S_{immediate}=100$.
@@
 Diagonal elements are removed to exclude self-comparisons.
-Figure 19: T-SNE Visualization.
+Figure 21: T-SNE Visualization of latent space.
 Each point denotes a 2D projection of a 768-dimensional latent embedding extracted from the world model.
@@
 The figure contains 8,000 data points collected from 8 distributed training processes, where each point corresponds to a single time step from the observation sequence (400 samples × 20 time steps per sample).
+Appendix F Detailed Related Work
+MCTS with Learned World Models.
+Planning within a learned latent space, popularized by MuZero (Schrittwieser et al., 2019), has become a dominant paradigm in reinforcement learning, building upon the success of AlphaZero (Silver et al., 2016; 2017).
+This approach enables a model to predict its own dynamics, rewards, and policies, allowing for powerful lookahead search within a compact, learned representation (Hubert et al., 2021; Antonoglou et al., 2021; Danihelka et al., 2022; Xuan et al., 2024).
+Recent advancements have incorporated Transformers as the backbone for these world models (Micheli et al., 2022; Robine et al., 2023; Pu et al., 2024; Zhou et al., 2024), significantly enhancing representational capacity and the ability to model long-horizon dependencies.
+However, the monolithic nature of these architectures, while powerful for single tasks, becomes a critical liability in heterogeneous multitask settings.
+They are prone to representational interference and plasticity collapse, where a single shared model struggles to accommodate the diverse and often conflicting properties of multiple tasks.
+Our work confronts this fundamental architectural bottleneck by systematically investigating and proposing a more robust design.
+Multi-Task Reinforcement Learning (MTRL).
+MTRL seeks to improve data efficiency and generalization by sharing knowledge across a distribution of tasks (Vithayathil Varghese & Mahmoud, 2020; D’Eramo et al., 2024).
+A common strategy is to use a shared backbone with task-specific heads to handle diverse action or observation spaces (Kumar et al., 2022; Hansen et al., 2023; Lee et al., 2022; Gallouédec et al., 2024).
+To mitigate inter-task interference within the shared backbone, various approaches have been proposed.
+Architectural methods aim to disentangle knowledge through context-based conditioning (Rakelly et al., 2019; Sodhani et al., 2021), learnable modulation modules that adapt network activations (Schmied et al., 2023), or explicit modularization with task-aware routing and parameter composition (Sun et al., 2022; He et al., 2023).
+In parallel, optimization-based methods focus on managing gradient conflicts at the training level by projecting gradients to avoid negative interference or re-weighting task losses dynamically (Fernando et al., 2023; Lin et al., 2024; Ma et al., 2023).
+However, these strategies have been predominantly studied in model-free RL or supervised settings, largely outside the domain of latent space planning, where the additional challenge of disentangling dynamics prediction is paramount.
+Few works have investigated the unique failure modes of multitask learning in this context.
+We address this gap by analyzing plasticity-related metrics to identify the root cause of performance collapse—an architectural bottleneck in the world model—and propose a novel design to resolve it.
+Sparse and Parameter-Efficient Architectures.
+To overcome the limitations of dense, monolithic models, we turn to sparse and parameter-efficient architectures.
+Sparse models like Mixture-of-Experts (MoE) (Dai et al., 2024) increase model capacity at a near-constant computational cost through conditional computation.
+This provides a natural architectural prior for multitask specialization, as different tasks can be routed to different "expert" sub-networks (Obando-Ceron et al., 2024).
+Concurrently, parameter-efficient fine-tuning (PEFT) methods, notably Low-Rank Adaptation (LoRA) (Hu et al., 2022), enable lightweight, task-specific adjustments.
+By freezing the bulk of the model and training only a small subset of injected parameters, LoRA can efficiently adapt a model to new tasks or data distributions (Wang et al., 2023; Yang et al., 2025; Zhang et al., 2025).
+We identify a critical synergy between these two paradigms: MoE provides coarse-grained, input-dependent specialization at an architectural level, while LoRA offers fine-grained, efficient adaptation for targeted plasticity.
+The integration of these two approaches is a promising but underexplored frontier for building highly adaptable models (Dou et al., 2023; Huang et al., 2024).
+Our work bridges this gap by unifying MoE and LoRA within a transformer-based world model, creating a system that is both architecturally specialized and dynamically adaptable for large-scale MTRL.
+Appendix G Detailed Conclusion and Future Work
+G.1 Conclusion
+In this paper, we diagnosed the plasticity collapse that plagues unified world models in the context of heterogeneous multi-task planning.
+Through a principled exploration across the input, model, and optimization dimensions, we introduced ScaleZero, a novel architecture that integrates a sparse Mixture-of-Experts (MoE) backbone.
+This design creates specialized computational pathways, directly counteracting the representational bottlenecks that lead to performance collapse.
+Extensive experiments on a diverse suite of 48 tasks across 26 Atari games, 18 DMControl environments, and 4 Jericho worlds show that a single ScaleZero agent matches or surpasses the performance of specialized single-task agents trained from scratch.
+We also proposed a LoRA-based Dynamic Parameter Scaling (DPS) strategy, which dynamically allocates model capacity in response to task progress.
+In a proof-of-concept evaluation on DMControl, this strategy achieved competitive performance while reducing total environment interactions by around 28.5%, demonstrating its strong potential for improving sample efficiency in large-scale MTRL.
+G.2 Future Work
+Our work points to several promising research directions for building more capable and efficient generalist agents:
+  • Comprehensive DPS Evaluation: The Dynamic Parameter Scaling (DPS) strategy showed promising results on DMControl.
+A crucial next step is to rigorously evaluate its performance on more complex and diverse domains like Atari and Jericho.
+This would involve assessing its generalization capabilities, scalability to a larger number of tasks, and the computational overhead of its adaptive mechanism.
+  • Deeper MoE-LoRA Synergy: Our current approach uses MoE for architectural specialization and LoRA for dynamic capacity scaling.
+A deeper investigation into the synergy between these two could yield further benefits.
+For instance, one could explore using LoRA to adapt the MoE gating network itself, or using MoE routing decisions to inform which LoRA adapters should be activated, creating a more tightly integrated and adaptive system.
+  • Hybrid Online-Offline Learning: Our framework currently operates in a purely online learning setting.
+A powerful extension would be to combine it with large-scale offline pre-training.
+By initializing our ScaleZero model with weights from a foundation model pre-trained on vast datasets, we could potentially leverage broad prior knowledge for superior initialization, faster adaptation to new tasks, and higher peak performance, paving the way for truly generalist agents.
+We believe these directions offer a clear path toward developing the next generation of multi-task decision-making systems.
+Appendix H The Use of Large Language Models
+In the preparation of this manuscript, the large language model Gemini 2.5 Pro was employed to assist with language refinement.
+The primary goal was to enhance the clarity, precision, and readability of the text.
+The authors have reviewed and approved the final version.